### v1

# Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions

11institutetext: Katherine J. Pearce 22institutetext: Oden Institute & Dept. of Mathematics, University of Texas at Austin, 22email: katherine.pearce@austin.utexas.edu 33institutetext: Per-Gunnar Martinsson 44institutetext: Oden Institute & Dept. of Mathematics, University of Texas at Austin, 44email: pgm@oden.utexas.edu
Randomized Algorithms for Low-Rank Matrix and Tensor Decompositions
Katherine J. Pearce and Per-Gunnar Martinsson
1 Introduction
Numerical linear algebra (NLA) is a cornerstone of applied mathematics, scientific computing, and data science, with nascent algorithmic implementations dating back to the beginning of recorded history. Algorithms to accomplish NLA tasks, such as solving linear systems, computing eigenvalues and singular values, orthogonalizing and factorizing matrices, among many others, have been used in applications ranging from ancient Babylonian land surveying (Brezinski2022_NLA, Chapter 2) to Nobel Prize-winning work on computer vision AlexNet_2012. Throughout its long history, NLA algorithmic development has been driven by two competing imperatives, accuracy and efficiency, with the extant challenge to strike the right balance between them.
Origins of RNLA   The relatively recent emergence of randomized algorithms marks a major paradigm shift in NLA. Though Monte Carlo methods developed by Ulam and von Neumann as part of the Manhattan Project were made public in the late 1940s and 1950s Metropolis1987; Ulam1949; vonNeumann1951techniques, for decades, the NLA community was largely skeptical of randomization. Numerical analysts questioned the lack of theoretical guarantees for the accuracy of outputs of randomized algorithms, as well as their reproducibility. The general consensus was that the increase in efficiency afforded by randomization meant a sacrifice in accuracy that made the algorithms inherently unreliable, particularly because Monte Carlo methods were often inaccurate as a byproduct of the central limit theorem, cf. (martinsson2020, Section 1.2).
However, since the 1980s, that skepticism has been replaced by an abundance of work to develop randomized algorithms with provable accuracy and practical efficiency, a subfield of NLA that has grown into a research area in its own right, known as randomized numerical linear algebra (RNLA). More details on historical milestones in the RNLA methods of the 20th century can be found in halko2011; Kannan2017; Mahoney2011Randomized; martinsson2020; Tropp2015_MatrixConcentration; woodruff2015.
RNLA in the 21st century   By the early 2000s, randomization was being utilized in ground-breaking applications, such as randomized sampling for web-searching AchlioptasFiatKarlinMcSherry2001; ChakrabartiDomGibsonKleinbergKumarRaghavanRajagopalanTomkins1999, including the notable Google PageRank algorithm BrinPage1998Anatomy; HenzingerHeydonMitzenmacherNajork1999; PageBrinMotwaniWinograd1998. Randomized sampling was also used to approximate matrix operations and compute low-rank matrix decompositions Achlioptas2001; Frieze2004; papadimitriou2000latent. By the mid-2000s, a unifying framework for RNLA had begun to emerge.
Over the past two decades, RNLA has developed into a firmly established research area, garnering widespread attention from both practitioners and theorists. In 2006, a three-part series laid the theoretical groundwork for randomized sampling in approximate matrix multiplication DrineasKannanMahoney2006, low-rank matrix decomposition (QR and SVD) DrineasKannanMahoney2006b, and CUR decomposition DrineasKannanMahoney2006c. The same year, the separate work of DeshpandeRademacherVempalaWang2006; Deshpande2006AdaptiveSA investigated adaptive sampling for low-rank approximation, and in sarlos2006improved, a relative error SVD algorithm was proposed that utilized randomized subspace embeddings.
Numerical analysts and theoretical computer scientists also provided compelling empirical evidence of the improved performance of RNLA algorithms for practical applications Deshpande2006AdaptiveSA; Friedland2005FastML; Frieze2004; Liberty2007; martinsson2006interpolation; RokhlinTygert2008. These results inspired subsequent investigations of randomized algorithms expressly for linear algebra tasks, such as preconditioning or rangefinding, paired with strong probabilistic guarantees Avron2010; HalkoMartinssonShkolniskyTygert2011; halko2011; MartinssonRokhlinTygert2011; Rokhlin2010. Twenty years later, the scope of research in RNLA now encompasses nearly every aspect of linear algebra, including eigenvalue problems bhattacharjee2023; BravermanKrishnanMusco2022; KressnerPlestenjak2024; LinSaadYang2016; Tropp2025_OSI; Tropp2022; YuanGuLi2018; general linear solvers AvronDruinskyGupta2015; Avron2010; BaboulinLiRouet2015; BalabanovGrigori2025; ChenEpperlyTroppWebber2025; NakatsukasaTropp2024; orthogonalization BalabanovGrigori2022; DuerschGu2017; FukayaKannanNakatsukasaYamamotoYanagisawa2020; GrigoriTimsit2024; JangGrigori2025; matrix function approximation AmselChenGreenbaumMuscoMusco2024; BurkeGuttel2024; ChenGreenbaumMuscoMusco2023; CortinovisKressnerNakatsukasa2024; GuttelSchweitzer2023; PerssonKressner2023; UbaruChenSaad2017; trace estimation AvronToledo2011; GrattonTitley2018; HallmanIpsenSaibaba2023; PerssonCortinovisKressner2022; RoostaKhorasaniAscher2015; UbaruSaad2018; and many, many more.
Why the R in RNLA? The rise in popularity of randomized algorithms is due in large part to the commensurate progress made in computing technology. In June 2006, the performance of the world’s top supercomputer BlueGene/L was measured at 2.806 Peta-flop/s (floating point operations per second) on the High Performance Linpack (HPL) benchmark, a program that solves a large dense system of linear equations by LU with partial pivoting. As of November 2025, that rate has grown exponentially to 1.742 Exa-flop/s, achieved by today’s top supercomputer El Capitan top500list.
Modern computing architectures have made it possible to solve increasingly large-scale problems; however, as evidenced by both theory and practice, the performance of classical algorithms is limited by asymptotic complexity and the “curse of dimensionality” bellman1961adaptive. As a point of reference, the amount of data collected by the Event Horizon Telescope Collaboration to generate the first image of a black hole in 2019 was over 5 Petabytes Whitwam2019, which is almost the full storage capacity of El Capitan. In other words, classical matrix algorithms with cubic cost in the input size, such as orthogonalization, may be untenable given the large-scale problems that are now considered in scientific investigations. For the best practical performance, it is necessary to develop algorithms and implementations that can fully leverage computing resources.
Perhaps the greatest benefit of randomization is that it gives us the ability to restructure algorithms to attenuate computational demands on whichever computing resources most hinder performance, such as flops, matrix access, or data movement. For example, randomization admits algorithms that require just a single pass over the input data halko2011; LiYin2020SinglePassLU; Mahoney2011Randomized; Murray2023RandNLA; Tropp2017Practical; Tropp2018-MorePracticalSketching; TroppStreaming2019; Woolfe2008FastRandomized, which is especially advantageous for large-scale problems or streaming data. Randomization can also significantly reduce the number of arithmetic operations required by classical algorithms; e.g., for dense $m\times n$ overdetermined least squares problems with $n\ll m$, randomization yields solutions in roughly $O(mn+n^{3})$ operations vs. $O(mn^{2})$ with classical approaches Avron2010; Epperly2024; LacottePilanci2021; OzaslanPilanciArikan2019; PilanciWainwright2016; RokhlinTygert2008. In short, randomized algorithms produce accurate solutions with high probability while reducing working storage and communication costs, offering improved asymptotic complexity, admitting straightforward parallelization and GPU acceleration, and enabling practical investigations that would be otherwise impossible. For an expanded discussion, see (martinsson2020, Sections 1.3-1.4).
RNLA for tensors   The successful incorporation of randomization into classical matrix algorithms inspired similar work to exploit randomization for fast tensor computations. Originally proposed in 1927 as a means of working with multi-way physical and chemical measurements Hitchcock1927TheEO; Hitchcock1927Multiple, tensors are now ubiquitous in scientific computing and data science. Applications that involve tensorial data include 3D image reconstruction Bendory2023Autocorrelation; Ghosh2022; Kam1980Reconstruction; Levin2018AbInitio; Zhang2024MomentMetrics; signal processing Chen2021_WirelessComm; CichockMandicLathauweri2015; Comon2001TensorDecompositions; Marot2008TensorDenoising; Miron2020; Sidiropoulos2017; quantum chemistry and physics Bischoff2012ManyBodyWaveFunctions; Khoromskaia2015TensorNM; Peng2020; Pierce2021; Szalay2015; machine learning Aidini2023FewShotTensorCompletion; Hao2013; Ji2019_ML; Panagakis2024; Tao2005; Zhou2016GroupComponentAnalysis; high-dimensional partial differential equations (PDEs) BachmayrSchneiderUschmajew2016; BoelensVenturiTartakovsky2018; DektorRodgersVenturi2021; Khoromskij2015TensorPDEs; LubichOseledetsVandereycken2015; among many others. For detailed surveys of tensors and their applications, we refer the reader to Ballard_Kolda_2025; FriedlandTammali2015LowRankApprox; Grasedyck2013; KoldaBader2009; papalexakis2017.
Unfortunately, tensor practitioners suffer harsher consequences of the curse of dimensionality. For a $d$-way array of uniform size $N$, the storage and computational costs scale as $N^{d}$, quickly becoming unmanageable. As such, randomization has been essential to the development of efficient algorithms for tensor computations, popularized by several foundational works in the mid-2000s CaiafaCichocki2009; delaVega2005; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR. Two decades later, tensor practitioners increasingly rely on randomization for efficient tensor computations that are backed by the strong accuracy guarantees of well-established randomized matrix algorithms. Randomized algorithms have been developed for nearly every type of low-rank tensor decomposition in the literature, including the CP format BattaglinoBallardKolda2018; Biagioni2015; Erichson_2020; MaSolomonik2021FastAccurateRandomizedTensor; Reynolds2016; VervlietDeLathauwer2016RandomizedBlockSampling; Wang2023, the Tucker format AhmadiAsl2021Randomized; CheWeiYan2025EfficientTucker; KressnerPerisa2017RecompressionHadamard; Minster2024; MinsterSaibabaKilmer2020; Tsourakakis2010; Zhou2014, the Tensor Train format CheWei2019; CheWeiYan2024; Fonat2019; HuberSchneiderWolf2017RandomizedTT_SVD; LiYuBatselier2022FasterTensorTrainSparse, as well as other specialized representations BallaniGrasedyckKluge2013BlackBoxHTucker; BatselierYuDanielWong2018TensorNetworkRandomizedSVD; CastilloHaddockHartsockEtAl2025RandomizedKaczmarzTPRODUCT; CheWeiYan2024SketchBasedMultiplicativeUpdating; Robeva2016; TarzanaghMichailidis2018; WangCuiLi2024SVD_TensorWheel; Yuan2019; ZhangSaibabaKilmerAeron2018RandomizedTensorSVD; Zhou2014NTD.
Outline   In the following sections, we highlight recent computational advances in randomized algorithms for low-rank matrix and tensor decompositions.
Sections 1-4 focus on the matrix setting. We begin with a review of requisite linear algebra material and classical matrix decomposition algorithms. We next survey existing methods of randomized dimension reduction in addition to promising new developments in randomized sketching and sampling. We end the first part of our manuscript by discussing recent work on randomized algorithms for low-rank matrix decompositions, which are the foundation of our randomized algorithms for tensors in part two.
Sections 5-6 contain the necessary tensor prerequisites, as well as overviews of the two major tensor formats that we consider in our work: CP and Tucker. We then summarize key developments in randomized algorithms to compute low-rank tensor decompositions in these formats.
We close with some final thoughts and an outlook on the future of RNLA.
Scope   Our intent with this survey is to bring attention to important recent developments in the fast-growing body of work on RNLA. This rapid progress means that regular appraisal of the current research landscape is important for guiding new work, drawing connections to prior work, and keeping RNLA practitioners abreast of major achievements and open problems; however, it also means that important subject areas and contributions will necessarily be omitted.
We quickly attempt to rectify this by pointing the interested reader to a few references on topics beyond the scope of our work. For matrices, we do not consider the vast body of work on randomized algorithms for eigenvalues and eigenvalue problems (see, e.g., HeKressnerPlestenjak2025; NakatsukasaTropp2024; Schneider2024_phd) or optimization (see, e.g., derezinski2020debiasing; na2023hessian). We also suppress the details of most theoretical results in favor of more background material to make our presentation self-contained; see, e.g., Camano2025_OSI; Kannan2017; martinsson2020. We refer the reader to Murray2023RandNLA for a survey of RNLA that centers on software implementations, and to DerezinskiMahoney2024 for a survey of RNLA from the perspective of machine learning. For tensors, we do not consider the tensor train format or other more specialized tensor decompositions and refer the reader to Ballard_Kolda_2025; Grasedyck2013.
2 Matrix Preliminaries
In this section, we summarize notation and key concepts in linear algebra that are frequently referenced in our work. We introduce our notational conventions in Sect. 2.1 and summarize several fundamental matrix decompositions and algorithms to compute them in Sect. 2.2-2.4.
2.1 Notation
Throughout this manuscript, we work over the real numbers $\mathbb{R}$ or complex numbers $\mathbb{C}$ (using $\mathbb{F}$ to signify either), and we let $[m]$ denote the integers $1,2,\ldots,m$.
Vectors are denoted by bold lowercase Roman or Greek letters (e.g., $\bm{\mathsf{x}},\bm{\mathsf{\omega}}$), whereas matrices are denoted by bold uppercase letters (e.g., $\bm{\mathsf{X}},\bm{\mathsf{\Omega}}$). We use $\bm{\mathsf{0}}$ to denote the zero matrix and $\bm{\mathsf{I}}$ for the identity matrix, with their dimensions made explicit by subscripts when needed. We generally reserve Greek letters for random vectors and matrices.
To refer to coordinates of vectors, we use parentheses with subscripts, i.e. $(\bm{\mathsf{x}})_{i}$ refers to the $i$th coordinate of $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn}$. We also adopt the notation of Golub and Van Loan golub2013 to reference submatrices. Namely, if $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, and $I=[i_{1},i_{2},\dots,i_{k}]\subset[m]$ and $J=[j_{1},j_{2},\dots,j_{l}]\subset[n]$ are (row and column, resp.) index sets, then $\bm{\mathsf{A}}(I,J)$ denotes the $k\times l$ matrix
$\bm{\mathsf{A}}(I,J)=\begin{bmatrix}\bm{\mathsf{A}}(i_{1},j_{1})&\bm{\mathsf{A}}(i_{1},j_{2})&\dots&\bm{\mathsf{A}}(i_{1},j_{l})\\ \bm{\mathsf{A}}(i_{2},j_{1})&\bm{\mathsf{A}}(i_{2},j_{2})&\dots&\bm{\mathsf{A}}(i_{2},j_{l})\\ \vdots&\vdots&&\vdots\\ \bm{\mathsf{A}}(i_{k},j_{1})&\bm{\mathsf{A}}(i_{k},j_{2})&\dots&\bm{\mathsf{A}}(i_{k},j_{l})\\ \end{bmatrix}$.
The abbreviation $\bm{\mathsf{A}}(I,:)$ refers to the submatrix $\bm{\mathsf{A}}(I,[n])$, analogously for $\bm{\mathsf{A}}(:,J)$.
A vector $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn}$ is measured in the Euclidean or $\ell_{2}$-norm $\|\bm{\mathsf{x}}\|_{2}=\left(\sum_{i}|(\bm{\mathsf{x}})_{i}|^{2}\right)^{1/2}$. A matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ may be equipped with the corresponding operator norm, $\|\bm{\mathsf{A}}\|_{2}=\sup_{\|\bm{\mathsf{x}}\|=1}\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|_{2}$, or with the Frobenius norm $\|\bm{\mathsf{A}}\|_{\text{F}}=(\sum_{i,j}|\bm{\mathsf{A}}(i,j)|^{2})^{1/2}$.
The (Hermitian) transpose of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ is denoted by $\bm{\mathsf{A}}^{*}$, and the Moore-Penrose pseudoinverse of $\bm{\mathsf{A}}$ is denoted by $\bm{\mathsf{A}}^{\dagger}$. A matrix $\bm{\mathsf{U}}$ is said to be orthonormal if its columns are orthonormal, i.e. $\bm{\mathsf{U}}^{*}\bm{\mathsf{U}}=\bm{\mathsf{I}}$.
The trace of a square matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptn\times n}$ is the sum of its diagonal elements $\sum_{i=1}^{n}\bm{\mathsf{A}}(i,i)$. For matrices $\bm{\mathsf{A}},\bm{\mathsf{B}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, we can define the inner product $\langle\bm{\mathsf{A}},\bm{\mathsf{B}}\rangle=\text{trace}(\bm{\mathsf{A}}^{*}\bm{\mathsf{B}})$, so $\|\bm{\mathsf{A}}\|^{2}_{\text{F}}=\langle\bm{\mathsf{A}},\bm{\mathsf{A}}\rangle=\textup{trace}(\bm{\mathsf{A}}^{*}\bm{\mathsf{A}})$.
[…]
2.2 The Singular Value Decomposition
Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits a singular value decomposition (SVD), given by
$\underset{m\times n}{\bm{\mathsf{A}}}=\underset{m\times r}{\bm{\mathsf{U}}}\underset{r\times r}{\hskip 5.69054pt\bm{\Sigma}\hskip 5.69054pt}\underset{r\times n}{\bm{\mathsf{V}}^{*}},$ (1)
where $r=\min(m,n)$, $\bm{\mathsf{U}}$ and $\bm{\mathsf{V}}$ are orthonormal, and $\mathbf{\Sigma}$ is diagonal. The columns $\{\bm{\mathsf{u}}_{i}\}_{i=1}^{r}$ and $\{\bm{\mathsf{v}}_{i}\}_{i=1}^{r}$ of $\bm{\mathsf{U}}$ and $\bm{\mathsf{V}}$, respectively, are called the left and right singular vectors of $\bm{\mathsf{A}}$. The diagonal elements $\{\sigma_{i}\}_{i=1}^{r}$ of $\mathbf{\Sigma}$ are the singular values of $\bm{\mathsf{A}}$, ordered so that $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$. The rank of $\bm{\mathsf{A}}$ is equal to the number of nonzero singular values.
Suppose that the rank of $\bm{\mathsf{A}}$ is at least $k$. To obtain a rank-$k$ approximation of $\bm{\mathsf{A}}$, we can truncate its SVD after the first $k$ terms, defining $\bm{\mathsf{A}}_{k}=\sum_{i=1}^{k}\sigma_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{v}}_{i}^{*}$. By the Eckart-Young theorem eckart1936, $\bm{\mathsf{A}}_{k}$ is the best possible rank-$k$ approximation of $\bm{\mathsf{A}}$, with approximation error given by the singular values:
$\|\bm{\mathsf{A}}-\bm{\mathsf{A}}_{k}\|_{2}=\sigma_{k+1}\ \ \textup{and}\ \ \|\bm{\mathsf{A}}-\bm{\mathsf{A}}_{k}\|_{\text{F}}=\left(\sum_{j=k+1}^{r}\sigma_{j}^{2}\right)^{1/2}.$ (2)
Frequently, the Golub-Reinsch algorithm GolubReinsch1970 is used to compute an SVD, as in the built-in MATLAB function
$\begin{bmatrix}\bm{\mathsf{U}},\bm{\Sigma},\bm{\mathsf{V}}\end{bmatrix}=\texttt{svd}(\bm{\mathsf{A}}).$
We adopt this notation when referring to the computation of the SVD, as well as $\texttt{svd}(\bm{\mathsf{A}},k)$ to represent the function outputting the truncation $\bm{\mathsf{A}}_{k}$. Golub-Reinsch and related algorithms rely on QR decompositions, covered next.
2.3 The QR Decomposition
Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits a QR decomposition given by
$\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}=\underset{m\times r}{\bm{\mathsf{Q}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{R}}\hskip 5.69054pt},$ (3)
where $r=\min(m,n)$, $\bm{\mathsf{Q}}$ is orthonormal and $\bm{\mathsf{R}}$ is upper triangular.
Using a greedy algorithm that builds the QR factorization incrementally (e.g., with Householder reflections Householder1958; cf. Sect 2.4), we can halt the factorization after computation of the first $k$ terms to obtain a “partial” QR factorization:
$\underset{m\times n}{\bm{\mathsf{A}}}\approx\underset{m\times k}{\bm{\mathsf{Q}}_{k}}\underset{k\times n}{\hskip 5.69054pt\bm{\mathsf{R}}_{k}\hskip 5.69054pt}.$ (4)
We let
$\begin{bmatrix}{\bm{\mathsf{Q}},\bm{\mathsf{R}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{Q}}_{k},\bm{\mathsf{R}}_{k}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}},k),$ (5)
denote functions that yield either the full or partial QR factorization of $\bm{\mathsf{A}}$, respectively. Occasionally, if we need only the factor $\bm{\mathsf{Q}}$, whose columns form an orthonormal basis for the range of $\bm{\mathsf{A}}$, we will use the notation
$\bm{\mathsf{Q}}=\texttt{col}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \bm{\mathsf{Q}}_{k}=\texttt{col}(\bm{\mathsf{A}},k).$ (6)
2.4 The Interpolative and CUR Decompositions
Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ of rank $k$ admits interpolative decompositions (ID) and a CUR decomposition of the form
$\begin{array}[]{ll}\bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}}\underset{k\times n}{\bm{\mathsf{R}}},&\hskip 14.22636pt\textit{Row ID}\\ \bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{C}}}\underset{k\times n}{\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}},&\hskip 14.22636pt\textit{Column ID}\\ \bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{C}}}\underset{k\times k}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt}\underset{k\times n}{\bm{\mathsf{R}}},&\hskip 14.22636pt\textit{CUR}\end{array}$ (10)
where
$\bm{\mathsf{R}}=\bm{\mathsf{A}}(I,:)$
for $k$ linearly independent rows indexed by $I$;
$\bm{\mathsf{C}}=\bm{\mathsf{A}}(:,J)$
for $k$ linearly independent columns indexed by $J$; and $\bm{\mathsf{U}}=\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$, though in practice, $\bm{\mathsf{U}}$ is often approximated by $\bm{\mathsf{A}}(I,J)^{\dagger}$. The columns of $\bm{\mathsf{C}}$ or rows of $\bm{\mathsf{R}}$ are known as skeletons in an ID or CUR decomposition.
Typically, constructing the CUR decomposition is more ill-conditioned than the IDs, but a stable construction can be attained through QR factorizations of $\bm{\mathsf{C}}$ and $\bm{\mathsf{R}}$. Namely, let $\bm{\mathsf{Q}}_{\bm{\mathsf{C}}}=\texttt{col}(\bm{\mathsf{C}})$ and $\bm{\mathsf{Q}}_{\bm{\mathsf{R}}}=\texttt{col}(\bm{\mathsf{R}}^{*})$. The CUR factorization is then given by
$\bm{\mathsf{A}}=\bm{\mathsf{C}}\bm{\mathsf{U}}\bm{\mathsf{R}}=\bm{\mathsf{Q}}_{\bm{\mathsf{C}}}\bm{\mathsf{Q}}_{\bm{\mathsf{C}}}^{*}\bm{\mathsf{A}}\bm{\mathsf{Q}}_{\bm{\mathsf{R}}}\bm{\mathsf{Q}}_{\bm{\mathsf{R}}}^{*}.$ (11)
This stabilization tactic also works for the construction of the IDs, if needed.
The accuracy of a rank-$k$ ID or CUR approximation depends on the choice of skeletons. It was shown in DeshpandeRademacher2010 that it is possible to construct an ID satisfying
$\|\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\|_{\text{F}}\leq\sqrt{k+1}\|\bm{\mathsf{A}}-\bm{\mathsf{A}}_{k}\|_{\text{F}},$ (12)
where the coefficient $\sqrt{k+1}$ is optimal and cannot be improved; here, $\bm{\mathsf{A}}_{k}$ is the optimal rank-$k$ approximation given by the truncated SVD.
In practice, greedy pivoting algorithms are frequently used to choose skeletons for ID and CUR decompositions. We discuss several popular choices in the following sections, including a recently proposed algorithm that achieves (12) with the best known complexity.
QR with Column Pivoting (CPQR)
Column-pivoted QR (CPQR) is the algorithm behind the qr function from Sect. 2.3. In particular, column pivoting may be required for numerical stability, and the outputs of qr($\bm{\mathsf{A}}$) = $\bm{\mathsf{Q}},\bm{\mathsf{R}},P]$ represent the factorization $\bm{\mathsf{A}}\bm{\mathsf{P}}=\bm{\mathsf{Q}}\bm{\mathsf{R}}$ for some permutation matrix $\bm{\mathsf{P}}\in^{n\times n}$ such that $\bm{\mathsf{P}}=\bm{\mathsf{I}}(:,P)$. This factorization can be achieved through a sequence of permutations and Householder reflections that rank-one update the “active submatrix” of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ by greedy pivoting, based on squared column norms.
Assuming $m\geq n$, let $\bm{\mathsf{A}}^{(t)}$ represent the resulting matrix after $t$ steps of pivoting and updating for $t=0,\ldots,n-2$ (with $\bm{\mathsf{A}}^{(0)}=\bm{\mathsf{A}}$). At the $(t+1)$-th step, CPQR determines the column pivot $j_{t+1}$ by finding the column of maximal Euclidean norm in the active submatrix,
$j_{t+1}=\arg\max_{t+1\leq j\leq n}\|\bm{\mathsf{A}}^{(t)}([t+1:m],j)\|_{2}.$
The selected $k$ (column) skeletons are then $j_{1},\ldots,j_{k}]$. (Row skeletons can be computed via CPQR on $\bm{\mathsf{A}}^{*}$.) Once a pivot entry is selected, column $j_{t+1}$ is swapped with column $t+1$, the normalized reflection vector $\bm{\mathsf{v}}$ is formed, and the lower right block of the active submatrix, starting from the $(t+1)$-st main diagonal entry, is updated as $\bm{\mathsf{A}}^{(t+1)}=\bm{\mathsf{A}}^{(t)}-2\bm{\mathsf{v}}\bm{\mathsf{v}}^{*}\bm{\mathsf{A}}^{(t)}.$
In practice, CPQR chooses pivots in an order that reveals the rank of $\bm{\mathsf{A}}$, so that the chosen skeletons form a good basis for its row or column space. While there are adversarial cases for which CPQR can fail to be rank-revealing (e.g., Kahan-type matrices kahan1966; peters1975), these are rarely encountered, cf. trefethen1990. There are variations of CPQR with better theoretical guarantees, such as rank-revealing QR Chan1987; Ilse1994 or strong rank-revealing QR GuEis1996, but they are more costly to implement.
LU with Partial Pivoting
The LU algorithm with row-wise partial pivoting (LUPP) yields a factorization
$\underset{m\times m}{\hskip 11.38109pt\bm{\mathsf{P}}\hskip 11.38109pt}\underset{m\times n}{\bm{\mathsf{A}}}=\underset{m\times n}{\bm{\mathsf{L}}}\underset{n\times n}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt},$ (13)
where $\bm{\mathsf{L}}$ is lower triangular with main diagonal entries equal to 1 and off-diagonal entries bounded in magnitude by 1, and $\bm{\mathsf{U}}$ is upper triangular. The function $\bm{\mathsf{L}},\bm{\mathsf{U}},\bm{\mathsf{P}}]=\texttt{lu}(\bm{\mathsf{A}})$ represents the factorization (13) obtained from row-wise partial pivoting.
More specifically, to compute an LU factorization, row pivots are selected from the active submatrix of $\bm{\mathsf{A}}$ that is updated via Schur complements (e.g., (golub2013, Algorithm 3.2.1) (trefethen1997, Algorithm 21.1)). Let $\bm{\mathsf{A}}^{(t)}$ be the resulting matrix after $t$ steps of pivoting and updating for $t=0,1,\ldots,n-2$. At the ($t+1$)-st step of LUPP, the largest-magnitude column entry of the active submatrix is selected as the next pivot entry:
$i_{t+1}=\arg\max_{t+1\leq i\leq m}|\bm{\mathsf{A}}^{(t)}(i,t+1)|.$
The active submatrix is then updated: defining vectors $\bm{\mathsf{r}}=[t+2:m]$ and $\bm{\mathsf{c}}=[t+2:n]$,
$\bm{\mathsf{A}}^{(t+1)}(\bm{\mathsf{r}},\bm{\mathsf{c}})=\bm{\mathsf{A}}^{(t)}(\bm{\mathsf{r}},\bm{\mathsf{c}})-\frac{\bm{\mathsf{A}}^{(t)}(\bm{\mathsf{r}},t+1)\bm{\mathsf{A}}^{(t)}(t+1,\bm{\mathsf{c}})}{\bm{\mathsf{A}}^{(t)}(t+1,t+1)},$ (14)
where $\bm{\mathsf{A}}^{(t)}(t+1,t+1)$ has been updated with $\bm{\mathsf{A}}^{(t)}(i_{t+1},t+1)$ following row permutation.
This LUPP algorithm leads to an exponential upper bound on the entries of $\bm{\mathsf{U}}$, which is tight for certain matrices (e.g., Kahan-type matrices kahan1966; peters1975; sorensen2014). Thus, LUPP can fail to be rank-revealing, but such adversarial inputs are rarely encountered in practice trefethen1990. There exist more sophisticated pivoting strategies to lend rank-revealing properties to the LU factorization Chan84; Pan2000; Zhao2005 with better error bounds, but at significantly greater computational cost. Fortunately, we will see in Sect. 4.3 that randomization can equip LUPP with rank-revealing properties comparable to CPQR even in adversarial cases.
Truncated SVD for Near-Optimal Skeletons
Previously, the tight error bound (12) had been achieved by expensive random volume sampling DeshpandeRademacher2010, or computing at least $k$ SVDs to choose each pivot from an updated residual subspace. Recently, in Osinsky2025, the optimal error in (12) is achieved through a single rank-$k$ truncated SVD, with total complexity $O(mnk)$.
Let $\bm{\mathsf{Z}}\in\mathbb{F}^{m\times n}$ be an arbitrary rank-$k$ approximation of $\bm{\mathsf{A}}\in\mathbb{F}^{m\times n}$. We want to find columns $\bm{\mathsf{C}}\in\mathbb{F}^{m\times k}$ of $\bm{\mathsf{A}}$ and an interpolation matrix $\bm{\mathsf{W}}\in\mathbb{F}^{k\times n}$ such that
$\|\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\|_{\text{F}}\leq\left\|\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{W}}\right\|_{\text{F}}\leq\sqrt{k+1}\left\|\bm{\mathsf{A}}-\bm{\mathsf{Z}}\right\|_{\text{F}}.$ (15)
For the optimal error, we can take $\bm{\mathsf{Z}}$ to be the truncated SVD. Otherwise, if $\bm{\mathsf{Z}}$ is not a rank-$k$ SVD, we compute one singular value decomposition $\bm{\mathsf{Z}}=\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}.$ Define
$\widetilde{\bm{\mathsf{A}}}=\bm{\mathsf{A}}-\bm{\mathsf{A}}\bm{\mathsf{V}}\bm{\mathsf{V}}^{*},$
noting that $\bm{\mathsf{Z}}-\bm{\mathsf{Z}}\bm{\mathsf{V}}\bm{\mathsf{V}}^{*}=\bm{\mathsf{0}}$, and
$\|\widetilde{\bm{\mathsf{A}}}\|_{\text{F}}=\|(\bm{\mathsf{A}}-\bm{\mathsf{Z}})(\bm{\mathsf{I}}-\bm{\mathsf{V}}\bm{\mathsf{V}}^{*})\|_{\text{F}}\leq\|\bm{\mathsf{A}}-\bm{\mathsf{Z}}\|_{\text{F}}.$
Crucially, the construction of $\widetilde{\bm{\mathsf{A}}}$ requires only $O(mnk)$ operations.
Letting $\bm{\mathsf{W}}=(\hat{\bm{\mathsf{V}}}^{*})^{-1}\bm{\mathsf{V}}^{*}$, where $\hat{\bm{\mathsf{V}}}^{*}\in\mathbb{F}^{k\times k}$ is the sub-matrix of $\bm{\mathsf{V}}^{*}\in\mathbb{F}^{k\times n}$ corresponding to the $k$ column skeletons in $\bm{\mathsf{C}}$, we have
$\|\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{W}}\|_{\text{F}}=\|\widetilde{\bm{\mathsf{A}}}-\widetilde{\bm{\mathsf{C}}}\bm{\mathsf{W}}\|_{\text{F}},$
for $\widetilde{\bm{\mathsf{C}}}=\bm{\mathsf{C}}-\bm{\mathsf{A}}\bm{\mathsf{V}}\hat{\bm{\mathsf{V}}}^{*}.$ Since $\widetilde{\bm{\mathsf{A}}}$ and $\bm{\mathsf{W}}$ are not affected by replacing $\bm{\mathsf{V}}^{*}$ with any unitary transformation $\bm{\mathsf{Q}}\bm{\mathsf{V}}^{*}$, we can assume unitary transformations are applied to $\bm{\mathsf{V}}^{*}$ so that $\hat{\bm{\mathsf{V}}}^{*}$ is always upper-triangular. (In Osinsky2025, this is accomplished via Householder reflections.)
Assuming we have selected the first i1i-1 columns for $i\geq 1$, the approximation error at the $i$th step is given by
$\widetilde{\bm{\mathsf{A}}}^{(i)}=\widetilde{\bm{\mathsf{A}}}^{(i-1)}-\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,i}(\bar{\bm{\mathsf{V}}}^{*})_{1:(k-i+1),i}^{\dagger}\bar{\bm{\mathsf{V}}}^{*},$ (16)
where $\bar{\bm{\mathsf{V}}}^{*}\in\mathbb{F}^{(k-i+1)\times n}$ are the last ki+1k-i+1 rows of $\bm{\mathsf{V}}^{*}$, cf. (Osinsky2025, Equation 5). We choose the next column skeleton as the one that minimizes the error at the $i$th step, which we can do by selecting the column $j\geq i$ that minimizes the ratio
$\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,j}\|_{2}/\|\bar{\bm{\mathsf{V}}}^{*}_{1:(k-i+1),j}\|_{2}.$ (17)
By the upper triangularity of $\widehat{\bm{\mathsf{V}}}^{*}$ and the relation (16),
$\sum_{\ell=1}^{n}\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,\ell}\|^{2}_{2}=\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}\|^{2}_{\text{F}},$ (18)
and
$\sum_{\ell=1}^{n}\|\bar{\bm{\mathsf{V}}}^{*}_{1:(k-i+1),\ell}\|^{2}_{2}=\|\bar{\bm{\mathsf{V}}}^{*}\|^{2}_{\text{F}}=k-i+1.$ (19)
Then, using the fact that $\sum_{\ell}\alpha_{\ell}/\sum_{\ell}\beta_{\ell}\geq\min_{\ell}\alpha_{\ell}/\beta_{\ell}$ for non-negative $\alpha_{\ell}$ and positive $\sum_{\ell}\beta_{\ell}$, we can upper bound (17) by the ratio of (18) and (19), to estimate the error in the $i$th step in terms of the (i1)(i-1)th step:
$\|\widetilde{\bm{\mathsf{A}}}^{(i)}\|^{2}_{\text{F}}=\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}-\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,i}\bar{\bm{\mathsf{V}}}^{*}_{1:(k-i+1),i}\bar{\bm{\mathsf{V}}}^{*}\|_{\text{F}}^{2}\leq\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}\|^{2}_{\text{F}}+\frac{1}{k-i+1}\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}\|^{2}_{\text{F}}.$
If we add this up over all previous steps, we have
$\|\widetilde{\bm{\mathsf{A}}}^{(i)}\|^{2}_{\text{F}}\leq\|\widetilde{\bm{\mathsf{A}}}\|^{2}_{\text{F}}+\frac{i}{k-i+1}\|\widetilde{\bm{\mathsf{A}}}\|^{2}_{\text{F}}.$ (20)
Letting $i=k$ in (20), we achieve the desired bound in (12) in $O(mnk)$ operations. Full details can be found in Osinsky2025.
3 Randomized Dimension Reduction: Sketching and Sampling
In this section, we discuss ways that randomization can be integrated into classical algorithms in numerical linear algebra to compute low-rank decompositions more efficiently. Namely, we will often rely on special types of linear maps that map the row or column space of an input matrix to a smaller-dimensional space in such a way that the original geometry is roughly preserved, called a dimension reduction map (DRM). If the linear map is drawn from a random matrix distribution that achieves geometry-preserving dimension reduction with high probability, it is known as a randomized DRM.
Randomized dimension reduction is frequently associated with randomized embeddings, linear maps drawn from random matrix distributions that satisfy the following condition with high probability. Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be a matrix of rank $k\leq\min(m,n)$ and $\varepsilon\in(0,1)$ a distortion parameter. A linear map $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an embedding of $\bm{\mathsf{A}}$ with distortion $\varepsilon$ if
$(1-\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|\leq\|\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}\bm{\mathsf{x}}\|\leq(1+\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|\ \ \forall\ \bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn}.$ (21)
Usually, $k\leq d\ll m,$ so that the application of $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ results in significant dimension reduction. The minimal value of $d$ required for $\bm{\mathsf{\Gamma}}$ to satisfy (21) depends on the type of embedding $\bm{\mathsf{\Gamma}}$, the rank $k$ of $\bm{\mathsf{A}}$, and the distortion parameter $\epsilon$. As a classical example, the Johnson-Lindenstrauss lemma jlind84 establishes that $d\geq 8\log k/\varepsilon^{2}$ is sufficient for a random matrix $\bm{\mathsf{\Gamma}}$ with standard Gaussian entries to embed $k$ points in m so that (21) is satisfied.
More generally, if $\mathcal{S}$ is any probability distribution over linear maps $\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$, then $\bm{\mathsf{\Gamma}}\sim\mathcal{S}$ is a randomized embedding if (21) holds over $\mathcal{S}$ for all $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ with at least constant probability. If a randomized embedding can be constructed to map an unknown subspace into a lower-dimensional space (i.e. without knowing anything about the geometry of the input space except its dimension), then it is said to be an oblivious subspace embedding sarlos2006improved. It is often the case in practice that randomized DRMs are assumed to be randomized embeddings, but we will see in Sect. 3.1 that satisfying (21) is sufficient but not necessary to ensure the smaller-dimensional space is a good proxy for the input space.
In Sect. 3.1, we review randomized sketching, a procedure that produces random linear combinations of input matrix coordinates, which we say is “coordinate-mixing.” We then focus on dimension reduction by randomized sampling of matrix coordinates in Sect. 3.2. We end this section with an important application of randomized dimension reduction that will be relevant for our discussion of tensor decompositions: solving overdetermined least squares problems. For detailed treatments of randomized dimension reduction, the reader is referred to Camano2025_OSI; halko2011; martinsson2019randomized; martinsson2020; woodruff2015.
3.1 Randomized Matrix Sketching
The application of a randomized dimension reduction map $\bm{\mathsf{\Gamma}}$ to $\bm{\mathsf{A}}$ is known as randomized sketching. We reserve the term randomized sketching for random linear maps that are coordinate-mixing, vs. coordinate-sampling. The matrix product $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ is said to be a row sketch of $\bm{\mathsf{A}}$, and a column sketch $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is defined analogously.
Ideally, the random matrix used for sketching should roughly preserve the geometry of the row or column space of $\bm{\mathsf{A}}$. Randomized dimension reduction maps $\bm{\mathsf{\Gamma}}$ that are frequently utilized for sketching include:
  • Gaussian matrices ((halko2011, Section 4.1), (martinsson2020, Section 8.3), (woodruff2015, Theorem 2.3)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ with entries
    $\bm{\mathsf{\Gamma}}(i,j)\sim\mathcal{N}(0,1/d),$
    i.i.d. Gaussian entries drawn from a normal distribution with mean 0 and standard deviation $1/d$.
  • Subsampled randomized trigonometric transforms (SRTT) ((halko2011, Section 4.6), (martinsson2020, Section 9.3), tropp2011, (woodruff2015, Theorem 2.4)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
    $\bm{\mathsf{\Gamma}}=\sqrt{\frac{m}{d}}\bm{\mathsf{\Pi}}_{m}\bm{\mathsf{F}}\bm{\mathsf{\Phi}}\bm{\mathsf{\Pi}}_{m\rightarrow d}.$
    Here, $\bm{\mathsf{\Pi}}_{m\rightarrow d}\in\mathbb{F}^{\hskip 0.56905ptm\times d}$ is a uniformly random selection of $d$ out of $m$ rows, whereas $\bm{\mathsf{\Pi}}_{m}\in\mathbb{F}^{\hskip 0.56905ptn\times n}$ is a random permutation of the $m$ rows. The matrix $\bm{\mathsf{F}}\in\mathbb{F}^{\hskip 0.56905ptm\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{\Phi}}:=\textup{diag}(\phi_{1},\ldots,\phi_{n})$ has i.i.d. Rademacher random variable entries, i.e. when $\mathbb{F}=$, entries of $\pm 1$ chosen independently with equal probability.
  • Sparse sign matrices ((martinsson2020, Section 9.2), nelson2013, (woodruff2015, Theorem 2.5)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
    $\bm{\mathsf{\Gamma}}=\sqrt{\frac{m}{\zeta}}\begin{bmatrix}{\bm{\mathsf{s}}_{1},\ldots,\bm{\mathsf{s}}_{m}}\end{bmatrix}$
    for sparsity parameter $2\leq\zeta\leq d$, where each column $\bm{\mathsf{s}}_{j}\in\mathbb{F}^{\hskip 0.56905ptd}$ for $j=1,\ldots,m$ is filled with $\zeta$ independent Rademacher random variables at uniformly random coordinates. In other words, each column is an independent random vector having $\zeta$ nonzero entries of $\pm 1$ (when $\mathbb{F}=$) in uniformly random positions.
  • CountSketch matrices (Charikar2004; meng2013low; nelson2013): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as a sparse sign matrix with sparsity parameter $\zeta=1$. We caution that CountSketch matrices on their own can perform poorly in practice among randomized DRMs (see the required sketching dimension $d$ in Table 1).
  • SparseStack matrices (Camano2025_OSI; Cohen2016ObliviousSubspace; KaneNelson2014): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
    $\bm{\mathsf{\Gamma}}=\frac{1}{\sqrt{\zeta}}\begin{bmatrix}\rho_{1,1}\bm{\mathsf{e}}_{s_{1,1}}&\cdots&\rho_{1,m}\bm{\mathsf{e}}_{s_{1,m}}\\ \rho_{2,1}\bm{\mathsf{e}}_{s_{2,1}}&\cdots&\rho_{2,m}\bm{\mathsf{e}}_{s_{2,m}}\\ \vdots&\ddots&\vdots\\ \rho_{\zeta,1}\bm{\mathsf{e}}_{s_{\zeta,1}}&\cdots&\rho_{\zeta,m}\bm{\mathsf{e}}_{s_{\zeta,m}}\end{bmatrix}$,
    where $d=\zeta b$ for some block size $b$ and sparsity parameter $\zeta$, $\rho_{i,j}$ are i.i.d. Rademacher variables, $\bm{\mathsf{e}}_{i}\in\mathbb{F}^{b}$ is the $i$th standard basis vector, and $s_{i,j}\in[b]$ is chosen uniformly. In other words, the SparseStack test matrix is comprised of $\zeta$ independent copies of a (scaled) CountSketch matrix, stacked on top of one another.
  • SparseCol matrices (Camano2025_OSI): For consistency with the previous definitions of row sketching matrices as subspace embeddings, we give the transpose of the SparseCol matrix in (Camano2025_OSI, Definition 4.1). Namely, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ is defined by its $d$ rows of sparse random vectors
    $\bm{\mathsf{\omega}}=\sqrt{\frac{m}{\zeta}}\sum_{i=1}^{\zeta}\rho_{i}\bm{\mathsf{e}}_{s_{i}}\in\mathbb{F}^{m},$ (22)
    where $\zeta$ is the sparsity parameter, $\rho_{1},\ldots\rho_{\zeta}$ are i.i.d. Rademacher random variables, the indices $s_{1},\ldots,s_{\zeta}$ are sampled uniformly from $[m]$ without replacement, and $\bm{\mathsf{e}}_{s_{i}}$ is the $s_{i}$-th standard basis vector. In other words, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ is the transpose of the SparseCol matrix $\bm{\mathsf{\Gamma}}^{*}\in\mathbb{F}^{m\times d}$ given by
    $\bm{\mathsf{\Gamma}}^{*}:=\frac{1}{\sqrt{k}}\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{d}\end{bmatrix}$, (23)
    where $\bm{\mathsf{\omega}}_{j}\sim\bm{\mathsf{\omega}}$ are i.i.d.
  • SparseRTT matrices ((Camano2025_OSI, Section 4.1), tropp2011): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ defined as
    $\bm{\mathsf{\Gamma}}=\bm{\mathsf{S}}\bm{\mathsf{F}}\bm{\mathsf{D}},$ (24)
    where $\bm{\mathsf{D}}\in\mathbb{F}^{m\times m}$ is a random diagonal matrix populated with i.i.d. Rademacher random variables, $\bm{\mathsf{F}}\in\mathbb{F}^{m\times m}$ is a discrete Fourier or Walsh-Hadamard transform, and $\bm{\mathsf{S}}\in\mathbb{F}^{d\times m}$ is the transpose of a SparseCol matrix with exactly $\zeta$ nonzero entries per column as in the SparseCol matrix above.
These distributions yield randomized DRMs that work well in practice, with theoretical guarantees summarized in Table 1. Gaussian test matrices are oblivious subspace embeddings that enjoy a superior lower bound on the requisite embedding dimension $d$ to guarantee (21), but the costs of storage and matrix-vector products for Gaussians are much more expensive than for “structured” embeddings like SRTT or sparse sign matrices. Put simply, Gaussian sketching makes for straightforward analysis, but randomized sketching with structured random matrices makes for better performance. The reader is referred to (martinsson2020, Sections 8,9) for detailed discussion.
Randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ Dimension $d$ for (21) Storage cost Mat-vec cost
Gaussian $d=\Omega\left(\nicefrac{{\mbox{\small$k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $O(md)$ $O(md)$
SRTT $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $O(m\log m)$ $O(m\log d)$
Sparse sign $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right),\ \zeta=\Omega\left(\nicefrac{{\mbox{\small$\log k$}}}{{\mbox{\small$\varepsilon$}}}\right)$ $O(\zeta m\log d)$ $O(\zeta m)$
CountSketch $d=\Omega\left(\nicefrac{{\mbox{\small$k^{2}$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $O(m\log d)$ $O(m)$
SparseStack $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right),\ \zeta=\Omega\left(\nicefrac{{\mbox{\small$\log k$}}}{{\mbox{\small$\varepsilon$}}}\right)$ $O(d\zeta)$ $O(\zeta m)$
Table 1: Summary of coordinate-mixing randomized DRMs commonly used for (row-)sketching rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, which are also oblivious subspace embeddings (OSE). Theoretical lower bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy (21) are provided, as well as asymptotic complexities to store $\bm{\mathsf{\Gamma}}$ and to apply $\bm{\mathsf{\Gamma}}$ to a vector $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptm}$, cf. Camano2025_OSI; martinsson2020; woodruff2015.
While the embedding dimensions in Table 1 guarantee the oblivious subspace embedding (OSE) property (21), it has been observed that smaller values of $d$ are often sufficient for many linear algebra tasks in practice when utilizing any of the randomized DRMs of Table 1 (with the exception of CountSketch), e.g., $d=k+10$, cf. Sect. 4.1. Recently, this observation was thoroughly investigated from the perspective of oblivious subspace injections, versus embeddings, which relaxes the OSE property to the following weaker one.
A random matrix $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an oblivious subspace injection (OSI) of $\bm{\mathsf{A}}$ with injectivity $\alpha$ if it satisfies the injectivity property
$\alpha\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|\leq\|\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}\bm{\mathsf{x}}\|\ \ \forall\ \bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn},$ (25)
and the isotropy property that $\mathbb{E}[\|\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}\bm{\mathsf{x}}\|^{2}]=\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|^{2}$ for all $\bm{\mathsf{x}}\in\mathbb{F}^{n}$. In other words, an OSI satisfies the lower bound requirement in (21), taking $\alpha=1-\varepsilon$, so that it does not annihilate any vector in the column space of $\bm{\mathsf{A}}$. Arguably, the “injectivity” parameter $\alpha$ is more crucial for the performance of randomized sketching than the “dilation” parameter $(1+\varepsilon)$ in the OSE definition Camano2025_OSI. All of the random matrices defined above satisfy the isotropy property, and if we focus instead on the weaker injectivity property of OSIs, we can obtain theoretical guarantees on the performance of randomized sketching that more closely align with what is observed in practice. In Table 2, we report the sketching dimension $d$ required to satisfy the OSI property, compared to the OSE property.
With the OSI property guiding our choice of random sketching matrix, we enjoy more favorable sketching dimensions for the random matrices considered above. In the OSI framework, Gaussian random matrices are still the gold standard for approximation accuracy, but suffer the same performance drawbacks in terms of storage and matrix-vector product, or mat-vec, costs. Structured random matrices offer significantly better performance, and specifically, sparse random matrices exhibit the best empirical performance, as demonstrated by many practical investigations Chen2025_sparseGPU; ChenakkodDereZinskiDongRudelson2024_OptimalEmbeddingSparse; EpperlyMeierNakatsukasa2024_FastRandomizedLeastSquares; halko2011; KaneNelson2014; martinsson2020; nelson2013; Tropp2015_MatrixConcentration; woodruff2015, and now also supported by the theoretical guarantees of Camano2025_OSI; Tropp2025_OSI summarized in Table 2.
In general, sparse random matrices are recommended for sketching a dense unstructured input matrix, due to the reduced storage costs and significant acceleration afforded by high-performance sparse arithmetic libraries. To that end, sketching with SparseRTT matrices, or sparsified SRTTs, is more computationally efficient than with standard SRTTs or Gaussians. Compared to the SRTT, the SparseRTT matrix can reduce data movement and better facilitate parallelism at the hardware level. However, the SparseStack matrix is typically faster and thus recommended over SparseRTTs whenever sparse linear algebra packages are available.
As an additional boon, the OSI values of $d$ that are reported in Table 2 are even still too pessimistic for practical implementation purposes. In particular, the results of Camano2025_OSI provide compelling evidence that SparseStack matrices are the superior choice of DRM in practice; it is conjectured that SparseStack matrices can achieve the $\frac{1}{2}$-OSI property with $d=O(k)$ and constant sparsity $\zeta=O(1)$. This conjecture is supported numerically, with $d=2k$ and $\zeta=4$ used in the experiments of Camano2025_OSI to demonstrate that the low-rank approximations obtained from the SparseStack sketches are of comparable accuracy to those obtained from Gaussian sketches.
Finally, we note that for input matrices that arise from tensors (cf. Sect. 5.3), we can construct Khatri-Rao sketching matrices that exploit underlying Kronecker product structure Camano2025_OSI. As we show in Sect. 6.1, these sketching matrices are formed as Kronecker products of isotropic random vectors. Often, these random vectors are drawn from (real) Gaussian and Rademacher distributions, but it is advised in Camano2025_OSI that the complex spherical distribution be used to avoid near-zero injectivity $\alpha$, which can occur in worst-case scenarios. For our sketching matrix constructions in Sect. 6, we do not assume any specific isotropic distribution unless otherwise specified.
DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ Dimension $d$ for $\varepsilon$-OSE (21) Dimension $d$ for $1/2$-OSI (25)
Gaussian $d=\Omega\left(\nicefrac{{\mbox{\small$k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $d=O(k)$,
SRTT $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $d=O(k\log k)$
Sparse sign $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right),\ \zeta=\Omega\left(\nicefrac{{\mbox{\small$\log k$}}}{{\mbox{\small$\varepsilon$}}}\right)$ $d=O(k),\zeta=O(\mu(\bm{\mathsf{A}})\cdot\log k)$
CountSketch $d=\Omega\left(\nicefrac{{\mbox{\small$k^{2}$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $d=O(k^{2})$
SparseStack $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right),\ \zeta=\Omega\left(\nicefrac{{\mbox{\small$\log k$}}}{{\mbox{\small$\varepsilon$}}}\right)$ $d=O(k),\zeta=O(\log k)$
SparseRTT $d=O(k),\zeta=O(\log^{3}k)^{\dagger}$ $d=O(k),\zeta=O(\log k)$
Table 2: Summary of coordinate-mixing randomized DRMs commonly used for row-sketching rank-$k$ $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, which are also oblivious subspace injections (OSI). Theoretical bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy the OSI property (25) vs. the OSE property (21) are provided, cf. Camano2025_OSI; Tropp2025_OSI. For OSIs, we enforce a constant injectivity parameter, i.e. $\alpha\geq 1/2$, which is generally not attainable for OSEs due to the dilation parameter. For the OSE column of SparseRTTs, the dagger $\dagger$ signifies that a different sparse matrix is used for the OSE property, cf. (Camano2025_OSI, Section 4.1).
3.2 Randomized Matrix Sampling
In the previous section, we observed that a random sketch $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ (or $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$) is a new matrix whose rows (or columns) are random linear combinations of the rows (or columns) of $\bm{\mathsf{A}}$. However, we can also construct randomized DRMs that sample coordinates of $\bm{\mathsf{A}}$ according to a probability distribution. We use the separate term randomized sampling to refer to these types of maps.
To illustrate, suppose that we want to draw $d$ random i.i.d. row indices of a matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ according to probabilities $(p_{1},\ldots,p_{m})$. Common sampling methods include:
  • Uniform ((martinsson2020, Section 9.6.2), (tropp2011b, Lemma 3.4), Kannan2017):
    $p_{i}=1/m.$
  • Squared (row) norms (Frieze2004):
    $p_{i}=\|\bm{\mathsf{a}}_{i}\|_{2}^{2}/\|\bm{\mathsf{A}}\|^{2}_{\text{F}},$
    where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$,
  • Row leverage scores (DrineasKannanMahoney2006; DrineasMahoneyMuthukrishnan2006, (martinsson2020, Section 9.6.3)):
    $\begin{split}p_{i}=l_{i}/n\ \textup{ for leverage score }l_{i}=\|\bm{\mathsf{a}}_{i}\|^{2}_{(\bm{\mathsf{A}}^{*}\bm{\mathsf{A}})^{-1}},\end{split}$ (26)
    where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$ and the norm is defined as $\|\bm{\mathsf{v}}\|_{\bm{\mathsf{B}}}=\sqrt{\bm{\mathsf{v}}^{*}\bm{\mathsf{B}}\bm{\mathsf{v}}}$.
  • Determinantal point processes (DPP) ((DerezinskiMahoney2021, Section 5.2), kulesza2012determinantal): If $S\subseteq[m]$, $|S|=d$,
    $p_{S}=\frac{\det\left((\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}\right)}{\det\left(\bm{\mathsf{I}}+\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\right)},$ (27)
    where $(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}$ is the sub-matrix of $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\in\mathbb{F}^{m\times m}$ indexed by $S$ and all indices $S$ are selected at once. However, in practice, vectors (rows) corresponding to indices $S$ are selected iteratively, as in the algorithm presented in Hough_DPP_2005. Let $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}=\sum_{i}\lambda_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ denote the eigendecomposition with $\lambda_{i}\in[0,1]$ for all $i$. Let $s_{i}\sim\textup{Bernoulli}(\lambda_{i})$, and let $\bm{\mathsf{U}}$ be the matrix with columns $\bm{\mathsf{u}}_{i}$. By (Hough_DPP_2005, Theorem 7) and (DerezinskiMahoney2021, Theorem 12), we can use $\sum_{i}s_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ as a proxy for DPP sampling as follows, letting $\bm{\mathsf{v}}_{i}^{*}$ represent the $i$th row of $\bm{\mathsf{U}}$:
    1. 1.
      Sample vector $\bm{\mathsf{v}}_{i}$ with $p_{i}\propto\|\bm{\mathsf{v}}_{i}\|_{2}^{2}$
    2. 2.
      Project all vectors $\bm{\mathsf{v}}_{j}$ onto the subspace orthogonal to $\bm{\mathsf{v}}_{i}$
    3. 3.
      Update the probabilities and repeat d1d-1 times.
  • Nuclear norm maximization (Altschuler2016GreedyCSS; Fornace2024; Ordozgoiti2018IterativeCSS; Ordozgoiti2019RegularizedCSS): While neither a traditional sketching or sampling method, we include nuclear norm maximization in this list because of its comparable performance to DPP sampling. Let $\bm{\mathsf{K}}=\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$, $S\subseteq[m]$, $|S|=d$, and define
    $\mathcal{L}_{\bm{\mathsf{K}}}(S)=\textup{trace}\left(\bm{\mathsf{K}}_{:,S}(\bm{\mathsf{K}}_{S,S})^{-1}\bm{\mathsf{K}}_{S,:}\right)=\textup{trace}\left((\bm{\mathsf{K}}^{2})_{S,S}(\bm{\mathsf{K}}_{S,S})^{-1}\right).$
    Maximizing $\mathcal{L}_{\bm{\mathsf{K}}}(S)$ over all possible subsets of $[m]$ of size $d$ is infeasible as in the DPP setting, so we compute approximate nuclear scores using random sketches of $\bm{\mathsf{K}}$ and $\bm{\mathsf{A}}$, and the identities
    $\textup{diag}(\bm{\mathsf{K}}^{2})=\mathbb{E}_{\bm{\mathsf{x}}\sim\mathcal{N}(\bm{\mathsf{0}},\bm{\mathsf{I}})}[(\bm{\mathsf{Kx}})^{\odot 2}],\hskip 5.69054pt\textup{diag}(\bm{\mathsf{K}})=\mathbb{E}_{\bm{\mathsf{x}}\sim\mathcal{N}(\bm{\mathsf{0}},\bm{\mathsf{I}})}[(\bm{\mathsf{\bm{\mathsf{A}}\bm{\mathsf{x}}}})^{\odot 2}],$
    where $\odot 2$ denotes element-wise squaring. Namely, we compute the averages
    $\textup{diag}(\bm{\mathsf{K}}^{2})/\textup{diag}(\bm{\mathsf{K}})\approx\frac{(\bm{\mathsf{K}}\bm{\mathsf{\Omega}})^{\odot 2}\bm{\mathsf{1}}}{(\bm{\mathsf{A}}\bm{\mathsf{\Phi}})^{\odot 2}\bm{\mathsf{1}}}\Rightarrow\bm{\mathsf{K}}^{2}_{\ell,\ell}/\bm{\mathsf{K}}_{\ell,\ell}\approx\frac{\|[\bm{\mathsf{K}}\bm{\mathsf{\Omega}}]_{\ell,:}\|^{2}_{2}}{\|[\bm{\mathsf{A}}\bm{\mathsf{\Phi}}]_{\ell,:}\|^{2}_{2}},$
    where $\bm{\mathsf{\Omega}},\bm{\mathsf{\Phi}}$ are Gaussian matrices with $O(\log m)$ columns. We then update $\bm{\mathsf{K}}$ and $\bm{\mathsf{A}}$ by Schur complements with respect to the selected index, and repeat d1d-1 times.
We can express the action of randomized sampling as a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$. The lower bounds on the minimal dimension required for a random sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) are summarized in Table 3, along with the computational complexity of computing the associated probabilities. Unlike the bounds in Table 1, the values of $d$ required for a randomized sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) depend not only on the rank $k$ of $\bm{\mathsf{A}}$ and the distortion parameter $\varepsilon$, but also, in some cases, on quantities associated with the matrix $\bm{\mathsf{A}}$ that are usually very expensive to compute.
For uniform sampling, the requisite embedding dimension $d$ depends on the coherence $\mu$ of $\bm{\mathsf{A}}$, defined as $\mu(\bm{\mathsf{A}})=\max_{1\leq i\leq m}\frac{n}{m}l_{i}$, or the maximum leverage score $l_{i}$ in (26). The coherence is a measure of how evenly information in the matrix is distributed; high coherence indicates that its key information is very localized and that uniform sampling may perform poorly as a result; see, e.g., Ipsen14 or (martinsson2020, Section 9.6.2). In spite of this drawback, uniform sampling has become a popular choice for large-scale linear algebra computations as a cost-effective oblivious subspace embedding (OSE) method. However, it is unclear how to best choose the embedding dimension $d$ in practice since the coherence is typically unavailable, cf. (martinsson2020, Section 9.6.4).
Squared norm sampling and leverage score sampling also yield randomized subspace embeddings, but not oblivious ones, though we can still analyze the embedding dimensions required for them to satisfy (21). For squared norm sampling, the value of $d$ depends on the condition number $\kappa$ of $\bm{\mathsf{A}}^{*}\bm{\mathsf{A}}$, defined as the ratio of its largest to smallest eigenvalue, which is not readily available in practice. The embedding dimension for leverage score sampling only depends on $k$ and $\epsilon$, but this independence comes at great computational expense to compute the sampling probabilities. In particular, we note that leverage scores are as expensive to directly compute as $\texttt{svd}(\bm{\mathsf{A}})$, rendering their direct computations infeasible for most practical purposes. However, there exist fast methods to approximate them, cf. AlaouiMahoney15; Chen21; DrineasMagdonMahoneyWoodruff12; MuscoMusco17.
DPP sampling is known to have nearly optimal theoretical guarantees, cf. derezinski2020improved. However, naive implementations incur a cost of $O(m^{3})$, which makes DPP sampling practically infeasible. In derezinski2019exact, the DPP-VFX algorithm is presented, which, given access to $\bm{\mathsf{A}}\bm{\mathsf{A}}^{\top}$, samples exactly from a DPP distribution with a pre-processing cost of $O(mk^{10}+k^{15})$. In particular, the DPP-VFX algorithm was the first to achieve a cost of $O(m\cdot\textup{poly}(k))$ for the first sample and $O(\textup{poly}(k))$ for subsequent samples; however, there still exists a runtime bottleneck of $\Omega(m)$ due to the computation of $m$ marginals for the associated sampling distribution. The same authors improved on DPP-VFX with the $\alpha$-DPP algorithm of calandriello2020sampling, which avoids the $\Omega(m)$ bottleneck by incorporating uniform sub-sampling before computing marginals.
While nuclear norm maximization is a greedy method of choosing columns based on posterior error minimization, the authors of Fornace2024 present a “matrix-free” algorithm for fast implementation with randomization. Its performance is comparable to that of DPP sampling, and nuclear norms could be adapted into a sampling distribution, so we include it in this section for side-by-side comparisons in Table 3. Namely, in Fornace2024, randomized sketching is used to compute the approximate scores that dictate column selection. We note that the requisite embedding dimension $d$ for nuclear norm maximization depends on the quantity $\nu=\frac{\textup{trace}(\bm{\mathsf{K}})-\textup{trace}^{(k)}(\bm{\mathsf{K}})}{\textup{trace}^{(k)}(\bm{\mathsf{K}})}$, where $\bm{\mathsf{K}}=\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$ and $\textup{trace}^{(k)}(\bm{\mathsf{K}})$ is the sum of its $k$ largest eigenvalues. For the complexity, we also include the dependence on the quantity $\phi$, which is the maximum cost of a mat-vec with $\bm{\mathsf{K}}$ or $\bm{\mathsf{A}}$, cf. (Fornace2024, Corollary 4.1).
Sampling distribution Dimension $d$ Complexity of distribution
Uniform $d=\Omega\left(\nicefrac{{\mbox{\small$\mu\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $O(1)$
Squared norms $d=\Omega\left(\nicefrac{{\mbox{\small$\kappa\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $O(mn)$
Leverage scores $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right)$ $O(mn^{2})$
[…]
Nuclear norm max $d=\Omega(k/\varepsilon+k-1)(\log(\nu^{-1})+\log(1/\varepsilon)))$ $O(md^{2}\log m+d\log m\psi)$
Table 3: Summary of randomized sampling distributions commonly used to sample rows of rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$. Theoretical lower bounds on the embedding dimension $d$ are given for $\bm{\mathsf{\Gamma}}$ to satisfy (21), as well as the asymptotic complexity of computing the associated probability distribution, cf. DerezinskiMahoney2021; martinsson2020. We note (by $\dagger$) that the cost of DPP sampling can be reduced, cf. (DerezinskiMahoney2021, Section 5), derezinski2019exact, calandriello2020sampling.
3.3 Application: Randomized DRMs for Overdetermined Least Squares
As an illustrative application of randomized DRMs that will also be useful for our tensor algorithms, we consider the overdetermined least squares problem, in which we solve
$\bm{\mathsf{x}}_{*}=\arg\min_{\bm{\mathsf{x}}\in\mathbb{R}^{n}}\|\bm{\mathsf{A}}\bm{\mathsf{x}}-\bm{\mathsf{b}}\|_{2},$ (28)
for given $\bm{\mathsf{A}}\in^{m\times n}$, $\bm{\mathsf{b}}\in\mathbb{R}^{m}$, and $m\gg n$. Using CPQR for (28) requires $O(mn^{2})$ operations, which is undesirable for large $m$.
A natural idea is to apply a randomized DRM to $\bm{\mathsf{A}}$ and $\bm{\mathsf{b}}$ in (28), a strategy known as “sketch-and-solve” sarlos2006improved. Namely, we draw a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ and solve a smaller sketched problem
$\widehat{\bm{\mathsf{x}}}=\arg\min_{\bm{\mathsf{x}}\in\mathbb{R}^{n}}\|\bm{\mathsf{\Gamma}}(\bm{\mathsf{A}}\bm{\mathsf{x}}-\bm{\mathsf{b}})\|_{2}.$ (29)
If $\bm{\mathsf{\Gamma}}$ is a randomized subspace embedding for the $k$-dimensional column space of $\begin{bmatrix}\bm{\mathsf{A}}&\bm{\mathsf{b}}\end{bmatrix}$ with distortion $\varepsilon\in(0,1)$, then
$(1-\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}_{*}-\bm{\mathsf{b}}\|_{2}\leq\|\bm{\mathsf{A}}\widehat{\bm{\mathsf{x}}}-\bm{\mathsf{b}}\|_{2}\leq(1+\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}_{*}-\bm{\mathsf{b}}\|_{2}$
when $d\approx k/\varepsilon^{2}$. Solving (29) requires just $O(dn^{2})$ operations, and often in practice, there are ways to avoid forming $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ and $\bm{\mathsf{\Gamma}}\bm{\mathsf{b}}$ explicitly.
Randomized DRMs are also utilized in other methods of solving (28). Sketched GMRES (NakatsukasaTropp2024, Algorithm 1.1) includes a $k$-truncated Arnoldi process to quickly compute an approximate basis of the Krylov subspace. Sketch-and-precondition methods are also commonly used, but we note that these methods were recently shown to be numerically unstable when initialized with the zero vector, cf. MeierNakatsukasaTownsendWebb2024. With iterative sketching methods, sketched least squares problems are solved until converged, and these methods were recently proven to be forward stable in Epperly2024. We discuss alternative randomized DRMs for overdetermined least squares problems for computing tensor decompositions in Sect. 6, which are based on the randomized algorithms for low-rank matrix decompositions that we discuss next.
4 Randomized Algorithms for Low-Rank Matrix Decompositions
In this section, we illustrate several ways that randomized DRMs are often used to compute low-rank matrix decompositions. We first introduce the rangefinder problem and show how it can be solved with randomized DRMs in Sect. 4.1. We then explain how randomized DRMS can be used to estimate the error in the corresponding low-rank approximation, which can be utilized for adaptive rangefinding. These ideas are foundational for randomized algorithms that have become ubiquitous for fast and reliable low-rank matrix approximation: the randomized SVD (Sect. 4.2) and randomized ID/CUR (Sect. 4.3).
4.1 The Randomized Rangefinder
At the heart of many linear algebraic tasks lies the rangefinder problem, the computation of a lower-dimensional space that sufficiently captures the action of a given matrix. More specifically, to solve the rangefinder problem for an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and target rank $k\leq\min(m,n)$, we compute an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns approximately span the column space of $\bm{\mathsf{A}}$. We then an obtain a rank-$k$ approximation given by $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$, with approximation error
$\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|=\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|,$ (30)
measured in either the Frobenius or spectral norm in our work. The Frobenius norm is often preferable for ease of analysis, but error bounds in the Frobenius norm can be less informative than bounds in the spectral norm, cf. (martinsson2020, Remark 2.1).
The rangefinder problem can be handled effectively with randomized sketching as shown in Algorithm 1. This algorithm produces an orthonormal matrix $\bm{\mathsf{Q}}$ via QR factorization of an $m\times\ell$ random sketch of $\bm{\mathsf{A}}$ for desired embedding dimension $\ell$. The total cost of Algorithm 1 includes the cost of simulating the $n\times\ell$ random test matrix, $O(mn\ell)$ operations to form the sketch, and $O(mk^{2})$ operations for the QR factorization. Sparsity or availability of fast matrix-vector primitives with $\bm{\mathsf{A}}$ can lend additional performance improvements.
Algorithm 1 Randomized Rangefinder
[…]
2:Orthonormal $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ such that $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$
3:Draw randomized DRM $\bm{\mathsf{\Omega}}^{*}\in\mathbb{F}^{\hskip 0.56905pt\ell\times n}$ for $k<\ell\leq\min(m,n)$
4:Form sketch $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$.
5:Compute $\bm{\mathsf{Q}}=\texttt{col}(\bm{\mathsf{Y}},k)$.
Implementation details
There are many considerations to be made when implementing the randomized rangefinder algorithm in practice. We summarize the important details here and refer to halko2011; martinsson2020 for more thorough treatments.
First, sketching, as opposed to sampling, should be used whenever possible. Gaussian matrices are popular for easy implementation and strong theoretical results, but structured random matrices (e.g., SparseStack) perform almost as well in practice. Second, the numerical accuracy of the rangefinder algorithm can be improved with power iteration Hestenes1952methods; IpsenWills2006; Saad1986GMRES or Krylov methods IpsenMeyer1998; Saad2022; SimonciniSzyld2003, in addition to numerically stable orthogonalization techniques like Householder reflections or rank-revealing QR Chan1987; Ilse1994.
One last important detail concerns the subspace dimension. If the target rank $k$ of the low-rank approximation to $\bm{\mathsf{A}}$ is known a priori, then a randomized DRM of size $\ell=k+p$ is usually sufficient with a small amount of oversampling $p$, such as $p=5$. (We deal with the situation where $k$ is unknown in Sect. 4.1.) For Gaussian test matrices with slight oversampling, (halko2011, Theorem 10.1) guarantees that the expected error of the randomized rangefinder is comparable to the best rank-$k$ approximation error from the truncated SVD of $\bm{\mathsf{A}}$, so long as its trailing singular values are small. This is generally achievable, for instance, by powering $\bm{\mathsf{A}}$ in the sketch $\bm{\mathsf{Y}}=(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})^{q}\bm{\mathsf{\Omega}}$ some small number of iterations $q$ RokhlinSzlamTygert2009.
The theoretical results for Gaussian random matrices often furnish the performance heuristics for structured random sketching matrices, which behave similarly to Gaussians in practice despite the more pessimistic bounds that appear in their analysis (halko2011, Theorem 11.2). This claim is strengthened by the simple addition of a posteriori “certificates of accuracy” upon termination of Algorithm 1, as explained in the next section.
Error estimation and certificates of accuracy
We now demonstrate how the approximation error (30) can be estimated reliably and efficiently with randomized sketching. This method of error estimation also serves as a “certificate of accuracy” for rangefinding, an assurance that the orthornomal basis given by $\bm{\mathsf{Q}}$ produces a satisfactory low-rank approximation of $\bm{\mathsf{A}}$.
Suppose that we have an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns potentially form a basis for the range of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, e.g., the output of Algorithm 1 for a given $k$. We estimate the approximation error $\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|$ using the following method.
Let $\bm{\mathsf{\Phi}}\in\mathbb{F}^{\hskip 0.56905ptn\times s}$ be a Gaussian matrix that is independent from any random matrix used to compute $\bm{\mathsf{Q}}$. In practice, $s$ is a small fixed value, such as $s=10$. We form a new random sketch, known as the auxiliary sample
$\bm{\mathsf{Z}}=\bm{\mathsf{A}}\bm{\mathsf{\Phi}},$ (31)
incurring only $s$ additional matrix-vector products with $\bm{\mathsf{A}}$. Using $\bm{\mathsf{Z}}$, we can obtain an inexpensive Gaussian sketch of the approximation error (30):
$(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{Z}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\bm{\mathsf{\Phi}}.$ (32)
We then take advantage of randomized methods to estimate the norm of (32), such as randomized trace estimation for the Frobenius norm or randomized Krylov methods for the Schatten $2p$-norm (a proxy for the the spectral norm with sufficiently large $p$). Recall that the Schatten $p$-norm of a matrix is the $\ell_{p}$-norm of its singular values, so that the Schatten $\infty$-norm coincides with the spectral norm and the Schatten 2-norm with the Frobenius norm.
To estimate the error in the Frobenius norm, we use the approximation
$\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|_{\text{F}}^{2}\approx\frac{1}{s}\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{Z}}\|_{\text{F}}^{2},$ (33)
relying on the theory of randomized trace estimation (recall that $\textup{trace}(\bm{\mathsf{X}}^{*}\bm{\mathsf{X}})=\|\bm{\mathsf{X}}\|_{\text{F}}^{2}$). Namely, we exploit the fact that for any matrix $\bm{\mathsf{B}}$ and isotropic random vector $\omega$, $\text{trace}(\bm{\mathsf{B}})=\bm{\mathsf{\omega}}^{*}\bm{\mathsf{B}}\bm{\mathsf{\omega}}$ in expectation. To reduce the variance, we take the average of $s$ random variables in (33), using the auxiliary sample $\bm{\mathsf{Z}}$.
Estimating the error in the spectral norm is not quite as simple. It is possible to construct unbiased estimators for the Schatten 2pp-norm for any positive integer $p$, as for the Frobenius norm above, but these estimators are expensive with large variance KongValiant2017; LiWoodruff2014. Randomized Krylov methods are often used instead to estimate the spectral norm LeykWoz1998; MartinssonRokhlinTygert2011; MuscoMusco2015; Tropp2022; WangZhangZhang2015. For example, we can estimate the spectral norm error with a randomized block Krylov method, letting $\bm{\mathsf{B}}=\begin{bmatrix}(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\end{bmatrix}\begin{bmatrix}(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\end{bmatrix}^{*}$,
$\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|^{2}_{2}\approx\max_{\bm{\mathsf{v}}\in\mathcal{K}(\bm{\mathsf{B}};\bm{\mathsf{\Phi}})}\frac{\bm{\mathsf{v}}^{*}\bm{\mathsf{B}}\bm{\mathsf{v}}}{\bm{\mathsf{v}}^{*}\bm{\mathsf{v}}},$ (34)
where $\mathcal{K}(\bm{\mathsf{B}};\bm{\mathsf{\Phi}})=\textup{span}\begin{bmatrix}\bm{\mathsf{\Phi}}\ &\bm{\mathsf{B}}\bm{\mathsf{\Phi}}\ &\bm{\mathsf{B}}^{2}\bm{\mathsf{\Phi}}\ &\cdots&\bm{\mathsf{B}}^{q}\bm{\mathsf{\Phi}}\end{bmatrix}$ is a block Krylov subspace of depth $q$. In general, $q\approx\log m$ is necessary to achieve the fast convergence rate of $1/q^{2}$ for Krylov methods, but smaller values of $q$ can be used in practice for matrices with fast spectral decay. Although Krylov methods enjoy quick convergence, numerical instability is a known shortcoming, so careful implementation is needed, cf. MuscoMusco2015; Tropp2018; WangZhangZhang2015.
Regardless of the choice of norm, an auxiliary sketch of the error with a Gaussian random test matrix $\bm{\mathsf{\Phi}}$ can provide us with an a posteriori “certificate of accuracy” for rangefinding (martinsson2020, Section 12.2). For instance, if a structured random matrix (e.g., SRTT, sparse sign) is used in Algorithm 1 instead of a Gaussian matrix, the computations are generally faster but lack formal guarantees on the accuracy of the computed solution. More generally, if we are unsure of the accuracy of a given solution to the rangefinder problem, we can take a small (independent) Gaussian sketch of the error as in (31). This auxiliary sketch is inexpensive and avails us of the powerful theory of Gaussian matrices, regardless of how the column space approximation was obtained. In the next section, we demonstrate how this method of error estimation can be utilized for adaptive rangefinding when the target rank is unknown a priori.
Adaptive rangefinding
Up to this point, we have assumed that the target approximation rank could be passed as an argument to the randomized rangefinder algorithm. Unfortunately, a satisfactory target rank for the approximation is frequently unknown a priori. In that case, we can implement the randomized rangefinder algorithm adaptively, using the error estimation techniques of the previous section to approximate the range of $\bm{\mathsf{A}}$ to a desired precision $\varepsilon$, forming $\bm{\mathsf{Q}}$ incrementally until that precision is achieved.
For ease of notation, let norm_est($\bm{\mathsf{Z}}$) denote the rangefinder error estimate obtained by one of the techniques in Sect. 4.1, e.g., (33). An adaptive version of the randomized rangefinder is given in Algorithm 2, which we quickly summarize.
Algorithm 2 Adaptive Randomized Rangefinder
[…]
2:Orthonormal $\bm{\mathsf{Q}}$ such that $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|\leq\varepsilon$ with high probability
3:$i=1$
4:Draw Gaussian random matrix $\bm{\mathsf{\Omega}}\in\mathbb{F}^{\hskip 0.56905ptn\times b}$
[…]
7:while $\texttt{norm\_est}(\bm{\mathsf{Y}})>\varepsilon$ do $\triangleright$ Error estimation of Sect. 4.1
8:  $i=i+1$
9:  Draw independent Gaussian $\bm{\mathsf{\Omega}}\in\mathbb{F}^{\hskip 0.56905ptn\times b}$
10:  $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$
11:  $\bm{\mathsf{Y}}=\bm{\mathsf{Y}}-\sum_{j=1}^{i-1}\bm{\mathsf{Q}}_{j}(\bm{\mathsf{Q}}_{j}^{*}\bm{\mathsf{Y}})$
12:  $\bm{\mathsf{Q}}_{i}=\texttt{col}(\bm{\mathsf{Y}})$
13:end while
[…]
Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and let $b$ be the number of columns to be processed at a time. The choice of block size $b$ affects the computational efficiency of the algorithm. If $b$ is too small, we are not fully reaping the benefits of blocked matrix operations that help reduce memory traffic and avoid communication bottlenecks Heavner2022; Orti1998. With a larger block size $b$, more floating-point operations can be cast as matrix-matrix multiplication (BLAS-3) vs. matrix-vector multiplication (BLAS-2) Blackford2002; Dongarra1990; Gallivan1987; Higham1990. However, a larger block size runs the risk of overshooting the target rank, meaning we have done more work than necessary for rangefinding. In practice, a block size $10\leq b\leq 100$ is usually appropriate.
In each iteration of the while-loop in Algorithm 2, the matrix $\bm{\mathsf{Y}}$ holds a sample of the residual error after the execution of line 9:
$\bm{\mathsf{Y}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\bm{\mathsf{\Omega}},$ (35)
where $\bm{\mathsf{Q}}$ is the basis that has been built up to the current iteration. Critically, the random test matrix $\bm{\mathsf{\Omega}}$ used in (35) is independent from the random matrices used to compute the current basis $\bm{\mathsf{Q}}$. We note that to compute the very first residual error norm_est($\bm{\mathsf{Y}}$) in line 5, we can draw a small auxiliary sample that then can be recycled in the second batch of samples in line 8.
Variations of Algorithm 2 have been designed to optimize performance for specific types of input matrices. If the input matrix $\bm{\mathsf{A}}$ is small enough to store in RAM, it can be explicitly updated in each iteration via $\bm{\mathsf{A}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}$ as in (martinsson2020, Algorithm 14). Alternatively, the version of YuGuLiLiuLi2017 effectively handles input matrices stored out-of-core, as well as sparse input matrices. If the spectrum of the input matrix decays slowly, power iteration can also be incorporated into Algorithm 2 to improve performance MartinssonVoronin2016.
4.2 Randomized SVD
Now equipped with the randomized rangefinder, we present a foundational algorithm for low-rank matrix approximation known as the randomized SVD halko2011. Pseudocode is provided in Algorithm 3. We quickly go through its steps to show that the error of the resulting approximate SVD is precisely the randomized rangefinder error.
Algorithm 3 Randomized SVD
[…]
2:Orthonormal matrices $\bm{\mathsf{U}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$, $\bm{\mathsf{V}}\in\mathbb{F}^{\hskip 0.56905ptn\times k}$, and diagonal $\bm{\mathsf{\Sigma}}\in\mathbb{F}^{\hskip 0.56905ptk\times k}$ such that $\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$
3:Compute $\bm{\mathsf{Q}}=\texttt{RandomizedRangefinder}(\bm{\mathsf{A}},k)$ $\triangleright$ Algorithm 1
4:Form $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptk\times n}$
5:Compute $\widehat{\bm{\mathsf{U}}},\bm{\mathsf{\Sigma}},\bm{\mathsf{V}}]=\texttt{svd}(\bm{\mathsf{B}})$
6:Form $\bm{\mathsf{U}}=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$
Given an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, our goal is to compute an approximate rank-$k$ truncated singular value decomposition
$\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ (36)
where $\bm{\mathsf{U}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ and $\bm{\mathsf{V}}\in\mathbb{F}^{\hskip 0.56905ptn\times k}$ are orthonormal matrices, and $\bm{\mathsf{\Sigma}}\in\mathbb{F}^{\hskip 0.56905ptk\times k}$ is a diagonal matrix of entries approximating the largest singular values of $\bm{\mathsf{A}}$.
Suppose that $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ is a solution to the rangefinder problem (not necessarily from Algorithm 1), which corresponds to a rank-$k$ approximation of $\bm{\mathsf{A}}$ given by $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$. Let $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptk\times n}$ and compute singular value decomposition
$\bm{\mathsf{B}}=\underset{k\times k}{\widehat{\bm{\mathsf{U}}}}\ \underset{k\times k}{\bm{\mathsf{\Sigma}}}\ \underset{k\times n}{\bm{\mathsf{V}}^{*}}.$ (37)
We observe that
$\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ $=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{B}}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|,$ (38)
for $\bm{\mathsf{U}}:=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$; i.e. the approximation error of $\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$ is exactly the approximation error of $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$.
Thus, if $\bm{\mathsf{Q}}=\texttt{RandomizedRangefinder}(\bm{\mathsf{A}},k)$ is the rangefinder solution computed by Algorithm 1, the accuracy of the randomized SVD is equal to that of the randomized rangefinder. In particular, this means that the error estimation techniques of Sect. 4.1 can also be used to estimate the error of the randomized SVD. Then if the target rank $k$ is unknown a priori, the adaptive randomized rangefinder (Algorithm 2) can be carried out first, which determines $k$ in the randomized SVD.
The total cost of Algorithm 3 includes the cost of the randomized rangefinder, plus $O(mnk)$ floating-point operations for $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$ in line 2. We note that the cost of the SVD in line 3 is just $O(k^{2}n)$, so the cost of Algorithm 3 after rangefinding is usually dominated by the matrix-matrix multiplication. Executing Algorithm 3 requires storage for $O((m+n)k)$ numbers.
Recent work has also been done on cost-effective a posteriori error estimation methods that use only the information at hand for computing the low-rank decomposition. In EpperlyTropp2024, an algorithm for a posteriori error estimation is presented for the randomized SVD in a matrix-free setting, where the input matrix $\bm{\mathsf{A}}$ is only accessible via mat-vecs. Given the rank-$k$ randomized SVD $\bm{\mathsf{X}}=\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\in\mathbb{F}^{m\times n}$ corresponding to Gaussian random matrix $\bm{\mathsf{\Omega}}=\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\bm{\mathsf{\omega}}_{2}&\ldots&\bm{\mathsf{\omega}}_{k}\end{bmatrix}$, we want to estimate the mean-square error
$\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})=\mathbb{E}\|\bm{\mathsf{A}}-\bm{\mathsf{X}}_{k-1}\|^{2}_{\text{F}},$
where $\bm{\mathsf{X}}_{k-1}\in\mathbb{F}^{m\times(k-1)}$ is a rank-(k1)(k-1) randomized SVD.
By (EpperlyTropp2024, Theorem 2.1), the following is an unbiased estimator for $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})$, which we use as a proxy for the mean-square error of the rank-$k$ approximation $\bm{\mathsf{X}}$, called the leave-one-out error estimator:
$\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})=\mathbb{E}\left[\frac{1}{k}\sum_{j=1}^{k}\|(\bm{\mathsf{A}}-\bm{\mathsf{X}}^{(j)})\bm{\mathsf{\omega}}_{j}\|^{2}\right],$ (39)
where $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{Q}}^{(j)}(\bm{\mathsf{Q}}^{(j)})^{*}\bm{\mathsf{A}}$ is called a replicate, formed from the matrix $\bm{\mathsf{Q}}^{(j)}$ outputted by the randomized rangefinder using $\bm{\mathsf{\Omega}}$ without its $j$-th column $\bm{\mathsf{\omega}}_{j}$. Implemented naively, the right-hand side of (39) is too expensive for practical use, as it would require $k$ additional randomized rangefinder applications. However, by (EpperlyTropp2024, Equation 4.2),
$\bm{\mathsf{Q}}^{(j)}(\bm{\mathsf{Q}}^{(j)})^{*}=\bm{\mathsf{Q}}(\bm{\mathsf{I}}-\bm{\mathsf{t}}_{j}\bm{\mathsf{t}}_{j}^{*})\bm{\mathsf{Q}}^{*},$ (40)
where $\bm{\mathsf{t}}_{j}$, $j=1,\ldots,k$, are the normalized columns of $(\bm{\mathsf{R}}^{*})^{-1}$ in the QR factorization defined by $\bm{\mathsf{Q}}$ in step 1 of the randomized SVD (now also storing the $\bm{\mathsf{R}}$ matrix). Then we can obtain each replicate $\bm{\mathsf{X}}^{(j)}$ inexpensively as
$\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{U}}(\bm{\mathsf{I}}-\widehat{\bm{\mathsf{U}}}^{*}\bm{\mathsf{t}}_{j}\bm{\mathsf{t}}_{j}^{*}\widehat{\bm{\mathsf{U}}})\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ (41)
which allows us to compute (39) without any additional mat-vecs or applications of Algorithm 3, via
$\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})\approx\frac{1}{k}\sum_{j=1}^{k}\left\|[(\bm{\mathsf{R}}^{*})^{-1}](:,j)\right\|^{-2}.$
4.3 Randomized ID/CUR
In both the randomized rangefinder and the randomized SVD, orthonormal matrices are computed whose columns form an approximate basis for the row or column space of an input matrix $\bm{\mathsf{A}}$. The entries of these orthonormal matrices are linear combinations of the entries of $\bm{\mathsf{A}}$. However, for some problems (e.g., if $\bm{\mathsf{A}}$ is sparse or non-negative), it may be preferable to have a “natural basis,” consisting of subsets of rows or columns of $\bm{\mathsf{A}}$, as in the ID/CUR factorizations of Sect. 2.4. In this section, we provide an overview of randomized methods to compute approximate ID/CUR factorizations based on the randomized rangefinder, including recent work to improve skeleton selection and performance DongRBRP2025; dong2022; Pearce2025; Pritchard2025.
With the randomized rangefinder
Recall from Sect. 2.4 that pivoting algorithms, such as column-pivoted QR (CPQR) or LU with partial pivoting, are frequently used in the deterministic setting to select skeletons for an ID/CUR factorization. CPQR produces skeletons that are close to optimal in practice Cheng2005; martinsson2006interpolation. In particular, this is true when CPQR is applied to a random sketch of $\bm{\mathsf{A}}$ with appropriate embedding dimension. It is also conjectured in (trefethen1990, Remark 4) and exhibited empirically in dong2022; Pearce2025 that randomized sketching equips LUPP with near-optimal skeleton selection capability as well, even in adversarial cases. We illustrate how randomized sketching with CPQR or LUPP can be used to compute a row ID, then discuss alternate skeleton selection strategies that are rank-adaptive or based on sampling instead of sketching.
More specifically, let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be an input matrix of exact rank $k$ for notational convenience. Let $\bm{\mathsf{Y}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ have columns that span the range of $\bm{\mathsf{A}}$, e.g., the output of the rangefinder algorithm. By definition,
$\bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{Y}}}\ \underset{k\times n}{\bm{\mathsf{F}}}$
for some matrix $\bm{\mathsf{F}}$. Compute a row ID of $\bm{\mathsf{Y}}$, either using CPQR via $\sim,\sim,I]=\texttt{qr}(\bm{\mathsf{Y}}^{*},k)$ as in Sect. 2.4 or LUPP via $\sim,\sim,I]=\texttt{lu}(\bm{\mathsf{Y}},k)$ as in Sect. 2.4, to obtain
$\bm{\mathsf{Y}}=\underset{m\times k}{\bm{\mathsf{X}}}\underset{k\times k}{\bm{\mathsf{Y}}(I,:)}.$ (42)
We now observe that index set $I$ and row interpolation matrix $\bm{\mathsf{X}}$ form a row ID of $\bm{\mathsf{A}}$:
$\bm{\mathsf{A}}=\bm{\mathsf{Y}}\bm{\mathsf{F}}=\left[\bm{\mathsf{X}}\bm{\mathsf{Y}}(I,:)\right]\bm{\mathsf{F}}=\bm{\mathsf{X}}\left[\bm{\mathsf{Y}}(I,:)\bm{\mathsf{F}}\right]=\bm{\mathsf{X}}\bm{\mathsf{A}}(I,:).$
Thus, to compute a row ID of $\bm{\mathsf{A}}$, it suffices to compute a row ID of a matrix $\bm{\mathsf{Y}}$ whose columns span the range of $\bm{\mathsf{A}}$.
We can construct such a matrix $\bm{\mathsf{Y}}\in\mathbb{F}^{m\times\ell}$ for a rank-$k$ approximation of $\bm{\mathsf{A}}$ using the randomized rangefinder with high probability with, e.g., $\ell=k+10.$ Additionally, since we do not need $\bm{\mathsf{Y}}$ to be an orthonormal matrix here, we can save on computational cost by observing that the range of $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is itself an accurate approximation to the range of $\bm{\mathsf{A}}$ with high probability (halko2011, Theorems 10.5 and 10.6). The randomized rangefinder error $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ using $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is equal to $\|\bm{\mathsf{A}}-\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}\bm{\mathsf{A}}\|$, where $\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}$ is the orthogonal projector onto the column space of $\bm{\mathsf{Y}}$ or, equivalently, the column space of $\bm{\mathsf{Q}}$. Pseudocode for this randomized row ID algorithm (with sketched CPQR) is provided in Algorithm 4.
Algorithm 4 Randomized Row ID with Sketched CPQR
[…]
2:Row indices $I$ and row interpolation matrix $\bm{\mathsf{X}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ such that $\bm{\mathsf{A}}\approx\bm{\mathsf{X}}\bm{\mathsf{A}}(I,:)$
3:Draw randomized embedding $\bm{\mathsf{\Omega}}^{*}\in\mathbb{F}^{\hskip 0.56905pt(k+p)\times n}$
4:Form $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}\in\mathbb{F}^{\hskip 0.56905ptm\times(k+p)}$
5:Compute $\sim,\bm{\mathsf{R}},P]=\texttt{qr}(\bm{\mathsf{Y}}^{*})$
6:Set $I=P(1:k)$
7:Set $\bm{\mathsf{X}}(P,:)=\begin{bmatrix}\bm{\mathsf{I}}\ &\ \ \bm{\mathsf{R}}_{1:k,1:k}^{-1}\bm{\mathsf{R}}_{1:k,k+1:m}\end{bmatrix}^{*}$
Adaptive algorithms for randomized ID/CUR
Frequently in applications, the target rank for a low-rank approximation is unknown a priori, in which case we turn to rank-adaptive methods. If using CPQR for skeleton selection, the adaptive version of the randomized rangefinder may be used to build an orthonormal basis incrementally from independent random sketches or samples.
However, if we are using LUPP as the skeleton selection method for improved efficiency, we want to estimate the approximation error without computing an orthonormal basis. We briefly illustrate how a row ID can be constructed with the rank-adaptive, LUPP-based method of Pearce2025. Given an error tolerance $\tau$ and input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{m\times n}$, after $t$ iterations of this algorithm, we obtain target rank $k_{t}$, row skeletons $I_{s}^{(t)}$ with $|I_{s}^{(t)}|=k_{t}$, and interpolation matrix $\bm{\mathsf{W}}^{(t)}$ such that
$\|\bm{\mathsf{A}}-\bm{\mathsf{W}}^{(t)}\bm{\mathsf{R}}^{(t)}\|_{\text{F}}<\tau,$
where $\bm{\mathsf{R}}^{(t)}=\bm{\mathsf{A}}(I_{s}^{(t)},:)$.
Fix a block size $b$ (e.g., $b=10$), draw a random matrix $\bm{\mathsf{\Omega}}^{(0)}\in\mathbb{F}^{n\times b}$, and form $\bm{\mathsf{Y}}^{(0)}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}^{(0)}$. We use LUPP to compute
$\bm{\mathsf{P}}^{(0)}\bm{\mathsf{Y}}^{(0)}=\bm{\mathsf{L}}^{(0)}\bm{\mathsf{U}}^{(0)},$
so that the first $b$ skeletons $I_{s}^{(0)}$ are the first $b$ indices of $\bm{\mathsf{P}}^{(0)}[1\ldots m]^{*}$. To estimate $\|\bm{\mathsf{A}}-\bm{\mathsf{W}}^{(0)}\bm{\mathsf{R}}^{(0)}\|_{\text{F}}$, we draw an independent random matrix $\bm{\mathsf{\Omega}}^{(1)}\in\mathbb{F}^{n\times b}$ from the same distribution as $\bm{\mathsf{\Omega}}^{(0)}$ and form $\bm{\mathsf{Y}}^{(1)}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}^{(1)}$. By (Pearce2025, Equation 20),
$\|\bm{\mathsf{A}}-\bm{\mathsf{W}}^{(0)}\bm{\mathsf{R}}^{(0)}\|_{\text{F}}$ $=\|(\bm{\mathsf{P}}^{(0)}\bm{\mathsf{A}})(b+1:m,:)-\bm{\mathsf{L}}_{2}^{(0)}(\bm{\mathsf{L}}_{1}^{(0)})^{-1}(\bm{\mathsf{P}}^{(0)}\bm{\mathsf{A}})(1:b,:)\|_{\text{F}}$
$\approx\|\bm{\mathsf{A}}\bm{\mathsf{\Omega}}^{(1)}-\bm{\mathsf{W}}^{(0)}\bm{\mathsf{R}}^{(0)}\bm{\mathsf{\Omega}}^{(1)}\|_{\text{F}}$
$=\|\bm{\mathsf{P}}^{(0)}\bm{\mathsf{Y}}^{(1)}(b+1:m,:)-\bm{\mathsf{L}}_{2}^{(0)}(\bm{\mathsf{L}}_{1}^{(0)})^{-1}(\bm{\mathsf{P}}^{(0)}\bm{\mathsf{Y}}^{(1)})(1:b,:)\|_{\text{F}},$
where $\bm{\mathsf{L}}^{(0)}_{1}=\bm{\mathsf{L}}^{(0)}(1:b,:)$ and $\bm{\mathsf{L}}^{(0)}_{2}=\bm{\mathsf{L}}^{(0)}(b+1:m,:)$. If the given tolerance is not met, we LU-factorize $\bm{\mathsf{Y}}^{(0)}\ \bm{\mathsf{Y}}^{(1)}]$ re-using the previous factorization, and repeat. In general, the error after t1t-1 iterations is given by the Schur complement
$\|\bm{\mathsf{S}}^{(t)}\|_{\text{F}}=\|(\bm{\mathsf{P}}^{(t-1)}\bm{\mathsf{Y}}^{(t)})(tb+1:m,:)-\bm{\mathsf{L}}^{(t-1)}_{2}(\bm{\mathsf{L}}_{1}^{(t-1)})^{-1}(\bm{\mathsf{P}}^{(t-1)}\bm{\mathsf{Y}}^{(t)})(1:tb,:)\|_{\text{F}}.$
A similar idea is presented in Pritchard2025 for a rank-adaptive CUR factorization, using a single small sketch of the residual matrix $\bm{\mathsf{S}}=\bm{\mathsf{A}}-\bm{\mathsf{C}}\bm{\mathsf{U}}\bm{\mathsf{R}}$. (Here, $\bm{\mathsf{U}}$ is computed as $\bm{\mathsf{A}}(I,J)^{\dagger}$ vs. $\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$ as in Sect. 2.4.) The column residual is initialized as a row sketch, $\bm{\mathsf{S}}^{(C)}_{0}=\bm{\mathsf{G}}\bm{\mathsf{A}}\in\mathbb{F}^{1.1b\times n}$, where typically $5\leq b\leq 250$, and the row and column skeletons are empty sets. While the error criterion is unmet, column skeletons $J$ are first computed from $\bm{\mathsf{S}}^{(C)}_{0}$, and the row residual is formed as $\bm{\mathsf{S}}^{(R)}_{1}=\bm{\mathsf{A}}(:,J)-\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}(:,J)=\bm{\mathsf{A}}(:,J)$. Row skeletons $I$ are then computed from $\bm{\mathsf{S}}^{(R)}_{1}$, and the column residual is formed as $\bm{\mathsf{S}}^{(C)}_{1}=\bm{\mathsf{G}}\bm{\mathsf{A}}-\bm{\mathsf{G}}\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}$, where now $\bm{\mathsf{C}}_{1}=\bm{\mathsf{A}}(:,J)$, $\bm{\mathsf{R}}_{1}=\bm{\mathsf{A}}(I,:)$, and $\bm{\mathsf{U}}_{1}=\bm{\mathsf{A}}(I,J)^{\dagger}$ are no longer empty. This procedure is repeated, with new skeletons appended to previously selected ones, until the desired approximation accuracy is achieved.
Algorithms based on coordinate sampling
There have been several recent advances in coordinate sampling-based methods for ID/CUR decompositions.
In ChenEpperlyTroppWebber2025, the randomly pivoted (RP) Cholesky algorithm is introduced for positive semi-definite matrices $\bm{\mathsf{A}}\in\mathbb{F}^{m\times m}$. Given a target rank $k$, the algorithm outputs a set of pivot indices $S=\{s_{1},\ldots,s_{k}\}$ akin to skeletons, and a matrix $\bm{\mathsf{F}}\in\mathbb{F}^{m\times k}$ that defines the Nyström approximation $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{F}}\bm{\mathsf{F}}^{*}$. We begin by defining $\bm{\mathsf{d}}=\textup{diag}(\bm{\mathsf{A}})$ and initializing the matrix $\bm{\mathsf{F}}$ as $\bm{\mathsf{0}}$.
After sampling each pivot $s_{i}\sim\bm{\mathsf{d}}/\sum_{j=1}^{m}\bm{\mathsf{d}}(j)$, we form $\bm{\mathsf{g}}=\bm{\mathsf{A}}(:,s_{i})$ and remove any overlap with previously selected columns via
$\bm{\mathsf{g}}=\bm{\mathsf{g}}-\bm{\mathsf{F}}(:,1:(i-1))\bm{\mathsf{F}}(s_{i},1:(i-1))^{*}.$
The matrix $\bm{\mathsf{F}}$ is then updated via $\bm{\mathsf{F}}(:,i)=\bm{\mathsf{g}}/\sqrt{\bm{\mathsf{g}}(s_{i})}$, and the vector $\bm{\mathsf{d}}$ of diagonal entries of the residual matrix is updated via $\bm{\mathsf{d}}=\bm{\mathsf{d}}-|\bm{\mathsf{F}}(:,i)|^{2}$. To achieve an $\varepsilon$-approximation, $d\geq k/\varepsilon+k\log(\frac{1}{\varepsilon\eta})$ columns are required, where $\eta$ is the relative trace-norm error, cf. (ChenEpperlyTroppWebber2025, Theorem 2.3). A blocked variant of RP Cholesky is also provided in (ChenEpperlyTroppWebber2025, Algorithm 3). Although RP Cholesky is efficient, simple to implement, and comes with strong theoretical guarantees, it can only be applied to positive semi-definite matrices. If $\bm{\mathsf{A}}$ is not positive semi-definite, then we would need to form the Gram matrix $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$ before applying RP Cholesky, which is computationally undesirable.
The recent work of DongRBRP2025, inspired by RP Cholesky, introduces robust blockwise random pivoting (RBRP) for matrix IDs. Like RP Cholesky, the RBRP algorithm incorporates both randomness and adaptiveness, in that each iteration of skeleton selection is informed by previous ones. Blockwise random pivoting replaces greedy pivoting based on squared column norms in CPQR with squared-norm sampling, and the “robustness” of RBRP refers to a filtering step that eliminates redundant points within each block of selected skeleton candidates. Namely, suppose that we sample $b$ row skeleton candidates $S_{t}$ for some fixed block size $b$. We then form
$\bm{\mathsf{V}}=\bm{\mathsf{A}}(S_{t},:)^{*}-\bm{\mathsf{Q}}^{(t-1)}((\bm{\mathsf{Q}}^{(t-1)})^{*}\bm{\mathsf{A}}(S_{t},:)^{*})\in\mathbb{F}^{n\times b},$
where $\bm{\mathsf{Q}}^{(t-1)}$ is an orthonormal basis for the previously selected skeletons. Given some filtering tolerance $\tau$, we use CPQR to compute the factorization
$\bm{\mathsf{V}}(:,P)=\bm{\mathsf{Q}}\bm{\mathsf{R}}\approx\bm{\mathsf{Q}}(:,1:b^{\prime})\bm{\mathsf{R}}(1:b^{\prime},:),$
so that
$\|\bm{\mathsf{V}}(:,P)-\bm{\mathsf{Q}}(:,1:b^{\prime})\bm{\mathsf{R}}(1:b^{\prime},:)\|^{2}_{\text{F}}=\|\bm{\mathsf{R}}(b^{\prime}+1:b,b^{\prime}+1:b)\|^{2}_{\text{F}}<\tau\|\bm{\mathsf{V}}\|^{2}_{\text{F}}.$
This leads to a time complexity of roughly $O(mnk+nk^{2})$ for resulting skeleton cardinality $k$, as RBRP is also rank-adaptive, and the optimal associated interpolation matrix $\bm{\mathsf{W}}=\bm{\mathsf{A}}\bm{\mathsf{A}}(S,:)^{\dagger}$ can be computed efficiently in $O(mk^{2})$ time.
5 Tensor Preliminaries
This section marks the beginning of our exploration of tensors. We begin with formal notation and definitions in Sect. 5.1, presenting tensors as natural higher-order generalizations of matrices, from 2-way to $d$-way arrays. As a result, many of the randomized methods in the previous sections developed for matrices can be extended to tensors, though there are subtleties with higher-order inputs, and many concepts can be obscured by the unwieldy notation that frequently accompanies them. We try to mitigate this by providing visualizations, including a table of notation (Table 4) and illustrations of the matrix and tensor products in Sect. 5.2. In Sect. 5.3, we point out some peculiarities specific to higher-order tensors, such as tensor rank and uniqueness of decompositions.
While there are many different ways to represent tensors, we focus on two fundamental tensor decomposition formats: the canonical polyadic (CP) and Tucker. In Sect. 5.3, we lay out key concepts and practicalities for each of them, including algorithms commonly used to compute them. We then discuss particular low-rank tensor decompositions that can be represented in each format, such as higher-order extensions of the matrix SVD and ID/CUR. Throughout this section, we try to emphasize that the classical matrix methods covered in previous sections underlie those for tensors.
Symbol Definition
$\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$ Tensor with $d$ modes
$\|\bm{\mathcal{X}}\|$ Frobenius norm $\left(\sum_{i_{1}=1}^{N_{1}}\ldots\sum_{i_{d}=1}^{N_{d}}\bm{\mathcal{X}}(i_{1},\ldots,i_{d})^{2}\right)^{1/2}$
$N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$ Number of entries of $\bm{\mathcal{X}}$
$N^{(-j)}=\prod_{k\neq j}N_{j}$ Number of entries of $\bm{\mathcal{X}}$ excluding mode $j$
$\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$ Mode-$j$ unfolding of $\bm{\mathcal{X}}$
$\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}$ Kronecker product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J},\bm{\mathsf{B}}\in\mathbb{F}^{K\times L}$
$\bm{\mathsf{A}}\odot\bm{\mathsf{B}}$ Khatri-Rao product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times K},\bm{\mathsf{B}}\in\mathbb{F}^{J\times K}$
$\bm{\mathsf{A}}*\bm{\mathsf{B}}$ Hadamard product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J},\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$
Table 4: Summary of tensor notation.
5.1 Notation and Definitions
A tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times\cdots\times N_{d}}$ is a $d$-way array with entries $\bm{\mathcal{X}}_{i_{1},i_{2},\ldots,i_{d}}$, for all $i_{1}\in[N_{1}]$, $i_{2}\in[N_{2}]$, $\ldots$, $i_{d}\in[N_{d}]$. A $d$-way tensor is said to be of order $d$, and each of its $d$ ways is referred to as a mode; a matrix, for instance, is a tensor of order 2, with mode 1 comprised of its rows and mode 2 comprised of its columns. A mode-$j$ fiber of a tensor is a vector obtained by fixing all indices except for the $j$th, e.g., a column of a matrix is a mode-1 fiber, and a row is a mode-2 fiber. The mode-$j$ slices of a tensor are the sub-tensors $\bm{\mathcal{X}}(:,\ldots,:,i_{j},:,\ldots,:)$ obtained by fixing a single index $i_{j}\in[N_{j}]$, e.g., slices of 3-mode tensors are matrices.
We define $N^{(-j)}=\prod_{k=1,k\neq j}^{d}N_{k}$ as the product of all dimensions except mode $j$, and $N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$. Through a rearrangement of entries known as matricization, a $d$-mode tensor can be “unfolded” into a matrix. The mode-$j$ unfolding of a $d$-mode tensor $\bm{\mathcal{X}}$, denoted $\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$, $j=1,\ldots,d$, is a matrix whose columns are the mode-$j$ fibers of $\bm{\mathcal{X}}$. Throughout this section, we assume that the $d$ modes are ordered $1,\ldots,d$ for notational simplicity, but for computational reasons, it may be beneficial to re-order the modes for processing.
In our work, the (Frobenius) norm of a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times\cdots\times N_{d}}$, denoted by $\|\bm{\mathcal{X}}\|$, is the square root of the sum of squares of all elements, analogous to the Frobenius norm for matrices. A tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times\cdots\times N_{d}}$ is said to be rank-one if it can be expressed as a vector outer product (denoted by $\circ$) of $d$ vectors, i.e. $\bm{\mathcal{X}}=\bm{\mathsf{v}}^{(1)}\circ\bm{\mathsf{v}}^{(2)}\circ\cdots\circ\bm{\mathsf{v}}^{(d)}$.
5.2 Matrix and Tensor Products
We now define operations involving matrices and tensors used throughout our work.
The Kronecker product of matrices $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J}$ and $\bm{\mathsf{B}}\in\mathbb{F}^{K\times L}$, denoted $\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}$, is the $(IK)\times(JL)$ matrix with entries
$(\bm{\mathsf{A}}\otimes\bm{\mathsf{B}})(K(i-1)+k,L(j-1)+\ell)=\bm{\mathsf{A}}(i,j)\bm{\mathsf{B}}(k,\ell).$
The Kronecker product then takes the form
$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,J}\end{bmatrix}}\otimes\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,L}\\ \vdots&\ddots&\vdots\\ b_{K,1}&\cdots&b_{K,L}\end{bmatrix}}$ $=\underset{\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&\cdots&a_{1,1}b_{1,L}&\cdots\cdots&a_{1,J}b_{1,1}&\cdots&a_{1,J}b_{1,L}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ a_{1,1}b_{K,1}&\cdots&a_{1,1}b_{K,L}&\cdots\cdots&a_{1,J}b_{K,1}&\cdots&a_{1,J}b_{K,L}\\ \vdots&&&&&&\vdots\\ \vdots&&&\ddots&&&\vdots\\ \vdots&&&&&&\vdots\\ a_{I,1}b_{1,1}&\cdots&a_{I,1}b_{1,L}&\cdots\cdots&a_{I,J}b_{1,1}&\cdots&a_{I,J}b_{1,L}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ a_{I,1}b_{K,1}&\cdots&a_{I,1}b_{K,L}&\cdots\cdots&a_{I,J}b_{K,1}&\cdots&a_{I,J}b_{K,L}\end{bmatrix}}$
$=\begin{bmatrix}a_{1,1}\bm{\mathsf{B}}&\cdots&a_{1,J}\bm{\mathsf{B}}\\ \vdots&\ddots&\vdots\\ a_{I,1}\bm{\mathsf{B}}&\cdots&a_{I,J}\bm{\mathsf{B}}\end{bmatrix}$.
The Khatri-Rao product of matrices $\bm{\mathsf{A}}\in\mathbb{F}^{I\times K}$ and $\bm{\mathsf{B}}\in\mathbb{F}^{J\times K}$, denoted $\bm{\mathsf{A}}\odot\bm{\mathsf{B}}$, is the $(IJ)\times K$ matrix with entries
$(\bm{\mathsf{A}}\odot\bm{\mathsf{B}})(:,k)=\bm{\mathsf{A}}(:,k)\otimes\bm{\mathsf{B}}(:,k)$
for $k=1,\ldots,K.$ The Khatri-Rao product takes the form
$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,K}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,K}\end{bmatrix}}\odot\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,K}\\ \vdots&\ddots&\vdots\\ b_{J,1}&\cdots&b_{J,K}\end{bmatrix}}$ $=\underset{\bm{\mathsf{A}}\odot\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&a_{1,2}b_{1,2}&\cdots&a_{1,K}b_{1,K}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,1}b_{J,1}&a_{1,2}b_{J,2}&\cdots&a_{1,K}b_{J,K}\\ \vdots&\vdots&&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ \vdots&\vdots&&\vdots\\ a_{I,1}b_{1,1}&a_{I,2}b_{1,2}&\cdots&a_{I,K}b_{1,K}\\ \vdots&\vdots&\ddots&\vdots\\ a_{I,1}b_{J,1}&a_{I,2}b_{J,2}&\cdots&a_{I,K}b_{J,K}\\ \end{bmatrix}}$
$=\begin{bmatrix}\bm{\mathsf{a}}_{1}\otimes\bm{\mathsf{b}}_{1}&\bm{\mathsf{a}}_{2}\otimes\bm{\mathsf{b}}_{2}&\cdots&\bm{\mathsf{a}}_{K}\otimes\bm{\mathsf{b}}_{K}\end{bmatrix}$,
where $\bm{\mathsf{a}}_{i}$ and $\bm{\mathsf{b}}_{i}$ are the columns of $\bm{\mathsf{A}}$ and $\bm{\mathsf{B}}$. There is a useful property relating the Kronecker and Khatri-Rao products: if $\bm{\mathsf{A}}\in\mathbb{F}^{K\times J}$, $\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$, $\bm{\mathsf{C}}\in\mathbb{F}^{I\times K}$, and $\bm{\mathsf{D}}\in\mathbb{F}^{J\times I}$,
$(\bm{\mathsf{C}}\otimes\bm{\mathsf{D}})(\bm{\mathsf{A}}\odot\bm{\mathsf{B}})=(\bm{\mathsf{C}}\bm{\mathsf{A}})\odot(\bm{\mathsf{D}}\bm{\mathsf{B}}).$
Additionally we have the following inequality of matrix norms:
$\|\bm{\mathsf{A}}\odot\bm{\mathsf{B}}\|^{2}_{\text{F}}\leq\|\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}\|^{2}=\|\bm{\mathsf{A}}\|^{2}\|\bm{\mathsf{B}}\|^{2}.$
The Hadamard product of matrices $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J}$ and $\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$, denoted $\bm{\mathsf{A}}*\bm{\mathsf{B}}$, is the $I\times J$ matrix of entries
$(\bm{\mathsf{A}}*\bm{\mathsf{B}})(i,j)=\bm{\mathsf{A}}(i,j)\bm{\mathsf{B}}(i,j).$
The Hadamard product can be visualized as element-wise multiplication,
$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,J}\end{bmatrix}}*\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,J}\\ \vdots&\ddots&\vdots\\ b_{I,1}&\cdots&b_{I,J}\end{bmatrix}}=\underset{\bm{\mathsf{A}}*\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&\cdots&a_{1,J}b_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}b_{I,1}&\cdots&a_{I,J}b_{I,J}\end{bmatrix}}.$
The tensor-times-matrix (TTM) product is the tensor resulting from the mode-$j$ product of a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{j}\times\cdots\times N_{d}}$ with a matrix $\bm{\mathsf{U}}\in\mathbb{F}^{K\times N_{j}}$, denoted by $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{j}\bm{\mathsf{U}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{j-1}\times K\times N_{j+1}\times\cdots\times N_{d}}$, with entries
$\bm{\mathcal{Y}}(i_{1},\ldots,i_{j-1},k,i_{j+1},\ldots,i_{d})=\sum_{i_{j}=1}^{N_{j}}\bm{\mathcal{X}}(i_{1},\ldots,i_{d})\bm{\mathsf{U}}(k,i_{j}).$
We can express this more concisely with mode-$j$ matricization: $\bm{\mathsf{Y}}_{(j)}=\bm{\mathsf{U}}\bm{\mathsf{X}}_{(j)}$.
The Multi-TTM product extends the TTM product to multiple modes. For $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{j}\times\cdots\times N_{d}}$ and matrices $\bm{\mathsf{U}}_{j}\in\mathbb{F}^{K_{j}\times N_{j}}$ for $j\in[d]$, a Multi-TTM is given by
$\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{U}}_{1}\times_{2}\bm{\mathsf{U}}_{2}\cdots\times_{d}\bm{\mathsf{U}}_{d}.$
The order of multiplication does not matter, cf. (Ballard_Kolda_2025, Proposition 3.19). Equivalently, after mode-$j$ matricization,
$\bm{\mathsf{Y}}_{(j)}=\bm{\mathsf{U}}_{j}\bm{\mathsf{X}}_{(j)}\left(\bm{\mathsf{U}}_{d}\otimes\cdots\otimes\bm{\mathsf{U}}_{j+1}\otimes\bm{\mathsf{U}}_{j-1}\otimes\cdots\otimes\bm{\mathsf{U}}_{1}\right)^{\top}.$
5.3 Tensor Decompositions
This work considers two fundamental formats of tensor decompositions: the canonical polyadic (CP) decomposition and the Tucker decomposition. We introduce them in Sect. 5.3 and 5.3, along with prototypical algorithms to compute them.
The CP decomposition has been known by many different names since its introduction in 1927: the polyadic form Hitchcock1927TheEO, PARAFAC (parallel factors) Harshman1970, CANDECOMP (canonical decomposition) carroll1970analysis; see also (KoldaBader2009, Table 3.1). It is often viewed as a higher-order extension of the matrix SVD KoldaBader2009.
The Tucker decomposition Tuck1963a; tucker1964; Tucker1966 is a representation of a tensor as a Multi-TTM product of a core tensor with a factor matrix for each mode. Like the CP decomposition, the Tucker decomposition has been known by several different names in the literature (KoldaBader2009, Table 4.1), and the Tucker decomposition can also be viewed as an extension of the matrix SVD, in the form of higher-order principal component analysis (PCA) kapteyn1986approach; kroonenberg1980principal; LeiboviciSabatier1998; Vasilescu2002.
We briefly summarize a few key properties of these decompositions in Sect. 5.3 before discussing low-rank tensor approximations that can be expressed in each format in Sect. 5.3.
The Canonical Polyadic (CP) Decomposition
The CP decomposition of a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$ expresses it as a sum of $R$ rank-one tensors:
$\bm{\mathcal{X}}=\sum_{r=1}^{R}\bm{\mathsf{a}}^{(1)}_{r}\circ\bm{\mathsf{a}}^{(2)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r},\hskip 11.38109pt\textup{where }\bm{\mathsf{a}}_{r}^{(j)}\in\mathbb{F}^{N_{j}}\ \textup{ for }j=1,\ldots,d.$ (43)
For example, the CP decomposition of a 3-mode tensor is visualized in Fig. 1.
$N_{2}\underbrace{\hskip 51.21504pt}$$N_{1}\underbrace{\hskip 34.14322pt}$$N_{3}\underbrace{\hskip 22.76228pt}$$\bm{\mathcal{X}}$$\bm{\mathsf{a}}_{1}$$\bm{\mathsf{b}}_{1}$$\bm{\mathsf{c}}_{1}$$++\cdots$$\bm{\mathsf{a}}_{2}$$\bm{\mathsf{b}}_{2}$$\bm{\mathsf{c}}_{2}$$\bm{\mathsf{a}}_{R}$$\bm{\mathsf{b}}_{R}$$\bm{\mathsf{c}}_{R}$
Figure 1: Rank-$R$ CP decomposition $\bm{\mathcal{X}}=\sum_{r=1}^{R}\bm{\mathsf{a}}_{r}\circ\bm{\mathsf{b}}_{r}\circ\bm{\mathsf{c}}_{r}$ for 3-mode tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$.
The factor matrices of the CP decomposition are formed from the vectors corresponding to each of the $d$ modes. Namely, the mode-$j$ factor matrix is given by
$\bm{\mathsf{A}}^{(j)}=\begin{bmatrix}\bm{\mathsf{a}}_{1}^{(j)}&\bm{\mathsf{a}}_{2}^{(j)}&\ldots&\bm{\mathsf{a}}_{R}^{(j)}\end{bmatrix}\in\mathbb{F}^{N_{j}\times R}.$
We can express (43) in terms of mode-$j$ unfoldings using the Khatri-Rao product:
$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}.$
The notation of Kolda2006 is frequently used as shorthand for the CP decomposition:
$\bm{\mathcal{X}}=\llbracket\bm{\mathsf{A}}^{(1)},\bm{\mathsf{A}}^{(2)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket:=\sum_{r=1}^{R}\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}$. (44)
If we enforce unit length, a weight vector $\bm{\lambda}\in\mathbb{F}^{R}$ may be incorporated into (44), so that
$\bm{\mathcal{X}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\bm{\mathsf{A}}^{(2)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket:=\sum_{r=1}^{R}\lambda_{r}\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}$. (45)
In practice, the CP decomposition is often computed by the alternating least squares (ALS) method carroll1970analysis; Harshman1970. Pseudocode for the general ALS procedure is given in Algorithm 5, which we quickly summarize.
Given a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the ALS procedure outputs a CP decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$ with $R$ rank-one components, once some convergence criteria is met. This is accomplished through a sequence of linear least squares problems, where all but one factor matrix is fixed:
$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{X}}_{(j)}-\widehat{\bm{\mathsf{A}}}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right\|^{2}_{\text{F}}.$ (46)
The solution of (46) is given by
$\bm{\mathsf{A}}^{(j)}:=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right]^{\dagger},$ (47)
where the columns of $\bm{\mathsf{A}}^{(j)}$ are normalized to determine $\lambda_{r}$ for $r=1,\ldots,R$. This process is repeated for each mode until some convergence criterion is met; see Sect. 5.3.
If Algorithm 5 is naively implemented, the per-iteration cost to compute a CP decomposition is dominated by $N=\textup{size}(\bm{\mathcal{X}})$, which is computationally undesirable. Fortunately, since its introduction in 1970, CP-ALS has benefited from many modifications for improved efficiency, cf. Sect. 6.1 for recent algorithmic developments.
One strategy that has garnered attention for efficient CP-ALS incorporates a preprocessing step to compress the input tensor before executing Algorithm 5; this method is known as CANDELINC BroAndersson1998PartII; Carroll1980. The compression is often achieved through a Tucker decomposition, which we discuss next.
Algorithm 5 Alternating Least Squares for CP Decomposition (CP-ALS)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, number of rank-one components $R$
2:CP decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
3:Initialize $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R}$, $j=1,\ldots,d$ $\triangleright$ e.g., $\bm{\mathsf{A}}^{(j)}=\texttt{randn}(N_{j},R)$
4:while not converged do
5:  for $j=1,\ldots,d$ do
6:   $\bm{\mathsf{A}}^{(j)}=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right]^{\dagger}$
7:   $\lambda_{r}=\left\|\bm{\mathsf{A}}^{(j)}(:,r)\right\|_{2}$, $r=1,\ldots,R$
8:   $\bm{\mathsf{A}}^{(j)}(:,r)=\bm{\mathsf{A}}^{(j)}(:,r)/\lambda_{r}$, $r=1,\ldots,R$
9:   Update $\widehat{\bm{\mathcal{X}}}=\llbracket\lambda;\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
10:  end for
11:end while
Tucker Decomposition
The Tucker decomposition expresses a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$ as a $d$-way Multi-TTM product:
$\bm{\mathcal{X}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\times_{2}\bm{\mathsf{A}}^{(2)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}=:\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\bm{\mathsf{A}}^{(2)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket,$ (48)
where $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$ is called the core tensor, and $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$ is the factor matrix for mode $j=1,\ldots,d$. We can express (48) in terms of mode-$j$ unfoldings by
$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\bm{\mathsf{G}}_{(j)}\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)^{\top}.$ (49)
A Tucker decomposition of a 3-mode tensor is visualized in Fig. 2.
[…]
Figure 2: Rank-$(R_{1},R_{2},R_{3})$ Tucker decomposition $\bm{\mathcal{X}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\times_{2}\bm{\mathsf{A}}^{(2)}\times_{3}\bm{\mathsf{A}}^{(3)}$ of $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$.
A prototypical algorithm to compute the Tucker decomposition is Tucker’s “Method 1” Tucker1966, which has come to be known as the higher-order SVD (HOSVD) DeLathauwer2000a, summarized in Algorithm 6. For an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$, each factor matrix $\bm{\mathsf{A}}^{(j)}$ is computed as the leading-$R_{j}$ left singular vectors of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$. If $R_{j}<\textup{rank}(\bm{\mathsf{X}}_{(j)})$, the decomposition is said to be “truncated.” When $R_{j}\ll N_{j}$, the Tucker decomposition is said to be a compressed representation of the input tensor. Once the factor matrices are computed, the core tensor is obtained as a Multi-TTM product of the original tensor with the factor matrix transposes. The computational cost of the HOSVD is $O(dn^{d+1}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity.
For improved efficiency, the sequentially truncated HOSVD (ST-HOSVD) is often used Hackbusch2019; Vannieuwenhoven2012. In the ST-HOSVD, once a factor matrix is computed for mode $j$, the input tensor is compressed via mode-$j$ TTM before the next factor matrix is computed; i.e., for input tensor $\bm{\mathcal{X}}$, we first compute $\bm{\mathsf{A}}^{(1)}$ and form $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{A}}^{(1)}$, then compute $\bm{\mathsf{A}}^{(2)}$ from $\bm{\mathcal{Y}}$, then form $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{2}\bm{\mathsf{A}}^{(2)}$, and so on. ST-HOSVD successively reduces the size of the tensor used in subsequent computations, and the core tensor is available immediately after computing the factor matrix for the last mode. The computational cost of ST-HOSVD is $O(\sum_{j=1}^{d}k^{j-1}n^{d-j+2}+k^{j}n^{d-j})$.
Algorithm 6 Higher-order SVD for Tucker Decomposition (HOSVD)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, factor matrix target ranks $R_{1},\ldots,R_{d}$
2:Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
3:for $j=1,\ldots,d$ do
4:  $\bm{\mathsf{A}}^{(j)},\sim,\sim]=\texttt{svd}(\bm{\mathsf{X}}_{(j)},R_{j})$
5:end for
[…]
Unlike the matrix SVD, the (ST-)HOSVD does not yield optimal approximation error. It has been shown that the relative error with (ST-)HOSVD is within $\sqrt{d}$ of an optimal Tucker decomposition of a given rank for a $d$-mode tensor, cf. (Ballard_Kolda_2025, Chapter 7). However, we can use the HOSVD output as a starting point for an ALS-style algorithm to compute Tucker decompositions, known as higher-order orthogonal iteration (HOOI) kroonenberg1980principal; DeLathauwer2000a. Pseudocode for HOOI is given in Algorithm 7.
As in CP-ALS, in HOOI, we seek to minimize the approximation error $\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|$, now over $\bm{\mathcal{G}}$ and $\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}$, subject to the constraints that $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1},\ldots,R_{d}}$ and that $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$ is orthogonal for $j=1,\ldots,d$. This tensorial minimization problem can be transformed into matrix least squares problems in each mode $j$, where we solve for each factor matrix via
$\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\widehat{\bm{\mathsf{A}}}^{\top}-\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}.$
Note the optimal solution to this least squares problem is given by the leading left singular vectors of each unfolding $\bm{\mathsf{X}}_{(j)}$ Andersson1998; DeLathauwer2000; Kolda2006. The core tensor can be computed as a column vector $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{x}}\right\|^{2}_{2}$, where $\bm{\mathsf{x}}$ is the vectorization or unfolding of tensor $\bm{\mathcal{X}}$ into a large column vector of size $N\times 1$. To get $\bm{\mathcal{G}}$, we “re-fold” $\bm{\mathsf{g}}\in\mathbb{F}^{\prod_{j}R_{j}}$ back into a tensor of dimension $\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$. If the factor matrices are computed as the leading $R_{j}$ left singular vectors of unfoldings, then the optimal core tensor is given by $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\left(\bm{\mathsf{A}}^{(1)}\right)^{\top}\times\cdots\times_{d}\left(\bm{\mathsf{A}}^{(d)}\right)^{\top}$. The per-iteration cost of Algorithm 7 is $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$.
While Algorithm 7 will converge to a local solution where the approximation error no longer decreases, it may not converge to the global optimum DeLathauwer2000; kroonenberg1980principal. Modified versions with better guarantees have been proposed, such as HOOI with Newton-Grassmann optimization Elden2009, but these come at greater computational cost. The Tucker decomposition may also be used as a preprocessing compression step to achieve better performance in CP-ALS, known as CANDELINC or a Tucker+CP approach BroAndersson1998PartII.
In the next section, we offer some practical guidelines for choosing a tensor format and target rank for a given problem, beginning with a discussion of tensor rank and ending with a comparison of the presented tensor formats.
Algorithm 7 Higher-order Orthogonal Iteration for Tucker Decomposition (HOOI)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, factor matrix ranks $R_{1},\ldots,R_{d}$
2:Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
3:Initialize $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$, $j=2,\ldots,d$
4:while not converged do
5:  for $j=1,\ldots,d$ do
6:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}$
7:  end for
8:  $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{x}}\right\|^{2}_{2}$ $\triangleright$ Column vectorization $\bm{\mathsf{x}}$ of $\bm{\mathcal{X}}$
9:  Re-fold vectorization $\bm{\mathsf{g}}$ into $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$
10:end while
Rank, Uniqueness, and the Choice of Tensor Representation
The rank of a tensor $\bm{\mathcal{X}}$ is the minimal number of rank-one tensors that are required for equality in (44):
$\textup{rank}(\bm{\mathcal{X}})=\min\{R\in\mathbb{N}\ |\ \bm{\mathcal{X}}=\llbracket\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket,\ \bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R}\ \forall j\in[d]\}.$ (50)
Mode ordering and TTM products with nonsingular matrices do not affect tensor rank. However, it is NP-hard to determine Hrastad1990.
In practice, the target ranks for low-rank tensor decompositions are chosen heuristically. For example, we might choose the smallest rank $R$ for a CP decomposition that significantly reduces the relative error, in comparison to rank R1R-1, cf. (Ballard_Kolda_2025, Section 9.4.1). Generally, the optimization procedure in CP-ALS is repeated until the error no longer decreases or a maximum number of iterations is reached.
For Tucker decompositions, it is useful to consider the $j$-rank of a tensor, denoted by $\textup{rank}_{j}(\bm{\mathcal{X}})$ and defined as the column rank of its mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$ kruskal1989rank; DeLathauwer2000a. If $R_{j}=\textup{rank}_{j}(\bm{\mathcal{X}})$ for $j=1,\ldots,d$, then we say that $\bm{\mathcal{X}}$ is a rank-$(R_{1},\ldots,R_{d})$ tensor (distinct from its rank, defined above). We can compute an exact rank-$(R_{1},\ldots,R_{d})$ Tucker decomposition for any tensor $\bm{\mathcal{X}}$ with $R_{j}=\textup{rank}_{j}(\bm{\mathcal{X}})$ for $j=1,\ldots,d.$ However, if $R_{j}<\textup{rank}_{j}(\bm{\mathcal{X}})$ for any $j$, the decomposition is inexact and more difficult to compute. In practice, we often prescribe a core $j$-rank for $j=1,\ldots,d$ that meets some error tolerance or compression ratio (Ballard_Kolda_2025, Section 4.2). The (ST-)HOSVD algorithm has the nice property that it computes a rank-$(R_{1},\ldots,R_{d})$ decomposition with an approximation error that is within $\sqrt{d}$ of the optimal rank-$(R_{1},\ldots,R_{d})$ approximation error.
Whether it is better to represent an input tensor in the CP or Tucker format is generally problem-dependent. Under mild assumptions, the CP decomposition of a tensor is unique, up to scaling and permutation of its rank-one components tenberge2002uniqueness; harshman1984data; Harshman1970; kruskal1977three; kruskal1989rank; sidiropoulos2000uniqueness. The same cannot be said for Tucker decompositions; the mode-$j$ factor matrix can be multiplied by a nonsingular matrix and the core tensor updated via mode-$j$ product of the original core tensor with the inverse of the nonsingular matrix. As a result, the factors of a Tucker decomposition are less interpretable in the context of the original data than those of a CP decomposition, cf. (Ballard_Kolda_2025, Section 9.6–9.8).
However, as previously mentioned, the Tucker decomposition is highly advantageous for data compression, even as a preprocessing step for a CP decomposition, e.g., CANDELINC BroAndersson1998PartII; Carroll1980. Despite its utility in data compression, the Tucker decomposition still suffers from the curse of dimensionality: letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the cost of storing a Tucker decomposition is $O(k^{d}+dkn)$. By contrast, the storage cost of a CP decomposition is linear in its rank, but its rank may be very large to achieve the desired approximation accuracy.
In short, the notion of rank for tensor decompositions is more nuanced than for matrices, and the choice of representation and target rank (or $j$-rank) should be informed by the desired trade-off between approximation accuracy and computational performance. For more detailed treatments, we refer the reader to Ballard_Kolda_2025; DeLathauwer2000a; Grasedyck2013; KoldaBader2009.
Low-Rank Tensor Factorizations
In the previous sections, we introduced two fundamental tensor representations, the CP and the Tucker decomposition, and we summarized their key properties. In this section, we survey low-rank tensor factorizations that can be expressed in each format, which we characterize primarily by whether or not the factor matrices are orthogonal. Namely, we assume that the standard Tucker decomposition has orthonormal factor matrices, whereas CP decompositions generally do not. Higher-order versions of ID/CUR may be expressed using orthogonal factor matrices or factor matrices that contain entries of the input tensor, so we further characterize them by whether the core subtensor or the factor matrices contain entries of the original tensor.
Orthogonal factor matrices
Recall from Sect. 5.3 that the (ST-)HOSVD and HOOI methods used to compute Tucker decompositions return orthonormal factor matrices $\bm{\mathsf{A}}^{(j)}$ for $j=1,\ldots,d$. Alternatively, there are algorithms for Tucker decompositions where orthogonality of factor matrices is not enforced, such as algorithms for the tensor ID/CUR factorizations in which factor matrices are computed as sub-matrices of unfoldings. Other examples beyond the scope of the current work include diagonalization algorithms for Tucker decompositions Tichavsky2012; Tichavsky2017 or algorithms that compute non-negative or sparse Tucker factorizations Kim2007; Morup2008. These algorithms may be preferable in specific application domains, such as image processing Sietsema2024; Zhou2015NTD, but in general, many tensor computations are greatly simplified when the Tucker factor matrices are orthogonal. Moreover, the (ST-)HOSVD with enforced orthogonality comes with quasi-optimality guarantees. As such, with the exception of the tensor ID/CUR, we focus on algorithms for Tucker decompositions that compute orthogonal (in fact, orthonormal) factor matrices. There is also historical precedence for this decision, as the seminal algorithms to compute Tucker decompositions enforced the orthogonality of factor matrices DeLathauwer2000a; DeLathauwer2000; Drineas2007tensorSVD; Kolda2001; tucker1964; Tucker1966; Vannieuwenhoven2012.
Finally, we note that the factor matrices in a CP decomposition are not orthogonal in general. Some tensors may admit CP decompositions with orthogonal factors Anandkumar2014; Robeva2016, called ODECO tensors; however, ODECO tensors occur with probability 0 Kolda2001; Kolda2003; Kolda2015b. This is in contrast with Tucker factor matrices, which can be transformed straightforwardly into orthogonal matrices, cf. (Ballard_Kolda_2025, Sections 5.4, 17.1.4). As such, in this work, we focus solely on Tucker decompositions in our discussion of randomized algorithms to compute orthogonal factor matrices in Sect. 6.
Tensor ID and CUR factorizations
As in the matrix setting (Sect. 2.4), a tensor ID or CUR factorization comprises actual entries of the input tensor. The advantages of ID/CUR factorizations for tensors are similar to those for matrices. The tensor ID/CUR factorization can be interpreted more easily in the context of the original tensorial data than factorizations involving multi-linear transformations of the data. Moreover, the ID/CUR preserves the structure of the input tensor (e.g., sparsity or non-negativity), and it can be computed efficiently with minimal storage requirements.
Analogous to the different versions of matrix ID/CUR (e.g., row, column, double-sided, CUR), there are several variants in the tensor setting that have been investigated since the ID/CUR was first extended to 3-mode tensors in Drineas2007tensorSVD. In general, these variants fall into one of two categories:
  1. 1.
    The core tensor (in the Tucker format) is a sub-tensor of the input tensor, or
  2. 2.
    The factor matrices (in the CP or Tucker format) are sub-matrices of unfoldings.
The work of MahoneyMaggioniDrineas2008_tensorCUR that introduced “tensor CUR factorizations” falls into the first category. In their algorithm, a $\textup{rank-}(N_{1},N_{2},R)$ core tensor is constructed by selecting $R$ indices from one mode and keeping all indices from the other two. Core tensor skeleton selection methods for 3-mode tensor CUR were generalized to $d$-mode tensors in Caiafa2010, for Tucker decompositions with the same $j$-rank in each mode. The “structure-preserving ST-HOSVD” of MinsterSaibabaKilmer2020, initially developed for sparse rank-$(R_{1},\ldots,R_{d})$ tensors, may also be interpreted as a tensor CUR factorization of the first category. The “Chidori” and “Fiber” tensor CUR factorizations with rank-$(R_{1},\ldots,R_{d})$ core tensors were characterized in Cai2021, inspired by those in Caiafa2010; these factorizations were utilized for robust tensorial principal component analysis in Cai2024. The recently proposed “CoreID” in Zhang2025 is similar to the CUR factorizations of Cai2021. In all of these methods, the entries of the associated factor matrices are computed based on the specific form of the factorization and on the skeletons selected for the core tensor, but the factor matrices are not themselves sub-matrices of unfoldings.
The second category above comprises those tensor ID/CUR factorizations in which the entries of the factor matrices are entries of the input tensor. The early algorithm of Drineas2007tensorSVD to compute a $\textup{rank-}(R_{1},\ldots,R_{d})$ “approximate tensor SVD” forms the factor matrices in a Tucker decomposition as randomly sampled columns of unfoldings; this factorization would later be termed a “higher-order ID” in Saibaba2016, where the mode-$j$ factor matrix in a Tucker decomposition is computed as column skeletons of the mode-$j$ unfolding selected by greedy pivoting. In FriedlandMehrmannMiedlarNkengla2011, an HOOI procedure is developed for 3-mode Tucker tensors that yields factor matrices of original tensor entries, and the “CURT” algorithm of ZhaoSong2018RETL; Song2019 computes tensor CUR decompositions in the Tucker format with strong accuracy guarantees. The CURT algorithm was recently generalized in Zhang2025 to compute rank-$(R_{1},\ldots,R_{d})$ Tucker decompositions where the core tensor is unconstrained and the factor matrices are column skeletons of mode unfoldings; this algorithm is also suitable for input tensors expressed in the CP format. The tensor ID algorithm of Biagioni2015 computes a CP decomposition of a $d$-mode tensor, in which factor matrices are comprised of entries of the input tensor, as does the more recent work of MalikBecker2020. We detail several algorithms that leverage randomization to compute tensor ID/CUR factorizations efficiently in Sect. 6.
6 Randomized Low-Rank Tensor Decompositions
As in the matrix setting, we are interested in algorithms that compute low-rank tensor decompositions accurately and efficiently, with provable performance guarantees that are substantiated by empirical evidence. Because many tensor algorithms fundamentally rely on classical matrix algorithms, we can frequently fall back on the ideas introduced in Sect. 4 to leverage randomization for efficient tensor decompositions. However, given the rapidly increasing number of applications that involve large-scale tensorial data, it is often impractical or infeasible to rely on basic implementations of the randomized rangefinder, e.g., with Gaussian sketching. The curse of dimensionality for $d$-way arrays is even more pernicious than for matrices, so great care must be taken to develop randomized algorithms that are not only accurate, but also scalable.
In this section, we survey new developments in fast randomized algorithms for low-rank tensor decompositions in the CP and Tucker formats. We draw connections and refer back to our presented matrix algorithms when relevant, augmenting them with emerging ideas from the recent proliferation of work on randomized algorithms for tensors. Sect. 6.1 explores how CP-ALS can be accelerated with randomization, specifically through fast sampling methods that exploit the structure of the overdetermined least squares problems solved in each iteration. Sect. 6.2 summarizes randomized methods to compute Tucker decompositions having orthogonal factor matrices, which naturally extend the randomized rangefinder to the tensor setting via the standard randomized (ST-)HOSVD and randomized HOOI, as well as a newly proposed algorithm for more efficient sketching. These randomized algorithms are contrasted with those for tensor ID/CUR decompositions in Sect. 6.3, which includes both the CP and Tucker formats. Our intent throughout this section is not to provide an exhaustive summary of all recent developments, but rather to broadly survey them to direct the interested reader to the relevant literature, and then hone in on key ideas that may stimulate future work in these emergent areas.
6.1 Randomized Algorithms for CP Decomposition
Over the last decade, many randomized algorithms have been proposed to compute CP decompositions more efficiently. We survey several notable algorithms, before summarizing the recent randomized CP-ALS procedure of Larsen2022, which underlies the high-performance algorithm of Bharadwaj2024, as an illustration of how randomization can be leveraged for efficient CP decompositions.
Many randomized algorithms for CP decompositions seek to exploit the structure of the Khatri-Rao product (KRP) in CP-ALS, cf. (47), for fast sketching and sampling. In Wang2015, the TensorSketch algorithm of Pham2013 is used to sketch the input tensor without explicitly forming it, then the KRP is efficiently computed. The “randomized ALS” algorithm of Reynolds2016 incorporates projections onto random tensors to improve the conditioning of the least squares problems in CP-ALS. A procedure to sample rows of the KRP near-optimally via fast leverage score approximations is presented in Cheng2016; however, the least squares problem is not sketched in this work (only the KRP). In BattaglinoBallardKolda2018, a Kronecker fast Johnson-Lindenstrauss transform (KFJLT) is applied to the KRP to reduce the coherence for better sampling performance, cf. Jin2021; MALIK2020120, which is then utilized in a sketched ALS problem without explicitly forming the KRP. The algorithm of Aggour2020 builds on BattaglinoBallardKolda2018 with options for adaptive sketch sizes and regularization terms. We now expand on the recent work of Larsen2022 on approximate leverage score sampling that exploits the structure of the KRP.
Randomized CP-ALS with Fast Leverage Score Sampling
Recall from Sect. 5.3 that within each iteration of CP-ALS (Algorithm 5), we compute a rank-$R$ CP decomposition that approximates our input tensor by solving $d$ least squares problems in succession. More precisely, given input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, in each iteration of CP-ALS, we solve a least squares problem in each mode $j=1,\ldots,d$, in which every factor matrix except $\bm{\mathsf{A}}^{(j)}$ is fixed, and $\bm{\mathsf{A}}^{(j)}$ is computed as the solution of
$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\widehat{\bm{\mathsf{A}}}\bm{\mathsf{Z}}^{\top}-\bm{\mathsf{X}}_{(j)}\right\|^{2}_{\text{F}}=\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{Z}}\widehat{\bm{\mathsf{A}}}^{\top}-\bm{\mathsf{X}}_{(j)}^{\top}\right\|^{2}_{\text{F}},$ (51)
where $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\in\mathbb{F}^{N^{(-j)}\times R}$.
If we solve (51) without exploiting any inherent problem structure, the cost of computing the QR decomposition of $\bm{\mathsf{Z}}$ is $O(R^{2}N^{(-j)})$, and the cost of applying it $N_{j}$ times brings the total cost to $O(R^{2}N^{(-j)}+RN)$.
In Larsen2022, a randomized sampling strategy based on fast leverage score approximation, CP-ARLS-LEV, is used to accelerate each least squares sub-problem (51). Namely, $S\ll N^{(-j)}$ rows of $\bm{\mathsf{Z}}$ are sampled according to an approximate leverage score distribution. The corresponding $S$ rows of $\bm{\mathsf{X}}_{(j)}^{\top}$ are then selected for a smaller “sketched” problem whose solution is a good approximation of the solution of (51) with high probability. Sketching reduces the cost of solving the $j$th least squares problem to $O(R^{2}S+RSN_{j})$. However, forming $\bm{\mathsf{Z}}$ or $\bm{\mathsf{X}}$ and computing leverage scores would incur a cost of $O(R^{2}N^{(-j)})$. To avoid this, in CP-ARLS-LEV, the leverage scores are estimated as the product of leverage scores of the respective factor matrices in the KRP, and the KRP structure is exploited so that only the $S$ selected rows of $\bm{\mathsf{Z}}$ are formed explicitly.
To this end, let $j\in[d]$ be fixed. We first observe that there is a bijection mapping each row $i\in[N]$ of $\bm{\mathsf{Z}}$ to a (d1)(d-1)-tuple $(i_{d},\ldots,i_{j+1},i_{j-1},\ldots,i_{1})$, with $i_{k}\in[N_{k}]$, $k\neq j,$ which indexes rows of factor matrices in the KRP:
$\bm{\mathsf{Z}}(i,:)=\bm{\mathsf{A}}^{(d)}(i_{d},:)*\cdots*\bm{\mathsf{A}}^{(j+1)}(i_{j+1},:)*\bm{\mathsf{A}}^{(j-1)}(i_{j-1},:)*\cdots*\bm{\mathsf{A}}^{(1)}(i_{1},:),$ (52)
where * again denotes the Hadamard product. Using (Cheng2016, Theorem 3.3), the leverage scores of the rows of $\bm{\mathsf{Z}}$ are bounded above by the product of the leverage scores of the rows of its constituent factor matrices:
$\ell_{i}(\bm{\mathsf{Z}})\leq\hat{\ell_{i}}(\bm{\mathsf{Z}}):=\prod_{k\neq j}\ell_{i_{k}}\left(\bm{\mathsf{A}}^{(k)}\right).$ (53)
From (53), we define a probability distribution on the rows of $\bm{\mathsf{Z}}$; namely, row $i$ is selected with probability
$p_{i}=\hat{\ell_{i}}(\bm{\mathsf{Z}})/R^{d-1}.$ (54)
Computing the probabilities in (54) explicitly would require the Kronecker product of leverage scores, costing $O(N^{(-j)})$ operations. However, (Larsen2022, Lemma 9) establishes that each $i_{k}$ in the (d1)(d-1)-tuple corresponding to row $i$ of $\bm{\mathsf{Z}}$ can be sampled independently, using the leverage scores of the rows of $\bm{\mathsf{A}}^{(k)}$, $k\neq j$, which cost $O(N_{k}R^{2})$ to compute. This gives a multi-index corresponding to the $i$-th row of $\bm{\mathsf{Z}}$ without computing the Kronecker product of leverage scores, and $\bm{\mathsf{Z}}(i,:)$ can be formed in O(R(d1))O(R(d-1)) operations using (52). There are additional modifications that can be made to the CP-ARLS-LEV algorithm to further improve performance; we refer the reader to Larsen2022 for details. The basic pseudocode for this procedure is given in Algorithm 8.
Algorithm 8 Alternating Randomized Least Squares for CP with Leverage Scores (CP-ARLS-LEV)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, number of rank-one components $R$, number of samples $S$
2:CP decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
3:Initialize $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R}$, $j=1,\ldots,d$ $\triangleright$ e.g., $\bm{\mathsf{A}}^{(j)}=\texttt{randn}(N_{j},R)$
4:for $j=1,\ldots,d$ do
5:  Initialize $\bm{\mathsf{p}}_{j}=\bm{\mathsf{\ell}}(\bm{\mathsf{A}}^{(j)})/R$, the vector of row leverage scores for $\bm{\mathsf{A}}^{(j)}$
6:end for
7:while not converged do
8:  for $j=1,\ldots,d$ do
9:   $I=\emptyset$
10:   while $|I|<S$ do
11:     for $k=1,\ldots,j-1,j+1,\ldots,d$ do
12:      Sample $i_{k}\sim\bm{\mathsf{p}}_{k}$ from $N_{k}]$
13:     end for
14:     Add $i=(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$ to II
15:   end while
16:   Form submatrix $\widehat{\bm{\mathsf{Z}}}=\bm{\mathsf{Z}}(I,:)$
17:   Form submatrix $\widehat{\bm{\mathsf{X}}}_{(j)}^{*}=\bm{\mathsf{X}}_{(j)}^{*}(I,:)$
18:   Compute $\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\|\widehat{\bm{\mathsf{Z}}}\widehat{\bm{\mathsf{A}}}^{*}-\widehat{\bm{\mathsf{X}}}_{(j)}^{*}\|^{2}_{\text{F}}$
19:   Update $\bm{\mathsf{p}}_{j}=\bm{\mathsf{\ell}}(\bm{\mathsf{A}}^{(j)})/R$
20:  end for
21:end while
The CP-ARLS-LEV algorithm enjoys strong theoretical guarantees. Namely, the resulting rank-$R$ CP decomposition is an $\varepsilon$-accurate approximation to the input tensor with high probability when the number of samples $S=O\left(R^{d-1}/\varepsilon\right)$. In particular, CP-ARLS-LEV improves on the sampling complexity required by the KFJLT-based method of Jin2021. A distributed-memory implementation of CP-ARLS-LEV is also developed in Bharadwaj2024, along with a distributed-memory version of an algorithm that performs random walks on a binary tree for fast leverage score estimation instead.
The recent work of Malik2022 improves on the sampling complexity of CP-ARLS-LEV by avoiding exponential dependence on $d$, using the fast leverage score estimation technique of DrineasMagdonMahoneyWoodruff12, but replacing the FJLT with recursive sketching Ahle2020. This strategy yields a sampling distribution closer to the exact leverage score distribution, but performs similarly to CP-ARLS-LEV in practice, and its implementation details are beyond the scope of the current manuscript; we refer the reader to Ahle2020; Malik2022. There has also been success in using stochastic gradient descent on the least squares objective function in (51), cf. KoldaHong2020; Wang2023; Yu2024. Randomized algorithms to compute low-rank Tucker decompositions as a pre-processing step for CP-ALS have also been investigated Erichson_2020; Yang2018; Zhou2014. We discuss several of these algorithms in the next section.
6.2 Randomized Algorithms for Tucker Decompositions with Orthogonal Factors
Because the Tucker decomposition provides a natural framework for data compression, methods based on the HOSVD (Algorithm 6) and HOOI (Algorithm 7) have been utilized in many data science and scientific computing applications. However, the Tucker decomposition suffers from the curse of dimensionality, thus classical algorithms can be prohibitively expensive for large problems. As randomized algorithms, particularly the randomized SVD (Algorithm 3), have become mainstays for fast and reliable matrix compression, it is natural to consider how randomization can be leveraged for fast and reliable tensor compression in the Tucker format. In this section, we survey recent work on randomized algorithms for Tucker decompositions with orthogonal factor matrices (as opposed to interpolating, cf. Sect. 5.3). We begin with an overview of major developments in randomized Tucker decompositions, before detailing two algorithms based on the randomized SVD: the randomized HOSVD and the randomized HOOI.
Randomized algorithms for Tucker decompositions have been investigated in many problem settings, generally involving either randomized sampling or sketching. One of the earliest works on randomized HOSVD and HOOI algorithms is Tsourakakis2010, called the MACH-HOSVD and MACH-HOOI, so named for the work of Achlioptas-McSherry Achlioptas2001 that inspired these randomized sampling algorithms. In MACH-HOSVD and MACH-HOOI, the input tensor is sparsified according to a coin flip for each nonzero entry; with probability $p$, a nonzero entry is kept and reweighted by $1/p$, and with probability 1p1-p, the entry is set to 0. The HOSVD and HOOI algorithms are then performed on the sparsified tensor, outputting a Tucker decomposition that approximates the original tensor well in practice, even for small values of $p$, e.g., $p=0.1$. Leverage score sampling of matrix unfoldings to sparsify the tensor is explored in Perros2015, and Tucker decompositions of sub-sampled tensors are also investigated in Oseledets2008; Vervliet2014BreakingTC. A detailed survey of these methods can be found in AhmadiAsl2021Randomized.
As the randomized rangefinder gained traction in the numerical linear algebra community, many investigations began to explore randomized sketching for fast Tucker decompositions. In Zhou2014, the RandTucker algorithm is presented, which relies on the randomized rangefinder to compute an HOSVD from Gaussian sketches of mode unfoldings. Formal analysis of RandTucker is performed in Erichson_2020, which incorporates power iteration into RandTucker for faster spectral decay; it is noted, though, that uniform sampling instead of Gaussian sketching performs similarly and may be preferable for large problems. Power iteration and randomized sketching are utilized in the related work of CheWeiYan2020; CheWeiYan2021 for randomized HOSVD and HOOI, which build on the rank-adaptive algorithm of CheWei2019 that iteratively computes an orthonormal basis for the range of each mode unfolding with $\varepsilon$-accuracy.
Rank-adaptive versions of the randomized (ST-)HOSVD are also presented in MinsterSaibabaKilmer2020, which apply the adaptive randomized rangefinder to each mode unfolding. In sun2020low, a single-pass algorithm based on the HOSVD is developed that significantly improves on the cost of storing the random sketching matrices, via Khatri-Rao products of random dimension reduction maps (called tensor random projections); details can also be found in TroppStreaming2019. The Sketch-STHOSVD algorithm of Dong2023Sketching is inspired by sun2020low but incorporates power iteration, whereas the Tucker-TS and Tucker-TTMTS algorithms of Malik2018 are single-pass variants of HOOI using TensorSketch pagh2013compressed. Cost-effective sketching for Tucker decompositions is also investigated in BucciRobol2024MultilinearNystrom, which replaces the randomized SVD step in computing approximate bases with generalized Nyström. In the recent work of Hashemi2025, the Randomized Tucker with Single-Mode Sketching (RTSMS) algorithm is introduced as a rank-adaptive method that sequentially builds approximate bases of unfoldings by applying small random sketches one mode at a time, similar to the randomized ST-HOSVD of MinsterSaibabaKilmer2020, which we describe next.
Randomized (ST-)HOSVD
In the HOSVD (Algorithm 6), we compute factor matrices for a Tucker decomposition as the left singular vectors of each mode unfolding. The randomized HOSVD is a natural extension, where we substitute the SVD of each unfolding with the randomized SVD (Algorithm 3). Pseudocode is given in Algorithm 9, this time for the ST-HOSVD to highlight the slight difference in its computation vs. the HOSVD; the R-HOSVD algorithm of MinsterSaibabaKilmer2020 is precisely the HOSVD algorithm, but using the randomized SVD. Note that this algorithm can be made adaptive by using the adaptive randomized rangefinder with a prescribed error tolerance for the low-rank decomposition of each unfolding.
Algorithm 9 Randomized Sequentially Truncated Higher-Order SVD (R-STHOSVD)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, target ranks $R_{1},\ldots,R_{d}$, oversampling parameter $p$
2:Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
3:Initialize $\bm{\mathcal{G}}=\bm{\mathcal{X}}$
4:for $j=1,\ldots,d$ do
5:  Draw Gaussian matrix $\bm{\mathsf{\Omega}}^{(j)}\in\mathbb{R}^{N^{(-j)}\times(R_{j}+p)}$
6:  Form sketch $\bm{\mathsf{Y}}^{(j)}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}^{(j)}$
7:  $\widehat{\bm{\mathsf{U}}},\widehat{\bm{\mathsf{\Sigma}}},\widehat{\bm{\mathsf{V}}}]=\texttt{svd}(\bm{\mathsf{Y}}^{(j)},R_{j})$
8:  Set $\bm{\mathsf{A}}^{(j)}=\widehat{\bm{\mathsf{U}}}$
9:  Update $\bm{\mathsf{G}}_{(j)}=\widehat{\bm{\mathsf{\Sigma}}}\widehat{\bm{\mathsf{V}}}^{\top}$
10:end for
If $\widehat{\bm{\mathcal{X}}}$ is the output of Algorithm 9 (or the randomized HOSVD, with Gaussian sketching), it is proven in (MinsterSaibabaKilmer2020, Theorem 3.1, 3.2) that the expected error is
$\mathbb{E}_{\{\bm{\mathsf{\Omega}}^{(j)}\}_{j=1}^{d}}\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|\leq\left(d+\frac{\sum_{j=1}^{d}R_{j}}{p-1}\right)^{1/2}\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}_{opt}\|.$ (55)
Letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity, the computational cost of R-STHOSVD as in Algorithm 9 is $O(\sum_{j=1}^{d}k^{j}n^{d-j+1}+k^{j}n^{d-j})$. By contrast, R-HOSVD (substituting the SVD with the randomized SVD in Algorithm 6) has computational cost $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$. In other words, the cost is reduced roughly by a factor of $n$ when using randomized (ST-)HOSVD vs. deterministic. Algorithmic performance can also be improved in the R-STHOSVD by processing the largest modes first, and while the error of a given output does depend on processing order, the worst-case expected error does not, cf. (MinsterSaibabaKilmer2020, Sect. 3.3).
Randomized HOOI
Recall from Section 5.3 that higher-order orthogonal iteration (HOOI) is a method of computing Tucker decompositions by iteratively solving least-squares problems for each mode’s factor matrix, which is typically more accurate than the HOSVD but more computationally expensive. Similar to our goal in CP-ALS, for a given input $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, we are seeking a rank-$(R_{1},\ldots,R_{d})$ Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}$ that minimizes
$\min_{\bm{\mathcal{G}},\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}}\|\bm{\mathcal{X}}-\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}\|,$ (56)
where $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$ and $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$. Through an appropriate transformation of the objective function, we can re-write the HOOI procedure in Algorithm 7 as two main steps that are repeated until convergence; cf. Malik2018:
  1. 1.
    For $j=1,\ldots,d$, solve $\bm{\mathsf{A}}^{(j)}=\underset{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}{\arg\min}\|\bm{\mathsf{Z}}^{(j)}\bm{\mathsf{G}}_{(j)}^{\top}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{X}}_{(j)}^{\top}\|^{2}$, where
    $\bm{\mathsf{Z}}^{(j)}=\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)},$
  2. 2.
    Update $\bm{\mathcal{G}}=\underset{\bm{\mathcal{Y}}\in\mathbb{F}^{R_{1}\times R_{d}}}{\arg\min}\|\bm{\mathsf{Z}}\bm{\mathsf{y}}-\bm{\mathsf{x}}\|^{2}$, where $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(d)}\otimes\cdots\bm{\mathsf{A}}^{(1)}$, and $\bm{\mathsf{y}},\bm{\mathsf{x}}$ are (column) vectorizations of $\bm{\mathcal{Y}}$, $\bm{\mathcal{X}}$, respectively.
We observe that each of these two steps involve large overdetermined least squares problems, and similar to the strategy in CP-ARLS-LEV, we will exploit the matrix product structures to solve smaller sketched least squares problems efficiently, now with TensorSketch, cf. (Malik2018, Algorithm 2 (TUCKER-TS)).
TensorSketch   We briefly summarize TensorSketch pagh2013compressed; Pham2013; AvronNguyenWoodruff2014; DiaoSongSunWoodruff2018, a specialized version of the CountSketch operator of Charikar2004. A CountSketch operator can be defined as a linear map $\bm{\mathsf{S}}:\mathbb{R}^{I}\rightarrow\mathbb{R}^{J}$ given by $\bm{\mathsf{S}}=\bm{\mathsf{P}}\bm{\mathsf{D}}$, where $\bm{\mathsf{P}}\in\mathbb{R}^{J\times I}$ is a matrix with $\bm{\mathsf{P}}(h(i),i)=1$ and all other entries equal to 0, for a random map $h:[I]\rightarrow[J]$ such that $\mathbb{P}[h(i)=j]=1/J$ for all $i\in[I]$, $j\in[J]$. The matrix $\bm{\mathsf{D}}\in\mathbb{R}^{I\times I}$ is a diagonal matrix with each entry equal to $\pm 1$ with equal probability.
In particular, a CountSketch operator $\bm{\mathsf{S}}$ can be rapidly applied to a matrix $\bm{\mathsf{A}}$ in $O(\textup{nnz}(\bm{\mathsf{A}}))$ operations, without explicitly forming $\bm{\mathsf{S}}$ or $\bm{\mathsf{A}}$. More specifically, if $\bm{\mathsf{S}}\in\mathbb{R}^{J\times I}$ and $\bm{\mathsf{A}}\in\mathbb{F}^{I\times L}$, typically with $J\ll I$, then we can form $\bm{\mathsf{Y}}=\bm{\mathsf{S}}\bm{\mathsf{A}}$ by hashing each row $i$ of $\bm{\mathsf{A}}$ with an integer $h$ sampled uniformly from $[J]$, assigning it a value $s$ of $\pm 1$ with equal probability, then summing the rows with the same hash value (i.e. adding $s\bm{\mathsf{A}}_{i,:}$ to the $h$th row of $\bm{\mathsf{Y}}$ if $\bm{\mathsf{A}}_{i,:}$ has hash value $h$).
A TensorSketch operator is a specific type of CountSketch operator, which can be rapidly applied to Kronecker products, e.g., $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)}\in\mathbb{F}^{n^{d}\times k^{d}}$. (We again assume here that $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$.) If $\bm{\mathsf{T}}:\mathbb{R}^{n^{d}}\rightarrow\mathbb{R}^{\ell}$ is a TensorSketch operator with embedding dimension $\ell\ll n^{d}$, then the cost of forming $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is only $O(\ell dnk+\ell k^{d})$, suppressing log factors, instead of $O(\ell n^{d}k^{d})$ for naive multiplication. This is accomplished by sketching each factor matrix with independent CountSketch operators, then convolving with Fast Fourier Transforms (FFTs).
To that end, let $h_{j}:[n]\rightarrow[\ell]$ for $j=1,\ldots,d$ be 3-wise independent hash functions, i.e. for any distinct $i_{1},i_{2},i_{3}\in[n]$, the hash codes $h_{j}(i_{1})$, $h_{j}(i_{2})$, $h_{j}(i_{3})$ are independent random variables, and $h_{j}(i)$ is uniformly distributed in $\ell$] for any fixed $i\in[n]$. Let $s_{j}:[n]\rightarrow\{+1,-1\}$ be 4-wise independent sign functions (so that analogously, $s_{j}(i_{1})$, $s_{j}(i_{2})$, $s_{j}(i_{3})$, $s_{j}(i_{4})$ are independent random variables with $s_{j}(i)\sim\textup{Uniform}\{\pm 1\}$. We form independent CountSketch matrices $\bm{\mathsf{S}}^{(j)}=\bm{\mathsf{P}}^{(j)}\bm{\mathsf{D}}^{(j)}$ for $j=1,\ldots,d$, using the hash functions $h_{j}$ to form $\bm{\mathsf{P}}^{(j)}$ and $s_{j}(i)=\bm{\mathsf{D}}^{(j)}_{i,i}$ for $i\in[n]$. The application of $\bm{\mathsf{S}}^{(j)}$ to each column $\bm{\mathsf{A}}^{(j)}_{:,r_{j}}$, for $r_{j}=1,\ldots,k$, can be represented by an $(\ell-1)$-degree polynomial
$\mathcal{P}^{(j)}_{r_{j}}(\omega)=\sum_{i=1}^{n}s_{j}(i)\bm{\mathsf{A}}^{(j)}_{i,r_{j}}\omega^{h_{j}(i)-1}=:\sum_{l=1}^{\ell}c^{(j)}_{l,r_{j}}\omega^{j-1},$ (57)
and we define $\bm{\mathsf{c}}_{r_{j}}^{(j)}$ to be the vector of coefficients $(c^{(j)}_{1,r_{j}},\ldots,c^{(j)}_{\ell,r_{j}})$, which are grouped in (57) by hash value.
Let $\bm{\mathsf{T}}$ be the TensorSketch operator, which is defined as the CountSketch operator formed from hash function $H$ and sign function $S$ given by
$H:[n]^{d}\rightarrow\ell,\ \ (i_{1},\ldots,i_{d})\mapsto\left(\sum_{j=1}^{d}(h_{j}(i_{j})-1)\mod\ell\right)+1,$ (58)
$S:[n]^{d}\rightarrow\{+1,-1\},\ \ (i_{1},\ldots,i_{d})\mapsto\prod_{j=1}^{d}s_{j}(i_{j}).$ (59)
Then the $r$th column of $\bm{\mathsf{T}}(\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)})=\bm{\mathsf{T}}\bm{\mathsf{Z}}$ for $r=1,\ldots,k^{d}$ can also be represented as a polynomial
$\mathcal{P}_{r}(\omega)$ $=\sum_{i=1}^{n^{d}}S(i_{1},\ldots,i_{d})\bm{\mathsf{Z}}_{i,r}\omega^{H(i_{1},\ldots,i_{d})},$ (60)
$=\sum_{i=1}^{n^{d}}s_{1}(i_{1})\cdots s_{d}(i_{d})\bm{\mathsf{A}}^{(1)}_{i_{1},r_{1}}\cdots\bm{\mathsf{A}}^{(d)}_{i_{d},r_{d}}\omega^{(h_{1}(i_{1})+\cdots+h_{d}(i_{d})-d)\mod\ell}$ (61)
$=\textup{FFT}^{-1}\left(\textup{FFT}(\bm{\mathsf{c}}^{(1)}_{r_{1}})*\cdots*\textup{FFT}(\bm{\mathsf{c}}_{r_{d}}^{(d)}\right),$ (62)
where * denotes the Hadamard product, and $i$ in (60) corresponds to the tuple $(i_{1},\ldots,i_{d})\in[n]^{d}$, and similarly $r$ corresponds to the tuple $(r_{1},\ldots,r_{d})\in[k]^{d}$. In other words,
$\bm{\mathsf{T}}\bm{\mathsf{Z}}=\textup{FFT}^{-1}\left(\left({\bigodot}_{j=1}^{d}\left(\textup{FFT}\left(\bm{\mathsf{S}}^{(j)}\bm{\mathsf{A}}^{(j)}\right)\right)^{\top}\right)^{\top}\right),$ (63)
where $\odot$ is the KRP, and the FFT is applied column-wise. Instead of the usual $O(\ell n^{d}k^{d})$ cost of matrix multiplication, the cost of $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is $O(\ell dnk+\ell k^{d}\log(\ell))$ using (63).
Using TensorSketch in HOOI We now turn our attention back to the randomized HOOI algorithm of Malik2018, which uses TensorSketch to reduce the size of each least squares subproblem in the two primary steps of HOOI listed above. The least squares problem to find the factor matrix for each mode $j=1,\ldots,d$ is highly overdetermined, involving a Kronecker product $\bm{\mathsf{Z}}^{(j)}$ of size $N^{(-j)}\times(\prod_{k\neq j}R_{k})$. The least squares problem to find the core tensor is also highly overdetermined, involving a Kronecker product of size $\prod_{j=1}^{d}N_{j}\times\prod_{j=1}^{d}R_{j}$, or $n^{d}\times k^{d}$ for convenience. To compute the factor matrices and core tensor, we form independent TensorSketches to solve smaller sketched least squares problems corresponding to both of these steps. Pseudocode for R-HOOI, equivalently (Malik2018, TUCKER-TS), is given in Algorithm 10. This algorithm can be implemented as a single-pass algorithm through the input tensor, cf. (Malik2018, Algorithm S3).
The dominant cost of Algorithm 10 (single-pass) is $O(dnk^{2}+\ell dk^{d}+\ell_{c}k^{d}+k^{2d})$, or roughly $k^{O(d)}$, versus the dominant cost of $O(n^{d})$ for deterministic HOOI (Algorithm 7), if we suppress $\log$ factors and assume sketching dimension $\ell=\ell_{1}=\cdots=\ell_{d}$. We refer the reader to (Malik2018, S3.2.6) for details on the complexity analysis. However, we observe that the single-pass R-HOOI algorithm as presented in Malik2018 uses the same TensorSketches for sketching the updated Kronecker products until some convergence criteria is met. In practice, the performance of R-HOOI as written in Algorithm 10 is curiously better in terms of error reduction than if new independent TensorSketch operators were drawn in every iteration, though it is unclear why this is the case. However, for the theoretical guarantees of DiaoSongSunWoodruff2018 to hold for TensorSketched least squares, it is technically necessary to form new independent TensorSketch operators in each iteration of R-HOOI. In Malik2018, the numerical results suggest that sketching dimensions $\ell=Ck^{d-1}$ and $\ell_{c}=Cr^{d}$ for constant $C>4$ (e.g., $C=10$) work well in practice.
Algorithm 10 Randomized Higher-order Orthogonal Iteration with TensorSketch (R-HOOI)
1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, factor matrix target ranks $R_{1},\ldots,R_{d}$, sketching dimensions $\ell_{1},\ldots,\ell_{d}$, $\ell_{c}$
2:Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
3:Initialize $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$, $j=2,\ldots,d$
4:Draw TensorSketch matrices $\bm{\mathsf{T}}^{(j)}\in\mathbb{R}^{\ell_{j}\times(\prod_{i\neq j}N_{i})}$ for $i,j\in[d]$ and $\bm{\mathsf{T}}^{(c)}\in\mathbb{R}^{\ell_{c}\times\prod_{j}N_{j}}$
5:while not converged do
6:  for $j=1,\ldots,d$ do
7:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{T}}^{(j)}\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{T}}^{(j)}\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}$
8:  end for
9:  $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bm{\mathsf{T}}^{(c)}\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{T}}^{(c)}\bm{\mathsf{x}}\right\|^{2}_{2}$ $\triangleright$ Column vectorization $\bm{\mathsf{x}}$ of $\bm{\mathcal{X}}$
10:  Re-fold vectorization $\bm{\mathsf{g}}$ into $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$
11:end while
Randomized TSMS
We end this section by briefly highlighting the recent work of Hashemi2025 on Randomized Tucker decompositions with Single-Mode Sketching (RTSMS). The RTSMS algorithm sketches the input tensor one mode at a time, i.e. sketching on the left of the mode-$j$ unfolding vs. sketching on the right, to avoid the computational bottleneck of multiplying along the larger dimension $N^{(-j)}=\prod_{i\neq j}N_{i}$. The algorithm is also single-pass, so that the input tensor is not revisited after the first sketch, and rank-adaptive if the target ranks for the factor matrices are unknown a priori.
To illustrate, suppose we want to compute a rank-$(R_{1},R_{2},R_{3})$ HOSVD that approximates an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$. (We omit rank-adaptivity and iterative refinement; the reader is referred to Hashemi2025 for practical implementation details.) We initialize tensors
$\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{X}}\ \textup{and}\ \bm{\mathcal{G}}^{\textup{new}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}^{(1)}\iff\bm{\mathsf{G}}^{\textup{new}}_{(1)}=\left(\bm{\mathsf{\Omega}}^{(1)}\right)^{\top}\bm{\mathsf{X}}_{(1)}\in\mathbb{F}^{\tilde{R_{1}}\times(N_{2}N_{3})}$
for some sketching dimension $\tilde{R_{1}}>R_{1}$. There exists a factor matrix $\bm{\mathsf{A}}^{(1)}\in\mathbb{F}^{N_{1}\times\tilde{R_{1}}}$ such that $\bm{\mathcal{G}}^{\textup{new}}\approx\bm{\mathcal{G}}^{\textup{old}}\times_{1}\bm{\mathsf{A}}^{(1)}$. In practice, this is computed as an approximate solution of the corresponding least squares problem, which is being sub-sampled from the right, cf. (Hashemi2025, Section 4.2). We then update $\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{G}}^{\textup{new}}$.
For the next mode, we first update $\bm{\mathcal{G}}^{\textup{new}}$ via mode-2 sketching:
$\bm{\mathsf{G}}^{\textup{new}}_{(2)}=\left(\bm{\mathsf{\Omega}}^{(2)}\right)^{\top}\bm{\mathsf{G}}_{(2)}^{\textup{old}}\in\mathbb{F}^{\tilde{R_{2}}\times(\tilde{R_{1}}N_{3})}$
for some sketching dimension $\tilde{R_{2}}>R_{2}$. To compute $\bm{\mathsf{A}}^{(2)}$, we solve a smaller sketched least squares problem of size $\tilde{R_{2}}\times(\tilde{R_{1}}N_{3})$, and repeat this for the final mode.
The computational complexity of RTSMS is dominated by the first TTM $\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}_{1}$. Letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the dominant cost is $O(kn^{d})$ to compute the first TTM, but the sketch size is only $k\times n$ vs. $k\times n^{d-1}$ for R-HOSVD. The least squares solves are also significantly reduced by sub-sampling and by computing a generalized Nyström decomposition vs. SVD for the factor matrices. The full tensor does not need to be sketched again in modes $j=2,\ldots,d$, and the core tensor is available immediately upon termination, as in the ST-HOSVD algorithm. Additional details can be found in Hashemi2025.
6.3 Randomized Algorithms for Tensor ID/CUR Decompositions
As we observed in Sect. 5.3, the tensor ID/CUR decomposition has been formulated in many ways in the literature, with no one standard representation. However, there exists a natural categorization of tensor ID/CUR decompositions, based on whether the factor matrices or core tensors in the computed decompositions are comprised of entries of the input tensor. In this section, we explore how randomization can be leveraged in prototypical algorithms to compute a Tucker ID/CUR decomposition from each category. We also discuss recent work on randomized algorithms for tensor ID/CUR in the CP format, which can be viewed as a form of input tensor pruning.
Randomized Algorithms for Tucker Core ID/CUR
We first consider randomized algorithms for tensor ID/CUR factorizations in the Tucker format of the first category, where the core tensor in the compressed representation is a sub-tensor of the input tensor. In particular, we focus on the tensor ID/CUR decompositions that were recently characterized in Cai2024; Cai2021, termed the “Chidori” and “Fiber” tensor CUR factorizations. This work was inspired by numerous investigations of randomized sampling algorithms for tensor CUR decompositions, including CaiafaCichocki2009; Caiafa2010; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR.
In Cai2021, given an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the rank-$(R_{1},\ldots,R_{d})$ tensor CUR factorization is defined as
$\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}(\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}^{\dagger})\times_{2}(\bm{\mathsf{C}}_{2}\bm{\mathsf{U}}_{2}^{\dagger})\times\cdots\times_{d}(\bm{\mathsf{C}}_{d}\bm{\mathsf{U}}_{d}^{\dagger}),$ (64)
where the rank-$(R_{1},\ldots,R_{d})$ core tensor is given by
$\bm{\mathcal{G}}=\bm{\mathcal{X}}(I_{1},\ldots,I_{d}),$ (65)
for index sets $I_{j}\subset[N_{j}]$, $j=1,\ldots,d$. The mode-$j$ factor matrix $\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger}$ is defined in terms of the mode-$j$ unfolding:
$\begin{split}\bm{\mathsf{C}}_{j}&=\bm{\mathsf{X}}_{(j)}(:,J_{j}),\ \ \textup{and}\\ \bm{\mathsf{U}}_{j}&=\bm{\mathsf{C}}_{j}(I_{j},:),\end{split}$ (66)
where $J_{j}\subset[N^{(-j)}]$. If indices $J_{j}$ are computed independently from indices $I_{j}$, the decomposition in (64) is called a “Fiber” CUR decomposition in Cai2021. If $J_{j}=\otimes_{k\neq j}I_{k}$, then (64) is called a “Chidori” CUR decomposition. By (Cai2021, Theorem 3.3), (HammHuang2020, Corollary 5.2), if the indices $I_{j}$ and $J_{j}$ are each sampled uniformly with $|I_{j}|=O(R_{j}\log N_{j})$ and $|J_{j}|=O(R_{j}\log N^{(-j)})$, then the computed Tucker decomposition is a rank-$(R_{1},\ldots,R_{d})$ approximation with high probability. We present pseudocode for either the Chidori or Fiber tensor CUR decomposition in Algorithm 11.
With uniform sampling, the computational complexity of computing either tensor CUR decomposition is dominated by the cost of the pseudoinverse of $\bm{\mathsf{U}}_{j}$, which is $O(k^{d+1}\log^{d}(n))$ for Chidori and $O(k^{2}\log^{2}(n))$ for Fiber, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$ as before. Recall that the HOSVD has complexity $O(kn^{d})$ by comparison.
Algorithm 11 Randomized Tucker ID/CUR with interpolating core
[…]
2:chidori or fiber decomposition $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{j=1}^{d}(\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger})$
3:for $j=1,\ldots,d$ do
4:  if chidori then
5:   Sample indices $I_{j}$ from $N_{j}]$, $|I_{j}|=S_{j}$ without replacement according to $\bm{\mathsf{p}}_{j}$
6:   Set $J_{j}=\otimes_{i\neq j}I_{i}$
7:  end if
8:  if fiber then
9:   Sample indices $I_{j}$ from $N_{j}]$, $|I_{j}|=S_{j}$ without replacement according to $\bm{\mathsf{p}}_{j}$
10:   Sample indices $J_{j}$ from $N^{(-j)}]$, $|J_{j}|=T_{j}$ without replacement according to $\bm{\mathsf{q}}_{j}$
11:  end if
12:  Form $\bm{\mathsf{C}}_{j}=\bm{\mathsf{X}}_{(j)}(:,J_{j})$
13:  Form $\bm{\mathsf{U}}_{j}=\bm{\mathsf{C}}_{j}(I_{j},:)$
14:end for
[…]
The recent work of Zhang2025 proposes a similar factorization to (64), called the “CoreID.” The CoreID algorithm is reminiscent of the sequentially-truncated HOSVD, specifically the “structure-preserving” ST-HOSVD developed for sparse input tensors in MinsterSaibabaKilmer2020. Namely, in the CoreID, as each mode’s factor matrix is computed, the input tensor is compressed along that mode before proceeding to the next one, so that each subsequent unfolding is reduced in size.
Through matricization, computing the CoreID is reduced to computing a randomized matrix row ID in each mode, cf. Algorithm 4. For $j=1,\ldots,d$, the mode-$j$ unfolding of the tensor is sketched by a random matrix $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}_{j}.$ Then a row ID of $\bm{\mathsf{Y}}_{j}$ is computed by some skeleton selection method, e.g., greedy pivoting or squared-norm sampling. The selected skeletons $I_{j}$ form the core sub-tensor, and the mode-$j$ factor matrix for the CoreID is defined as the associated interpolation matrix. The tensor is then compressed along mode $j$ as in ST-HOSVD before proceeding to the next mode. Sketching is the dominating cost in this algorithm, which is why many of the numerical investigations in Zhang2025 consider structured input tensors for fast sketching, e.g., a sparse input tensor or a CP-factorized input tensor.
Randomized Algorithms for Tucker Factor ID/CUR
We now turn our attention to randomized algorithms for tensor ID/CUR factorizations in the Tucker format in which the factor matrices contain entries of the input tensor. We illustrate such factorizations with the “higher-order ID” of Saibaba2016, which we contrast with the “SatID” of Zhang2025.
For a given $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the higher-order ID (HOID) of Saibaba2016 is given by
$\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{C}}_{1}\times_{2}\bm{\mathsf{C}}_{2}\times\cdots\times_{d}\bm{\mathsf{C}}_{d},$ (67)
where the rank-$(R_{1},\ldots,R_{d})$ core tensor $\bm{\mathcal{G}}$ is defined as
$\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{C}}_{1}^{\dagger}\times_{2}\bm{\mathsf{C}}_{2}^{\dagger}\times\cdots\times_{d}\bm{\mathsf{C}}_{d}^{\dagger},$ (68)
and each factor matrix $\bm{\mathsf{C}}_{j}\in\mathbb{F}^{N_{j}\times R_{j}}$ consists of column skeletons of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$. The core tensor in (68) is optimal for the factorization in (67), cf. DeLathauwer2000. For the theoretical guarantees of Saibaba2016, the strong rank-revealing QR algorithm of gu1996 is used for skeleton selection on a random sketch $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{\Omega}}_{j}\bm{\mathsf{X}}_{(j)}$, where $\bm{\mathsf{\Omega}}_{j}$ is an $(R_{j}+p)\times N_{j}$ random matrix for some small oversampling parameter $p$. However, in practice, CPQR or LUPP in the randomized matrix ID (e.g., Algorithm 4, with input $\bm{\mathsf{X}}_{(j)}^{*}$) performs similarly and is less computationally expensive. Pseudocode is provided in Algorithm 12.
Algorithm 12 Randomized Tucker ID/CUR with interpolating factors
[…]
2:Tensor ID $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{j=1}^{d}\bm{\mathsf{C}}_{j}$
3:for $j=1,\ldots,d$ do
4:  Compute rank-$R_{j}$ randomized column ID $\bm{\mathsf{X}}_{(j)}\approx\bm{\mathsf{C}}_{j}\bm{\mathsf{W}}_{j}$ $\triangleright$ e.g., Algorithm 4 on $\bm{\mathsf{X}}_{(j)}^{*}$
5:end for
6:Compute core $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{j=1}^{d}\bm{\mathsf{C}}_{j}^{\dagger}$
For each mode $j$, the cost of sketching is $O(R_{j}N_{j}N^{(-j)})=O(R_{j}N)$. Skeleton selection as in Saibaba2016 costs $O(R_{j}^{2}N^{(-j)})$. The cost of computing the core tensor is $O\left(\sum_{j=1}^{d}\left[R_{j}^{2}N_{j}+\prod_{i\leq j}R_{i}\cdot N^{(-j)}\right]\right)$. Letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$ again, the total cost is roughly $O(dkn^{d}+dk^{d-1}n^{d-1})$. A sequentially-truncated version of the HOID (ST-HOID), analogous to the ST-HOSVD, is given in (Saibaba2016, Algorithm 4).
A very similar decomposition is developed in Zhang2025, termed the “SatID,” which can be more efficiently computed. In particular, for an $n^{d}$-sized tensor, a QR-based algorithm to compute skeletons for the mode-$j$ factor matrix as in Saibaba2016 requires $n^{d-1}$ column norms, which is not optimal for structured tensors and can even have larger storage costs than the input tensor. By contrast, the SatID algorithm is based on a randomized matrix ID computed by marginalized norm sampling, which can be viewed as an extension of the original tensor CUR construction of Drineas2007tensorSVD, which uses norm sampling.
We briefly summarize the marginalized norm sampling procedure of Zhang2025. As in Sect. 6.1, for a fixed mode $j$, we can uniquely identify a column index $i\in[N^{(-j)}]$ of $\bm{\mathsf{X}}_{(j)}$ with a (d1d-1)-tuple $(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$. Similar to CP-ARLS-LEV in Larsen2022, the idea is to accelerate sampling of $i\in[N^{(-j)}]$, now based on column norm sampling of $\bm{\mathsf{X}}_{(j)}$.
Suppose we have already selected some subset of column indices $S$ of $\bm{\mathsf{X}}=\bm{\mathsf{X}}_{(j)}$ to determine the mode-$j$ column skeletons comprising $\bm{\mathsf{C}}_{j}$, and we want to select the next column. Norm sampling (cf. Sect. 3.2) requires the computation of column scores
$d_{i}^{(S)}=\|\bm{\mathsf{X}}_{:,i}-\bm{\mathsf{Q}}_{S}\bm{\mathsf{Q}}_{S}^{\top}\bm{\mathsf{X}}_{:,i}\|^{2}=\min_{\bm{\mathsf{v}}}\|\bm{\mathsf{X}}_{:,S}\bm{\mathsf{v}}-\bm{\mathsf{X}}_{:,i}\|^{2},\ \ i\in[N^{(-j)}],$ (69)
where $\bm{\mathsf{Q}}_{S}$ is an orthonormal basis for the range of $\bm{\mathsf{X}}_{:,S}$. If $X=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{d})$ is a random variable representing the next column index, then using (69),
$\mathbb{P}(X=i)\propto d_{i}^{(S)}.$
The key observation is that
$\mathbb{P}(X_{1}=i_{1})\propto\sum_{i_{2},\ldots,i_{d}}d_{i_{1},i_{2},\ldots,i_{d}}^{(S)},$
so that we can sample $i_{1}$ from its marginal distribution, then sample $i_{2}$ given $i_{1}$, and continue for all of $i=(i_{1},i_{2},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$. This sum can be computed efficiently via the equation derived in (Zhang2025, Eq. 18):
$\mathbb{P}(X_{1}=i_{1})\propto\left\|\widetilde{\bm{\mathsf{X}}}_{:,i_{1}}\right\|^{2},$ (70)
[…]
For structured inputs, such as sparse tensors, this procedure can be modified for better performance. The total complexity (including randomized sketching) is $O(\textup{nnz}(\bm{\mathcal{X}})+\ell n^{2}k^{2})$ for a rank-$(k,\ldots,k)$ SatID of an $n^{d}$-sized tensor with sketch size $\ell$; see (Zhang2025, Algorithm 7) for details.
Randomized Algorithms for CP Factor ID/CUR
We now consider tensor ID/CUR decompositions in the CP format, which necessarily fall into the category of decompositions in which the factor matrices contain entries of the input tensor. These decompositions can be interpreted as the result of pruning the input tensor, leaving only its most important rank-one components.
The seminal work of Biagioni2015 introduces a method to obtain rank-$k$ CP-tensor IDs from rank-$R$ CP-factorized tensors, given by
$\bm{\mathcal{X}}=\sum_{r=1}^{R}\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}.$ (71)
To better illustrate the algorithm, we express (71) in a vectorized form, where $\bm{\mathsf{X}}_{r}\in\mathbb{F}^{N\times 1}$ is a vectorization of the rank-one component $\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}$. Then we can consider the matrix
$\bm{\mathsf{X}}=\begin{bmatrix}\bm{\mathsf{X}}_{1}\ \vline&\cdots&\vline\ \bm{\mathsf{X}}_{R}\end{bmatrix}\in\mathbb{F}^{N\times R},$ (72)
and within this framework, we are seeking a rank-$k$ matrix column ID of $\bm{\mathsf{X}}$, corresponding to a rank-$k$ CP decomposition $\widehat{\bm{\mathcal{X}}}$ denoted by
$\widehat{\bm{\mathcal{X}}}=\sum_{r=1}^{k}\bm{\mathsf{X}}_{s_{r}},$ (73)
for $k$ column skeletons of (72) indexed by $\{s_{1},\ldots,s_{k}\}\subset[R]$. In particular, we want to select these $k$ indices without explicitly forming the matrix $\bm{\mathsf{X}}$.
To do this, the randomized CP compression algorithm of Biagioni2015 uses rank-one random tensors $\bm{\mathsf{r}}_{r}^{(1)}\circ\cdots\circ\bm{\mathsf{r}}_{r}^{(d)}$, with vectorized form $\bm{\mathsf{R}}_{r}$ for $r=1,\ldots,\ell$ and $\ell=k+p$ for some small oversampling parameter $p$. An $\ell\times R$ random sample matrix is formed via
$\bm{\mathsf{Y}}=\begin{bmatrix}\langle\bm{\mathsf{R}}_{1},\bm{\mathsf{X}}_{1}\rangle&\langle\bm{\mathsf{R}}_{1},\bm{\mathsf{X}}_{2}\rangle&\cdots&\langle\bm{\mathsf{R}}_{1},\bm{\mathsf{X}}_{R}\rangle\\ \vdots&\vdots&&\vdots\\ \langle\bm{\mathsf{R}}_{\ell},\bm{\mathsf{X}}_{1}\rangle&\langle\bm{\mathsf{R}}_{\ell},\bm{\mathsf{X}}_{2}\rangle&\cdots&\langle\bm{\mathsf{R}}_{\ell},\bm{\mathsf{X}}_{R}\rangle\end{bmatrix}=:\bm{\mathsf{R}}\bm{\mathsf{X}},$ (74)
whose rank-$k$ column ID is then computed to obtain skeletons $\{s_{1},\ldots,s_{k}\}$ as in (73). We note that the matrix $\bm{\mathsf{Y}}$ can be computed efficiently without ever explicitly forming $\bm{\mathsf{R}}$ or $\bm{\mathsf{X}}$, since each $\bm{\mathsf{Y}}(l,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}^{(j)}_{l},\bm{\mathsf{a}}^{(j)}_{r}\rangle$. However, we note that the elements of $\bm{\mathsf{R}}$ are not independent, and so theoretical guarantees on the randomized matrix decomposition computed with $\bm{\mathsf{Y}}$ are not given, though it is demonstrated in Biagioni2015 that it performs well in practice. Pseudocode is provided in Algorithm 13.
Algorithm 13 Randomized CP ID/CUR with interpolating factors
1:CP tensor $\bm{\mathcal{X}}=\sum_{r=1}^{R}\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}$, target rank $k$, oversampling $p$
2:Rank-$k$ CP tensor $\widehat{\bm{\mathcal{X}}}=\sum_{i=1}^{k}\bm{\mathsf{a}}^{(1)}_{s_{i}}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{s_{i}}$
3:Draw $\ell=k+p$ random rank-one tensors $\bm{\mathsf{r}}_{i}^{(1)}\circ\cdots\circ\bm{\mathsf{r}}_{i}^{(d)}$
4:Form $\bm{\mathsf{Y}}\in\mathbb{F}^{\ell\times R}$ via $\bm{\mathsf{Y}}(\ell,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}_{\ell}^{(j)},\bm{\mathsf{a}}_{r}^{(j)}\rangle$
5:Compute rank-$k$ column ID of $\bm{\mathsf{Y}}$ for skeletons $\{s_{1},\ldots,s_{k}\}$ $\triangleright$ e.g., Algorithm 4 applied to $\bm{\mathsf{Y}}^{*}$
To analyze the cost of this algorithm, let $n=N_{1}=\cdots=N_{d}$. Forming the $\ell$ random tensors costs $O(t_{R}\ell dn)$, where $t_{R}$ is the cost of generating a single random number. The cost of computing each entry of $\bm{\mathsf{Y}}$ is $O(dn)$, thus the total cost to compute $\bm{\mathsf{Y}}$ is $O(\ell R\cdot dn)$. The rank-$k$ matrix ID of $\bm{\mathsf{Y}}$ comes at a cost of $O(k\ell R)$, and forming the rank-$k$ tensor ID from the rank-one terms of $\bm{\mathcal{X}}$ requires $O(dkn)$ additional operations. In particular, the method of Biagioni2015 is faster than a rank-$k$ ALS procedure by a factor of $k\cdot n_{iter}$, where $n_{iter}$ is the number of iterations of ALS.
The recent work of MalikBecker2020 constructs an identical rank-$k$ CP approximation to a rank-$R$ CP-factorized tensor $\bm{\mathcal{X}}$, but uses the TensorSketch operator instead of rank-one random tensors. The complexity of the TensorSketch-based algorithm of MalikBecker2020 is dominated by the sketching step, which costs $O(d(nR+R\ell\log\ell))$; the ID of the $\ell\times R$ matrix $\bm{\mathsf{Y}}$ costs the same as above. In addition to its theoretical performance guarantees, the algorithm’s practical performance is demonstrated through numerical experiments, in which the TensorSketch-based method exhibits a runtime speed-up of roughly one order of magnitude over other randomized CP-tensor ID algorithms, including the random projection method of Biagioni2015.
7 Concluding Remarks
We end our survey with an outlook on the future of RNLA, which will continue to be an indispensable tool for scientific computing and data science in the face of increasingly large problems. We have seen throughout this work that randomization is not only useful for dimension reduction, but also admits algorithmic implementations that can be optimized for specific hardware or problem constraints, such as parallelization for GPU computing or factorizing streaming data. Randomization is especially critical for the development of scalable tensor algorithms.
Future work may touch on several different aspects of computing fast and accurate matrix and tensor decompositions with randomized algorithms. Probabilistic guarantees are still needed for the resulting error of many randomized algorithms that are used in practice, as well as tighter bounds that more closely align with observed behavior. For example, it is generally an open question as to what the weakest assumptions are for a particular type of randomized DRM to achieve a desired approximation accuracy. The question can be further refined by the specific problem setting, e.g., the weakest assumptions to obtain a certain accuracy of rank-structured matrix decomposition or sketched least squares solutions.
As stronger theoretical foundations are developed for new and existing randomized algorithms, high-performance software is also needed to fully leverage available computing resources. For example, there remains a need for standardized RNLA kernels and interfaces that can offer both practical speed and front-end accessibility, e.g., in Python or Julia. RNLA software libraries like RandBLAS or RandLAPACK are continually being updated with new functionality, such as mixed-precision arithmetic or sparse linear algebra packages. Additionally, there is an ongoing effort to develop high-performance implementations of randomized tensor algorithms that take advantage of GPU acceleration. This includes the growing body of work on tensor decompositions other than those presented here, some of which can break the curse of dimensionality.
In summary, RNLA has become an established area of applied mathematics that will prove increasingly important as larger problems are tackled and the size and complexity of scientific data grow. As such, there remain many interesting open questions to investigate, which build on the fundamental concepts surveyed in this work.

### v2

@@
 1 Introduction
-Numerical linear algebra (NLA) is a cornerstone of applied mathematics, scientific computing, and data science, with nascent algorithmic implementations dating back to the beginning of recorded history.
-Algorithms to accomplish NLA tasks, such as solving linear systems, computing eigenvalues and singular values, orthogonalizing and factorizing matrices, among many others, have been used in applications ranging from ancient Babylonian land surveying (Brezinski2022_NLA, Chapter 2) to Nobel Prize-winning work on computer vision AlexNet_2012.
+Numerical linear algebra (NLA) is a cornerstone of applied mathematics, scientific computing, and data science.
 Throughout its long history, NLA algorithmic development has been driven by two competing imperatives, accuracy and efficiency, with the extant challenge to strike the right balance between them.
-Origins of RNLA   The relatively recent emergence of randomized algorithms marks a major paradigm shift in NLA.
-Though Monte Carlo methods developed by Ulam and von Neumann as part of the Manhattan Project were made public in the late 1940s and 1950s Metropolis1987; Ulam1949; vonNeumann1951techniques, for decades, the NLA community was largely skeptical of randomization.
-Numerical analysts questioned the lack of theoretical guarantees for the accuracy of outputs of randomized algorithms, as well as their reproducibility.
-The general consensus was that the increase in efficiency afforded by randomization meant a sacrifice in accuracy that made the algorithms inherently unreliable, particularly because Monte Carlo methods were often inaccurate as a byproduct of the central limit theorem, cf.
-(martinsson2020, Section 1.2).
-However, since the 1980s, that skepticism has been replaced by an abundance of work to develop randomized algorithms with provable accuracy and practical efficiency, a subfield of NLA that has grown into a research area in its own right, known as randomized numerical linear algebra (RNLA).
-More details on historical milestones in the RNLA methods of the 20th century can be found in halko2011; Kannan2017; Mahoney2011Randomized; martinsson2020; Tropp2015_MatrixConcentration; woodruff2015.
-RNLA in the 21st century   By the early 2000s, randomization was being utilized in ground-breaking applications, such as randomized sampling for web-searching AchlioptasFiatKarlinMcSherry2001; ChakrabartiDomGibsonKleinbergKumarRaghavanRajagopalanTomkins1999, including the notable Google PageRank algorithm BrinPage1998Anatomy; HenzingerHeydonMitzenmacherNajork1999; PageBrinMotwaniWinograd1998.
-Randomized sampling was also used to approximate matrix operations and compute low-rank matrix decompositions Achlioptas2001; Frieze2004; papadimitriou2000latent.
-By the mid-2000s, a unifying framework for RNLA had begun to emerge.
+Randomization has always played a role in NLA, for instance in sometimes providing a good starting point for a Krylov iteration.
+Another interesting early result was a proposal to randomly precondition linear systems to avoid the need for pivoting, which increases practical speed by reducing data movement parker1995random.
+However, randomized methods formed a small niche within the mainstream literature; one may speculate that concerns about reproducibility played a role, and perhaps also concerns that randomized methods could not be executed in a high accuracy mode, since that is a characteristic of Monte Carlo methods based on the central limit theorem.
+(For additional discussion of the role of randomization in NLA, we refer to Kannan2017; Mahoney2011Randomized; martinsson2020.) This survey is concerned with a set of techniques that we refer to as randomized numerical linear algebra (RNLA) that emerged in the early 2000s, and have in the years since attracted significant interest.
+RNLA in the 21st century: By the early 2000s, randomization had led to major developments in ground-breaking applications, such as the notable Google PageRank algorithm BrinPage1998Anatomy; HenzingerHeydonMitzenmacherNajork1999; PageBrinMotwaniWinograd1998.
+Randomized sampling was also being used to compute low-rank matrix decompositions Achlioptas2001; Frieze2004; papadimitriou2000latent.
 Over the past two decades, RNLA has developed into a firmly established research area, garnering widespread attention from both practitioners and theorists.
-In 2006, a three-part series laid the theoretical groundwork for randomized sampling in approximate matrix multiplication DrineasKannanMahoney2006, low-rank matrix decomposition (QR and SVD) DrineasKannanMahoney2006b, and CUR decomposition DrineasKannanMahoney2006c.
-The same year, the separate work of DeshpandeRademacherVempalaWang2006; Deshpande2006AdaptiveSA investigated adaptive sampling for low-rank approximation, and in sarlos2006improved, a relative error SVD algorithm was proposed that utilized randomized subspace embeddings.
+In 2006, a three-part series of papers laid the theoretical groundwork for randomized sampling in approximate matrix multiplication DrineasKannanMahoney2006, low-rank matrix decomposition (QR and SVD) DrineasKannanMahoney2006b, and CUR decomposition DrineasKannanMahoney2006c.
+Around the same time, DeshpandeRademacherVempalaWang2006; Deshpande2006AdaptiveSA investigated adaptive sampling for low-rank approximation; sarlos2006improved proposed a relative error SVD algorithm that utilized randomized subspace embeddings; and 2006_martinsson_random1_orig introduced what would be later referred to as the “randomized SVD (RSVD)”.
 Numerical analysts and theoretical computer scientists also provided compelling empirical evidence of the improved performance of RNLA algorithms for practical applications Deshpande2006AdaptiveSA; Friedland2005FastML; Frieze2004; Liberty2007; martinsson2006interpolation; RokhlinTygert2008.
 These results inspired subsequent investigations of randomized algorithms expressly for linear algebra tasks, such as preconditioning or rangefinding, paired with strong probabilistic guarantees Avron2010; HalkoMartinssonShkolniskyTygert2011; halko2011; MartinssonRokhlinTygert2011; Rokhlin2010.
-Twenty years later, the scope of research in RNLA now encompasses nearly every aspect of linear algebra, including eigenvalue problems bhattacharjee2023; BravermanKrishnanMusco2022; KressnerPlestenjak2024; LinSaadYang2016; Tropp2025_OSI; Tropp2022; YuanGuLi2018; general linear solvers AvronDruinskyGupta2015; Avron2010; BaboulinLiRouet2015; BalabanovGrigori2025; ChenEpperlyTroppWebber2025; NakatsukasaTropp2024; orthogonalization BalabanovGrigori2022; DuerschGu2017; FukayaKannanNakatsukasaYamamotoYanagisawa2020; GrigoriTimsit2024; JangGrigori2025; matrix function approximation AmselChenGreenbaumMuscoMusco2024; BurkeGuttel2024; ChenGreenbaumMuscoMusco2023; CortinovisKressnerNakatsukasa2024; GuttelSchweitzer2023; PerssonKressner2023; UbaruChenSaad2017; trace estimation AvronToledo2011; GrattonTitley2018; HallmanIpsenSaibaba2023; PerssonCortinovisKressner2022; RoostaKhorasaniAscher2015; UbaruSaad2018; and many, many more.
+Twenty years later, the scope of research in RNLA now encompasses nearly every aspect of linear algebra, including eigenvalue problems BravermanKrishnanMusco2022; KressnerPlestenjak2024; Tropp2022; general linear solvers BalabanovGrigori2025; ChenEpperlyTroppWebber2025; NakatsukasaTropp2024; orthogonalization BalabanovGrigori2022; FukayaKannanNakatsukasaYamamotoYanagisawa2020; JangGrigori2025; matrix function approximation AmselChenGreenbaumMuscoMusco2024; CortinovisKressnerNakatsukasa2024; PerssonKressner2023; trace estimation HallmanIpsenSaibaba2023; PerssonCortinovisKressner2022; UbaruSaad2018; and many, many more.
 Why the R in RNLA?
 The rise in popularity of randomized algorithms is due in large part to the commensurate progress made in computing technology.
-In June 2006, the performance of the world’s top supercomputer BlueGene/L was measured at 2.806 Peta-flop/s (floating point operations per second) on the High Performance Linpack (HPL) benchmark, a program that solves a large dense system of linear equations by LU with partial pivoting.
-As of November 2025, that rate has grown exponentially to 1.742 Exa-flop/s, achieved by today’s top supercomputer El Capitan top500list.
-Modern computing architectures have made it possible to solve increasingly large-scale problems; however, as evidenced by both theory and practice, the performance of classical algorithms is limited by asymptotic complexity and the “curse of dimensionality” bellman1961adaptive.
-As a point of reference, the amount of data collected by the Event Horizon Telescope Collaboration to generate the first image of a black hole in 2019 was over 5 Petabytes Whitwam2019, which is almost the full storage capacity of El Capitan.
-In other words, classical matrix algorithms with cubic cost in the input size, such as orthogonalization, may be untenable given the large-scale problems that are now considered in scientific investigations.
+In June 2006, the performance of the world’s top supercomputer BlueGene/L was measured at 2.806 Peta-flops/s (floating point operations per second) on the High Performance Linpack (HPL) benchmark, a program that solves a large dense system of linear equations by LU with partial pivoting.
+As of November 2025, that rate has grown to 1.742 Exa-flops/s, achieved by today’s top supercomputer El Capitan.
+Modern computing architectures have made it possible to solve increasingly large-scale problems; however, as evidenced by both theory and practice, the performance of classical algorithms is limited by asymptotic complexity and by the cost of data movement.
+As a point of reference, the amount of data collected by the Event Horizon Telescope Collaboration to generate the first image of a black hole in 2019 was over 5 Petabytes, which is almost the full storage capacity of El Capitan.
+In other words, classical matrix algorithms with cubic cost in the input size, such as orthogonalization, are untenable given the large-scale problems now considered in scientific investigations.
 For the best practical performance, it is necessary to develop algorithms and implementations that can fully leverage computing resources.
 Perhaps the greatest benefit of randomization is that it gives us the ability to restructure algorithms to attenuate computational demands on whichever computing resources most hinder performance, such as flops, matrix access, or data movement.
-For example, randomization admits algorithms that require just a single pass over the input data halko2011; LiYin2020SinglePassLU; Mahoney2011Randomized; Murray2023RandNLA; Tropp2017Practical; Tropp2018-MorePracticalSketching; TroppStreaming2019; Woolfe2008FastRandomized, which is especially advantageous for large-scale problems or streaming data.
-Randomization can also significantly reduce the number of arithmetic operations required by classical algorithms; e.g., for dense $m\times n$ overdetermined least squares problems with $n\ll m$, randomization yields solutions in roughly $O(mn+n^{3})$ operations vs.
-$O(mn^{2})$ with classical approaches Avron2010; Epperly2024; LacottePilanci2021; OzaslanPilanciArikan2019; PilanciWainwright2016; RokhlinTygert2008.
+For example, randomization admits algorithms that require just a single pass over the input data LiYin2020SinglePassLU; Tropp2018-MorePracticalSketching; TroppStreaming2019, which is especially advantageous for large-scale problems or streaming data.
+Randomization can also significantly reduce the number of arithmetic operations required by classical algorithms; e.g., for dense $m\times n$ overdetermined least squares problems with $m\gg n$, randomization yields solutions in roughly $O(mn+n^{3})$ operations vs.
+$O(mn^{2})$ with classical approaches Epperly2024; LacottePilanci2021; OzaslanPilanciArikan2019; RokhlinTygert2008.
 In short, randomized algorithms produce accurate solutions with high probability while reducing working storage and communication costs, offering improved asymptotic complexity, admitting straightforward parallelization and GPU acceleration, and enabling practical investigations that would be otherwise impossible.
-For an expanded discussion, see (martinsson2020, Sections 1.3-1.4).
-RNLA for tensors   The successful incorporation of randomization into classical matrix algorithms inspired similar work to exploit randomization for fast tensor computations.
+For an expanded discussion, see martinsson2020.
+RNLA for tensors: The successful incorporation of randomization into classical matrix algorithms inspired similar work to exploit randomization to accelerate tensor computations.
 Originally proposed in 1927 as a means of working with multi-way physical and chemical measurements Hitchcock1927TheEO; Hitchcock1927Multiple, tensors are now ubiquitous in scientific computing and data science.
-Applications that involve tensorial data include 3D image reconstruction Bendory2023Autocorrelation; Ghosh2022; Kam1980Reconstruction; Levin2018AbInitio; Zhang2024MomentMetrics; signal processing Chen2021_WirelessComm; CichockMandicLathauweri2015; Comon2001TensorDecompositions; Marot2008TensorDenoising; Miron2020; Sidiropoulos2017; quantum chemistry and physics Bischoff2012ManyBodyWaveFunctions; Khoromskaia2015TensorNM; Peng2020; Pierce2021; Szalay2015; machine learning Aidini2023FewShotTensorCompletion; Hao2013; Ji2019_ML; Panagakis2024; Tao2005; Zhou2016GroupComponentAnalysis; high-dimensional partial differential equations (PDEs) BachmayrSchneiderUschmajew2016; BoelensVenturiTartakovsky2018; DektorRodgersVenturi2021; Khoromskij2015TensorPDEs; LubichOseledetsVandereycken2015; among many others.
+Applications that involve tensorial data include 3D image reconstruction Bendory2023Autocorrelation; Ghosh2022; Zhang2024MomentMetrics; signal processing Chen2021_WirelessComm; Miron2020; Sidiropoulos2017; quantum chemistry and physics Bischoff2012ManyBodyWaveFunctions; Peng2020; Pierce2021; machine learning Aidini2023FewShotTensorCompletion; Ji2019_ML; Panagakis2024; high-dimensional partial differential equations (PDEs) BoelensVenturiTartakovsky2018; DektorRodgersVenturi2021; Khoromskij2015TensorPDEs; among many others.
 For detailed surveys of tensors and their applications, we refer the reader to Ballard_Kolda_2025; FriedlandTammali2015LowRankApprox; Grasedyck2013; KoldaBader2009; papalexakis2017.
-Unfortunately, tensor practitioners suffer harsher consequences of the curse of dimensionality.
-For a $d$-way array of uniform size $N$, the storage and computational costs scale as $N^{d}$, quickly becoming unmanageable.
-As such, randomization has been essential to the development of efficient algorithms for tensor computations, popularized by several foundational works in the mid-2000s CaiafaCichocki2009; delaVega2005; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR.
-Two decades later, tensor practitioners increasingly rely on randomization for efficient tensor computations that are backed by the strong accuracy guarantees of well-established randomized matrix algorithms.
-Randomized algorithms have been developed for nearly every type of low-rank tensor decomposition in the literature, including the CP format BattaglinoBallardKolda2018; Biagioni2015; Erichson_2020; MaSolomonik2021FastAccurateRandomizedTensor; Reynolds2016; VervlietDeLathauwer2016RandomizedBlockSampling; Wang2023, the Tucker format AhmadiAsl2021Randomized; CheWeiYan2025EfficientTucker; KressnerPerisa2017RecompressionHadamard; Minster2024; MinsterSaibabaKilmer2020; Tsourakakis2010; Zhou2014, the Tensor Train format CheWei2019; CheWeiYan2024; Fonat2019; HuberSchneiderWolf2017RandomizedTT_SVD; LiYuBatselier2022FasterTensorTrainSparse, as well as other specialized representations BallaniGrasedyckKluge2013BlackBoxHTucker; BatselierYuDanielWong2018TensorNetworkRandomizedSVD; CastilloHaddockHartsockEtAl2025RandomizedKaczmarzTPRODUCT; CheWeiYan2024SketchBasedMultiplicativeUpdating; Robeva2016; TarzanaghMichailidis2018; WangCuiLi2024SVD_TensorWheel; Yuan2019; ZhangSaibabaKilmerAeron2018RandomizedTensorSVD; Zhou2014NTD.
-Outline   In the following sections, we highlight recent computational advances in randomized algorithms for low-rank matrix and tensor decompositions.
-Sections 1-4 focus on the matrix setting.
-We begin with a review of requisite linear algebra material and classical matrix decomposition algorithms.
+Unfortunately, tensor practitioners suffer harsh consequences of the curse of dimensionality.
+For a $d$-way array of uniform size $n$, the storage costs scale as $O(n^{d})$, and most algorithms for analyzing the data scale significantly worse.
+The key contribution of randomization in this context has been to reveal new algorithms whose complexity scale linearly or close to linearly with the amount of input data, and that involve minimal data movement.
+This work was inspired by the foundational works CaiafaCichocki2009; delaVega2005; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR from the late 2000s.
+Two decades later, tensor practitioners increasingly rely on randomization for efficient tensor computations, backed by the strong accuracy guarantees of well-established randomized matrix algorithms.
+Randomized algorithms have been developed for nearly every type of low-rank tensor decomposition in the literature, including the CP format MaSolomonik2021FastAccurateRandomizedTensor; VervlietDeLathauwer2016RandomizedBlockSampling; Wang2023, the Tucker format AhmadiAsl2021Randomized; CheWeiYan2025EfficientTucker; KressnerPerisa2017RecompressionHadamard, the Tensor Train format CheWeiYan2024; HuberSchneiderWolf2017RandomizedTT_SVD; LiYuBatselier2022FasterTensorTrainSparse, as well as other specialized representations BallaniGrasedyckKluge2013BlackBoxHTucker; BatselierYuDanielWong2018TensorNetworkRandomizedSVD; CastilloHaddockHartsockEtAl2025RandomizedKaczmarzTPRODUCT; Robeva2016; TarzanaghMichailidis2018; WangCuiLi2024SVD_TensorWheel; Zhou2014NTD.
+Scope: Our intent with this survey is to bring attention to recent work extending the ideas of RNLA to the world of tensors.
+We have chosen to include a thorough review of the matrix case in order to make the survey self-contained, and to introduce notation and ideas that will be brought to bear in the tensor regime.
+(This also provided us the opportunity to review some recent advances on topics such as “fast” random embeddings.) To keep the survey focused, we had to omit several important and interesting topics.
+For matrices, we do not consider the vast body of work on randomized algorithms for eigenvalues and eigenvalue problems (see, e.g., HeKressnerPlestenjak2025; NakatsukasaTropp2024; Schneider2024_phd) or optimization (see, e.g., derezinski2020debiasing; na2023hessian).
+We also suppress the details of most theoretical results in favor of more background material; see, e.g., Camano2025_OSI; Kannan2017; martinsson2020.
+We refer the reader to Murray2023RandNLA for a software-focused survey of RNLA, and to DerezinskiMahoney2024 for a survey of RNLA from the perspective of machine learning.
+For tensors, we do not consider the tensor train format or other more specialized tensor decompositions and refer the reader to Ballard_Kolda_2025; Grasedyck2013.
+Outline: Sect.
+24 focus on the matrix setting, beginning with a review of requisite linear algebra material and classical matrix decomposition algorithms.
 We next survey existing methods of randomized dimension reduction in addition to promising new developments in randomized sketching and sampling.
-We end the first part of our manuscript by discussing recent work on randomized algorithms for low-rank matrix decompositions, which are the foundation of our randomized algorithms for tensors in part two.
-Sections 5-6 contain the necessary tensor prerequisites, as well as overviews of the two major tensor formats that we consider in our work: CP and Tucker.
-We then summarize key developments in randomized algorithms to compute low-rank tensor decompositions in these formats.
-We close with some final thoughts and an outlook on the future of RNLA.
-Scope   Our intent with this survey is to bring attention to important recent developments in the fast-growing body of work on RNLA.
-This rapid progress means that regular appraisal of the current research landscape is important for guiding new work, drawing connections to prior work, and keeping RNLA practitioners abreast of major achievements and open problems; however, it also means that important subject areas and contributions will necessarily be omitted.
-We quickly attempt to rectify this by pointing the interested reader to a few references on topics beyond the scope of our work.
-For matrices, we do not consider the vast body of work on randomized algorithms for eigenvalues and eigenvalue problems (see, e.g., HeKressnerPlestenjak2025; NakatsukasaTropp2024; Schneider2024_phd) or optimization (see, e.g., derezinski2020debiasing; na2023hessian).
-We also suppress the details of most theoretical results in favor of more background material to make our presentation self-contained; see, e.g., Camano2025_OSI; Kannan2017; martinsson2020.
-We refer the reader to Murray2023RandNLA for a survey of RNLA that centers on software implementations, and to DerezinskiMahoney2024 for a survey of RNLA from the perspective of machine learning.
-For tensors, we do not consider the tensor train format or other more specialized tensor decompositions and refer the reader to Ballard_Kolda_2025; Grasedyck2013.
+We end the first part of our manuscript by discussing recent work on randomized algorithms for low-rank matrix decompositions, which are the foundation of our randomized algorithms for tensors.
+Sect.
+5 contains the necessary tensor prerequisites, as well as overviews of the two major tensor formats that we consider in our work: CP and Tucker.
+We then summarize key developments in randomized algorithms to compute low-rank tensor decompositions in these formats in Sect.
+6.
+We close with a few final thoughts on RNLA in Sect.
+7.
 2 Matrix Preliminaries
-In this section, we summarize notation and key concepts in linear algebra that are frequently referenced in our work.
+We begin by summarizing notation and key concepts in linear algebra that are frequently referenced in our work.
 We introduce our notational conventions in Sect.
-2.1 and summarize several fundamental matrix decompositions and algorithms to compute them in Sect.
-2.2-2.4.
+2.1 and summarize several fundamental matrix decompositions, as well as algorithms to compute them, in Sect.
+2.2-2.5.
 2.1 Notation
@@
 We also adopt the notation of Golub and Van Loan golub2013 to reference submatrices.
-Namely, if $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, and $I=[i_{1},i_{2},\dots,i_{k}]\subset[m]$ and $J=[j_{1},j_{2},\dots,j_{l}]\subset[n]$ are (row and column, resp.) index sets, then $\bm{\mathsf{A}}(I,J)$ denotes the $k\times l$ matrix
-$\bm{\mathsf{A}}(I,J)=\begin{bmatrix}\bm{\mathsf{A}}(i_{1},j_{1})&\bm{\mathsf{A}}(i_{1},j_{2})&\dots&\bm{\mathsf{A}}(i_{1},j_{l})\\ \bm{\mathsf{A}}(i_{2},j_{1})&\bm{\mathsf{A}}(i_{2},j_{2})&\dots&\bm{\mathsf{A}}(i_{2},j_{l})\\ \vdots&\vdots&&\vdots\\ \bm{\mathsf{A}}(i_{k},j_{1})&\bm{\mathsf{A}}(i_{k},j_{2})&\dots&\bm{\mathsf{A}}(i_{k},j_{l})\\ \end{bmatrix}$.
+Namely, if $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, and $I=[i_{1},i_{2},\dots,i_{k}]\subseteq[m]$ and $J=[j_{1},j_{2},\dots,j_{l}]\subseteq[n]$ are (row and column, resp.) index sets, then $\bm{\mathsf{A}}(I,J)$ denotes
+$\bm{\mathsf{A}}(I,J)=\begin{bmatrix}\bm{\mathsf{A}}(i_{1},j_{1})&\bm{\mathsf{A}}(i_{1},j_{2})&\dots&\bm{\mathsf{A}}(i_{1},j_{l})\\ \vdots&\vdots&&\vdots\\ \bm{\mathsf{A}}(i_{k},j_{1})&\bm{\mathsf{A}}(i_{k},j_{2})&\dots&\bm{\mathsf{A}}(i_{k},j_{l})\\ \end{bmatrix}\in\mathbb{F}^{k\times\ell}.$
 The abbreviation $\bm{\mathsf{A}}(I,:)$ refers to the submatrix $\bm{\mathsf{A}}(I,[n])$, analogously for $\bm{\mathsf{A}}(:,J)$.
@@
 The diagonal elements $\{\sigma_{i}\}_{i=1}^{r}$ of $\mathbf{\Sigma}$ are the singular values of $\bm{\mathsf{A}}$, ordered so that $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$.
-The rank of $\bm{\mathsf{A}}$ is equal to the number of nonzero singular values.
-Suppose that the rank of $\bm{\mathsf{A}}$ is at least $k$.
+The rank of $\bm{\mathsf{A}}$ is the number of nonzero singular values.
 To obtain a rank-$k$ approximation of $\bm{\mathsf{A}}$, we can truncate its SVD after the first $k$ terms, defining $\bm{\mathsf{A}}_{k}=\sum_{i=1}^{k}\sigma_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{v}}_{i}^{*}$.
@@
 We adopt this notation when referring to the computation of the SVD, as well as $\texttt{svd}(\bm{\mathsf{A}},k)$ to represent the function outputting the truncation $\bm{\mathsf{A}}_{k}$.
-Golub-Reinsch and related algorithms rely on QR decompositions, covered next.
 2.3 The QR Decomposition
-Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits a QR decomposition given by
-$\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}=\underset{m\times r}{\bm{\mathsf{Q}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{R}}\hskip 5.69054pt},$ (3)
-where $r=\min(m,n)$, $\bm{\mathsf{Q}}$ is orthonormal and $\bm{\mathsf{R}}$ is upper triangular.
-Using a greedy algorithm that builds the QR factorization incrementally (e.g., with Householder reflections Householder1958; cf.
-Sect 2.4), we can halt the factorization after computation of the first $k$ terms to obtain a “partial” QR factorization:
-$\underset{m\times n}{\bm{\mathsf{A}}}\approx\underset{m\times k}{\bm{\mathsf{Q}}_{k}}\underset{k\times n}{\hskip 5.69054pt\bm{\mathsf{R}}_{k}\hskip 5.69054pt}.$ (4)
-We let
-$\begin{bmatrix}{\bm{\mathsf{Q}},\bm{\mathsf{R}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{Q}}_{k},\bm{\mathsf{R}}_{k}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}},k),$ (5)
-denote functions that yield either the full or partial QR factorization of $\bm{\mathsf{A}}$, respectively.
-Occasionally, if we need only the factor $\bm{\mathsf{Q}}$, whose columns form an orthonormal basis for the range of $\bm{\mathsf{A}}$, we will use the notation
+Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits a QR decomposition $\bm{\mathsf{A}}=\bm{\mathsf{Q}}\bm{\mathsf{R}}$ where $\bm{\mathsf{Q}}$ is an orthonormal matrix of size $m\times r$, and where $\bm{\mathsf{R}}$ is an $r\times n$ upper triangular matrix, where $r=\min(m,n)$.
+In what follows, we are particularly interested in column pivoted QR factorizations, which take the form
+$\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}\underset{n\times n}{\bm{\mathsf{P}}}=\underset{m\times r}{\bm{\mathsf{Q}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{R}}\hskip 5.69054pt},$ (3)
+for some permutation matrix $\bm{\mathsf{P}}$.
+The point here is that it is often possible to pick a permutation matrix $\bm{\mathsf{P}}$ in such a way that the truncation of (3) to the dominant terms forms a good low rank approximation to $\bm{\mathsf{A}}$.
+Common algorithms such as Golub-Businger with standard column pivoting perform well in practice, and more advanced pivoting techniques come with provable guarantees gu1996, cf.
+Sect 2.5.
+These methods are greedy, and can be halted after a fixed number $k$ steps.
+We use the notation
+$\begin{bmatrix}{\bm{\mathsf{Q}},\bm{\mathsf{R}},\bm{\mathsf{P}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{Q}}_{k},\bm{\mathsf{R}}_{k},\bm{\mathsf{P}}}\end{bmatrix}=\texttt{qr}(\bm{\mathsf{A}},k),$ (4)
+to refer to functions for computing either a full or a partial QR factorization.
+It is sometimes convenient to store the permutation using an index vector $I$ such that $\bm{\mathsf{P}}=\bm{\mathsf{I}}(\colon,I)$, where $\bm{\mathsf{I}}$ is the identity matrix.
+If we need only the factor $\bm{\mathsf{Q}}$, whose columns form an orthonormal basis for the range of $\bm{\mathsf{A}}$, we use the notation
 $\bm{\mathsf{Q}}=\texttt{col}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \bm{\mathsf{Q}}_{k}=\texttt{col}(\bm{\mathsf{A}},k).$ (5)
+2.4 The LU Decomposition
+Every matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ admits an LU decomposition given by
+$\underset{m\times m}{\bm{\mathsf{P}}}\underset{m\times n}{\hskip 5.69054pt\bm{\mathsf{A}}\hskip 5.69054pt}=\underset{m\times r}{\bm{\mathsf{L}}}\underset{r\times n}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt},$ (6)
+where $r=\min(m,n)$, $\bm{\mathsf{P}}$ is a permutation matrix, $\bm{\mathsf{L}}$ is lower triangular and $\bm{\mathsf{R}}$ is upper triangular.
+We can halt the factorization after computation of the first $k$ terms to obtain a “partial” LU factorization $\bm{\mathsf{A}}\approx\bm{\mathsf{L}}_{k}\bm{\mathsf{U}}_{k}$.
+$\begin{bmatrix}{\bm{\mathsf{L}},\bm{\mathsf{U}},P}\end{bmatrix}=\texttt{lu}(\bm{\mathsf{A}})\ \ \textup{and}\ \ \begin{bmatrix}{\bm{\mathsf{L}}_{k},\bm{\mathsf{U}}_{k},I}\end{bmatrix}=\texttt{lu}(\bm{\mathsf{A}},k),$ (7)
+denote functions that yield either the full or partial LU factorization, respectively, where $\bm{\mathsf{P}}$ is stored as a permutation vector $P$ and $I=P(1:k)$.
 2.5 The Interpolative and CUR Decompositions
@@
 $\begin{array}[]{ll}\bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}}\underset{k\times n}{\bm{\mathsf{R}}},&\hskip 14.22636pt\textit{Row ID}\\ \bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{C}}}\underset{k\times n}{\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}},&\hskip 14.22636pt\textit{Column ID}\\ \bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{C}}}\underset{k\times k}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt}\underset{k\times n}{\bm{\mathsf{R}}},&\hskip 14.22636pt\textit{CUR}\end{array}$ (11)
-where
-$\bm{\mathsf{R}}=\bm{\mathsf{A}}(I,:)$
-for $k$ linearly independent rows indexed by $I$;
-$\bm{\mathsf{C}}=\bm{\mathsf{A}}(:,J)$
-for $k$ linearly independent columns indexed by $J$; and $\bm{\mathsf{U}}=\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$, though in practice, $\bm{\mathsf{U}}$ is often approximated by $\bm{\mathsf{A}}(I,J)^{\dagger}$.
-The columns of $\bm{\mathsf{C}}$ or rows of $\bm{\mathsf{R}}$ are known as skeletons in an ID or CUR decomposition.
+where $\bm{\mathsf{R}}=\bm{\mathsf{A}}(I,:)$ for $k$ linearly independent rows indexed by $I$; $\bm{\mathsf{C}}=\bm{\mathsf{A}}(:,J)$ for $k$ linearly independent columns indexed by $J$; and $\bm{\mathsf{U}}=\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$, though in practice, $\bm{\mathsf{U}}$ is often approximated by $\bm{\mathsf{A}}(I,J)^{\dagger}$.
+The columns of $\bm{\mathsf{C}}$ or rows of $\bm{\mathsf{R}}$ are known as skeletons.
 Typically, constructing the CUR decomposition is more ill-conditioned than the IDs, but a stable construction can be attained through QR factorizations of $\bm{\mathsf{C}}$ and $\bm{\mathsf{R}}$.
@@
 where the coefficient $\sqrt{k+1}$ is optimal and cannot be improved; here, $\bm{\mathsf{A}}_{k}$ is the optimal rank-$k$ approximation given by the truncated SVD.
-In practice, greedy pivoting algorithms are frequently used to choose skeletons for ID and CUR decompositions.
-We discuss several popular choices in the following sections, including a recently proposed algorithm that achieves (12) with the best known complexity.
-QR with Column Pivoting (CPQR)
-Column-pivoted QR (CPQR) is the algorithm behind the qr function from Sect.
-2.3.
-In particular, column pivoting may be required for numerical stability, and the outputs of qr($\bm{\mathsf{A}}$) = $\bm{\mathsf{Q}},\bm{\mathsf{R}},P]$ represent the factorization $\bm{\mathsf{A}}\bm{\mathsf{P}}=\bm{\mathsf{Q}}\bm{\mathsf{R}}$ for some permutation matrix $\bm{\mathsf{P}}\in^{n\times n}$ such that $\bm{\mathsf{P}}=\bm{\mathsf{I}}(:,P)$.
-This factorization can be achieved through a sequence of permutations and Householder reflections that rank-one update the “active submatrix” of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ by greedy pivoting, based on squared column norms.
+In practice, greedy pivoting algorithms, based on QR, LU, or the SVD, are frequently used to choose skeletons for ID and CUR decompositions.
+We discuss several in the following sections.
+ID/CUR from QR with Column Pivoting (CPQR)
 Assuming $m\geq n$, let $\bm{\mathsf{A}}^{(t)}$ represent the resulting matrix after $t$ steps of pivoting and updating for $t=0,\ldots,n-2$ (with $\bm{\mathsf{A}}^{(0)}=\bm{\mathsf{A}}$).
-At the $(t+1)$-th step, CPQR determines the column pivot $j_{t+1}$ by finding the column of maximal Euclidean norm in the active submatrix,
+At the $(t+1)$-th step, CPQR chooses the column pivot $j_{t+1}$ as the column of maximal Euclidean norm in the active submatrix,
 $j_{t+1}=\arg\max_{t+1\leq j\leq n}\|\bm{\mathsf{A}}^{(t)}([t+1:m],j)\|_{2}.$
@@
 In practice, CPQR chooses pivots in an order that reveals the rank of $\bm{\mathsf{A}}$, so that the chosen skeletons form a good basis for its row or column space.
-While there are adversarial cases for which CPQR can fail to be rank-revealing (e.g., Kahan-type matrices kahan1966; peters1975), these are rarely encountered, cf.
+While there are adversarial cases for which CPQR can fail to be rank-revealing (e.g., Kahan-type matrices kahan1966), these are rarely encountered; cf.
 trefethen1990.
 There are variations of CPQR with better theoretical guarantees, such as rank-revealing QR Chan1987; Ilse1994 or strong rank-revealing QR GuEis1996, but they are more costly to implement.
-LU with Partial Pivoting
-The LU algorithm with row-wise partial pivoting (LUPP) yields a factorization
-$\underset{m\times m}{\hskip 11.38109pt\bm{\mathsf{P}}\hskip 11.38109pt}\underset{m\times n}{\bm{\mathsf{A}}}=\underset{m\times n}{\bm{\mathsf{L}}}\underset{n\times n}{\hskip 5.69054pt\bm{\mathsf{U}}\hskip 5.69054pt},$ (13)
-where $\bm{\mathsf{L}}$ is lower triangular with main diagonal entries equal to 1 and off-diagonal entries bounded in magnitude by 1, and $\bm{\mathsf{U}}$ is upper triangular.
-The function $\bm{\mathsf{L}},\bm{\mathsf{U}},\bm{\mathsf{P}}]=\texttt{lu}(\bm{\mathsf{A}})$ represents the factorization (13) obtained from row-wise partial pivoting.
-More specifically, to compute an LU factorization, row pivots are selected from the active submatrix of $\bm{\mathsf{A}}$ that is updated via Schur complements (e.g., (golub2013, Algorithm 3.2.1) (trefethen1997, Algorithm 21.1)).
+ID/CUR from LU with Partial Pivoting (LUPP)
+For an ID or CUR decomposition using LUPP, row pivots are selected from the active submatrix of $\bm{\mathsf{A}}$, which is updated after each pivot selection via Schur complements (e.g., (golub2013, Algorithm 3.2.1)).
 Let $\bm{\mathsf{A}}^{(t)}$ be the resulting matrix after $t$ steps of pivoting and updating for $t=0,1,\ldots,n-2$.
@@
 where $\bm{\mathsf{A}}^{(t)}(t+1,t+1)$ has been updated with $\bm{\mathsf{A}}^{(t)}(i_{t+1},t+1)$ following row permutation.
-This LUPP algorithm leads to an exponential upper bound on the entries of $\bm{\mathsf{U}}$, which is tight for certain matrices (e.g., Kahan-type matrices kahan1966; peters1975; sorensen2014).
-Thus, LUPP can fail to be rank-revealing, but such adversarial inputs are rarely encountered in practice trefethen1990.
-There exist more sophisticated pivoting strategies to lend rank-revealing properties to the LU factorization Chan84; Pan2000; Zhao2005 with better error bounds, but at significantly greater computational cost.
-Fortunately, we will see in Sect.
-4.3 that randomization can equip LUPP with rank-revealing properties comparable to CPQR even in adversarial cases.
-Truncated SVD for Near-Optimal Skeletons
+If expressed as the output of the function $\begin{bmatrix}\bm{\mathsf{L}},\bm{\mathsf{U}},P\end{bmatrix}=\texttt{lu}(\bm{\mathsf{A}})$, the row skeletons at the $(t+1)$-th step are given by $I=P(1:t+1)$.
+This algorithm leads to an exponential upper bound on the entries of $\bm{\mathsf{U}}$, which is tight for certain matrices (e.g., Kahan-type matrices kahan1966).
+Thus, LUPP can fail to be rank-revealing, but again this is rare in practice trefethen1990.
+There are more sophisticated pivoting strategies to lend rank-revealing properties to the LU factorization Chan84; Pan2000, but at significantly greater computational cost.
+We will see that randomization can equip LUPP with rank-revealing properties comparable to CPQR even in adversarial cases.
+ID/CUR from a Single Truncated SVD
 Previously, the tight error bound (13) had been achieved by expensive random volume sampling DeshpandeRademacher2010, or computing at least $k$ SVDs to choose each pivot from an updated residual subspace.
@@
 $\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,j}\|_{2}/\|\bar{\bm{\mathsf{V}}}^{*}_{1:(k-i+1),j}\|_{2}.$ (17)
-By the upper triangularity of $\widehat{\bm{\mathsf{V}}}^{*}$ and the relation (16),
+By (16) and the upper triangularity of $\widehat{\bm{\mathsf{V}}}^{*}$,
 $\sum_{\ell=1}^{n}\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,\ell}\|^{2}_{2}$ $=\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}\|^{2}_{\text{F}},$ (18)
-and
 $\sum_{\ell=1}^{n}\|\bar{\bm{\mathsf{V}}}^{*}_{1:(k-i+1),\ell}\|^{2}_{2}$ $=\|\bar{\bm{\mathsf{V}}}^{*}\|^{2}_{\text{F}}=k-i+1.$ (19)
@@
 $\|\widetilde{\bm{\mathsf{A}}}^{(i)}\|^{2}_{\text{F}}=\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}-\widetilde{\bm{\mathsf{A}}}^{(i-1)}_{1:m,i}\bar{\bm{\mathsf{V}}}^{*}_{1:(k-i+1),i}\bar{\bm{\mathsf{V}}}^{*}\|_{\text{F}}^{2}\leq\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}\|^{2}_{\text{F}}+\frac{1}{k-i+1}\|\widetilde{\bm{\mathsf{A}}}^{(i-1)}\|^{2}_{\text{F}}.$
-If we add this up over all previous steps, we have
+Adding up over all previous steps, we have
 $\|\widetilde{\bm{\mathsf{A}}}^{(i)}\|^{2}_{\text{F}}\leq\|\widetilde{\bm{\mathsf{A}}}\|^{2}_{\text{F}}+\frac{i}{k-i+1}\|\widetilde{\bm{\mathsf{A}}}\|^{2}_{\text{F}}.$ (20)
-Letting $i=k$ in (20), we achieve the desired bound in (12) in $O(mnk)$ operations. Full details can be found in Osinsky2025.
+Letting $i=k$ in (20), we achieve the desired bound in (13) in $O(mnk)$ operations.
 3 Randomized Dimension Reduction: Sketching and Sampling
@@
 If the linear map is drawn from a random matrix distribution that achieves geometry-preserving dimension reduction with high probability, it is known as a randomized DRM.
-Randomized dimension reduction is frequently associated with randomized embeddings, linear maps drawn from random matrix distributions that satisfy the following condition with high probability.
+Randomized DRMs are often associated with randomized embeddings, linear maps drawn from random distributions that satisfy the following condition with high probability.
 Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be a matrix of rank $k\leq\min(m,n)$ and $\varepsilon\in(0,1)$ a distortion parameter.
@@
 $(1-\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|\leq\|\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}\bm{\mathsf{x}}\|\leq(1+\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|\ \ \forall\ \bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn}.$ (21)
-Usually, $k\leq d\ll m,$ so that the application of $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ results in significant dimension reduction.
+Usually, $k\leq d\ll m$.
 The minimal value of $d$ required for $\bm{\mathsf{\Gamma}}$ to satisfy (21) depends on the type of embedding $\bm{\mathsf{\Gamma}}$, the rank $k$ of $\bm{\mathsf{A}}$, and the distortion parameter $\epsilon$.
-As a classical example, the Johnson-Lindenstrauss lemma jlind84 establishes that $d\geq 8\log k/\varepsilon^{2}$ is sufficient for a random matrix $\bm{\mathsf{\Gamma}}$ with standard Gaussian entries to embed $k$ points in m so that (21) is satisfied.
+As a classical example, the Johnson-Lindenstrauss lemma jlind84 establishes that $d\geq 8\log k/\varepsilon^{2}$ is sufficient for $\bm{\mathsf{\Gamma}}$ with standard Gaussian entries to embed $k$ points in m to satisfy (21).
 More generally, if $\mathcal{S}$ is any probability distribution over linear maps $\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$, then $\bm{\mathsf{\Gamma}}\sim\mathcal{S}$ is a randomized embedding if (21) holds over $\mathcal{S}$ for all $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ with at least constant probability.
 If a randomized embedding can be constructed to map an unknown subspace into a lower-dimensional space (i.e.
-without knowing anything about the geometry of the input space except its dimension), then it is said to be an oblivious subspace embedding sarlos2006improved.
-It is often the case in practice that randomized DRMs are assumed to be randomized embeddings, but we will see in Sect.
-3.1 that satisfying (21) is sufficient but not necessary to ensure the smaller-dimensional space is a good proxy for the input space.
+without knowing anything about the input space except its dimension), then it is said to be an oblivious subspace embedding (OSE) sarlos2006improved.
+Many times in practice, randomized DRMs are treated as embeddings, but we will see in Sect.
+3.1 that satisfying (21) is sufficient but not necessary for the smaller-dimensional space to be a good proxy for the input.
 In Sect.
-3.1, we review randomized sketching, a procedure that produces random linear combinations of input matrix coordinates, which we say is “coordinate-mixing.” We then focus on dimension reduction by randomized sampling of matrix coordinates in Sect.
+3.1, we review randomized sketching, a procedure that produces random linear combinations of input matrix coordinates, which we say is “coordinate-mixing.” We then focus on randomized sampling of matrix coordinates in Sect.
 3.2.
-We end this section with an important application of randomized dimension reduction that will be relevant for our discussion of tensor decompositions: solving overdetermined least squares problems.
-For detailed treatments of randomized dimension reduction, the reader is referred to Camano2025_OSI; halko2011; martinsson2019randomized; martinsson2020; woodruff2015.
+We end with an important application of randomized dimension reduction that will be relevant for our discussion of tensor decompositions: solving overdetermined least squares problems.
+For detailed treatments, the reader is referred to Camano2025_OSI; halko2011; martinsson2019randomized; martinsson2020; woodruff2015.
 3.1 Randomized Matrix Sketching
-The application of a randomized dimension reduction map $\bm{\mathsf{\Gamma}}$ to $\bm{\mathsf{A}}$ is known as randomized sketching.
-We reserve the term randomized sketching for random linear maps that are coordinate-mixing, vs.
-coordinate-sampling.
-The matrix product $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ is said to be a row sketch of $\bm{\mathsf{A}}$, and a column sketch $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is defined analogously.
-Ideally, the random matrix used for sketching should roughly preserve the geometry of the row or column space of $\bm{\mathsf{A}}$.
-Randomized dimension reduction maps $\bm{\mathsf{\Gamma}}$ that are frequently utilized for sketching include:
-  • Gaussian matrices ((halko2011, Section 4.1), (martinsson2020, Section 8.3), (woodruff2015, Theorem 2.3)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ with entries
-    $\bm{\mathsf{\Gamma}}(i,j)\sim\mathcal{N}(0,1/d),$
-    i.i.d. Gaussian entries drawn from a normal distribution with mean 0 and standard deviation $1/d$.
+The application of a randomized DRM $\bm{\mathsf{\Gamma}}$ to $\bm{\mathsf{A}}$ is known as randomized sketching.
+We refer to $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ as a row sketch of $\bm{\mathsf{A}}$, with a column sketch $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ defined analogously.
+We generally reserve the term “sketching” for random linear maps that are coordinate-mixing, as opposed to methods that simply sub-sample columns.
+Ideally, the random matrix used for sketching should roughly preserve the geometry of the row or column space of $\bm{\mathsf{A}}$. Commonly used randomized DRMs $\bm{\mathsf{\Gamma}}$ include:
+  • Gaussian matrices ((halko2011, Section 4.1), (martinsson2020, Section 8.3), (woodruff2015, Theorem 2.3)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ with entries $\bm{\mathsf{\Gamma}}(i,j)\sim\mathcal{N}(0,1/d).$
   • Subsampled randomized trigonometric transforms (SRTT) ((halko2011, Section 4.6), (martinsson2020, Section 9.3), tropp2011, (woodruff2015, Theorem 2.4)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
-    $\bm{\mathsf{\Gamma}}=\sqrt{\frac{m}{d}}\bm{\mathsf{\Pi}}_{m}\bm{\mathsf{F}}\bm{\mathsf{\Phi}}\bm{\mathsf{\Pi}}_{m\rightarrow d}.$
+    $\bm{\mathsf{\Gamma}}=\sqrt{m/d}\ \bm{\mathsf{\Pi}}_{m}\bm{\mathsf{F}}\bm{\mathsf{\Phi}}\bm{\mathsf{\Pi}}_{m\rightarrow d}.$
     Here, $\bm{\mathsf{\Pi}}_{m\rightarrow d}\in\mathbb{F}^{\hskip 0.56905ptm\times d}$ is a uniformly random selection of $d$ out of $m$ rows, whereas $\bm{\mathsf{\Pi}}_{m}\in\mathbb{F}^{\hskip 0.56905ptn\times n}$ is a random permutation of the $m$ rows.
-The matrix $\bm{\mathsf{F}}\in\mathbb{F}^{\hskip 0.56905ptm\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{\Phi}}:=\textup{diag}(\phi_{1},\ldots,\phi_{n})$ has i.i.d.
-Rademacher random variable entries, i.e.
-when $\mathbb{F}=$, entries of $\pm 1$ chosen independently with equal probability.
-  • Sparse sign matrices ((martinsson2020, Section 9.2), nelson2013, (woodruff2015, Theorem 2.5)): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
-    $\bm{\mathsf{\Gamma}}=\sqrt{\frac{m}{\zeta}}\begin{bmatrix}{\bm{\mathsf{s}}_{1},\ldots,\bm{\mathsf{s}}_{m}}\end{bmatrix}$
-    for sparsity parameter $2\leq\zeta\leq d$, where each column $\bm{\mathsf{s}}_{j}\in\mathbb{F}^{\hskip 0.56905ptd}$ for $j=1,\ldots,m$ is filled with $\zeta$ independent Rademacher random variables at uniformly random coordinates.
-In other words, each column is an independent random vector having $\zeta$ nonzero entries of $\pm 1$ (when $\mathbb{F}=$) in uniformly random positions.
-  • CountSketch matrices (Charikar2004; meng2013low; nelson2013): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as a sparse sign matrix with sparsity parameter $\zeta=1$.
-We caution that CountSketch matrices on their own can perform poorly in practice among randomized DRMs (see the required sketching dimension $d$ in Table 1).
+$\bm{\mathsf{F}}\in\mathbb{F}^{\hskip 0.56905ptm\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{\Phi}}:=\textup{diag}(\phi_{1},\ldots,\phi_{n})$ has i.i.d.
+Rademacher entries.
+  • Sparse sign matrices ((martinsson2020, Section 9.2), nelson2013, woodruff2015): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
+    $\bm{\mathsf{\Gamma}}=\sqrt{m/\zeta}\begin{bmatrix}{\bm{\mathsf{s}}_{1},\ldots,\bm{\mathsf{s}}_{m}}\end{bmatrix}$
+    for sparsity parameter $2\leq\zeta\leq d$, where each column $\bm{\mathsf{s}}_{j}\in\mathbb{F}^{\hskip 0.56905ptd}$, $j=1,\ldots,m$ contains $\zeta$ i.i.d.
+Rademacher random variables at uniformly random coordinates.
+  • CountSketch matrices (Charikar2004; meng2013low; nelson2013): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as a sparse sign matrix with sparsity $\zeta=1$.
+We advise against using CountSketch matrices on their own.
   • SparseStack matrices (Camano2025_OSI; Cohen2016ObliviousSubspace; KaneNelson2014): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ defined as
@@
   • SparseCol matrices (Camano2025_OSI): For consistency with the previous definitions of row sketching matrices as subspace embeddings, we give the transpose of the SparseCol matrix in (Camano2025_OSI, Definition 4.1).
-Namely, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ is defined by its $d$ rows of sparse random vectors
+Namely, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ is defined by its $d$ rows given by
     $\bm{\mathsf{\omega}}=\sqrt{\frac{m}{\zeta}}\sum_{i=1}^{\zeta}\rho_{i}\bm{\mathsf{e}}_{s_{i}}\in\mathbb{F}^{m},$ (22)
@@
 Rademacher random variables, the indices $s_{1},\ldots,s_{\zeta}$ are sampled uniformly from $[m]$ without replacement, and $\bm{\mathsf{e}}_{s_{i}}$ is the $s_{i}$-th standard basis vector.
-In other words, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ is the transpose of the SparseCol matrix $\bm{\mathsf{\Gamma}}^{*}\in\mathbb{F}^{m\times d}$ given by
-    $\bm{\mathsf{\Gamma}}^{*}:=\frac{1}{\sqrt{k}}\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{d}\end{bmatrix}$, (23)
-    where $\bm{\mathsf{\omega}}_{j}\sim\bm{\mathsf{\omega}}$ are i.i.d.
+In other words, $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ is the transpose of the SparseCol matrix $\bm{\mathsf{\Gamma}}^{*}\in\mathbb{F}^{m\times d}$ given by $\bm{\mathsf{\Gamma}}^{*}:=\frac{1}{\sqrt{k}}\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{d}\end{bmatrix}$, $\bm{\mathsf{\omega}}_{j}\sim\bm{\mathsf{\omega}}$ i.i.d.
   • SparseRTT matrices ((Camano2025_OSI, Section 4.1), tropp2011): $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ defined as
@@
     where $\bm{\mathsf{D}}\in\mathbb{F}^{m\times m}$ is a random diagonal matrix populated with i.i.d.
-Rademacher random variables, $\bm{\mathsf{F}}\in\mathbb{F}^{m\times m}$ is a discrete Fourier or Walsh-Hadamard transform, and $\bm{\mathsf{S}}\in\mathbb{F}^{d\times m}$ is the transpose of a SparseCol matrix with exactly $\zeta$ nonzero entries per column as in the SparseCol matrix above.
+Rademacher random variables, $\bm{\mathsf{F}}\in\mathbb{F}^{m\times m}$ is a discrete Fourier transform, and $\bm{\mathsf{S}}\in\mathbb{F}^{d\times m}$ is the transpose of a SparseCol matrix with exactly $\zeta$ nonzero entries per column.
 These distributions yield randomized DRMs that work well in practice, with theoretical guarantees summarized in Table 1.
-Gaussian test matrices are oblivious subspace embeddings that enjoy a superior lower bound on the requisite embedding dimension $d$ to guarantee (21), but the costs of storage and matrix-vector products for Gaussians are much more expensive than for “structured” embeddings like SRTT or sparse sign matrices.
-Put simply, Gaussian sketching makes for straightforward analysis, but randomized sketching with structured random matrices makes for better performance.
-The reader is referred to (martinsson2020, Sections 8,9) for detailed discussion.
+Gaussian test matrices are OSEs that enjoy a superior lower bound on the requisite embedding dimension $d$ for (21), but the costs of storage and matrix-vector products for Gaussians are much more expensive than for “structured” embeddings like SRTTs or sparse sign matrices.
+Put simply, Gaussian sketching makes for straightforward analysis, but structured random matrix sketching makes for better performance.
+The reader is referred to (martinsson2020, Sections 8,9) for details.
 Randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ Dimension $d$ for (21) Storage cost Mat-vec cost
@@
 SparseStack $d=\Omega\left(\nicefrac{{\mbox{\small$k\log k$}}}{{\mbox{\small$\varepsilon^{2}$}}}\right),\ \zeta=\Omega\left(\nicefrac{{\mbox{\small$\log k$}}}{{\mbox{\small$\varepsilon$}}}\right)$ $O(d\zeta)$ $O(\zeta m)$
-Table 1: Summary of coordinate-mixing randomized DRMs commonly used for (row-)sketching rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, which are also oblivious subspace embeddings (OSE).
-Theoretical lower bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy (21) are provided, as well as asymptotic complexities to store $\bm{\mathsf{\Gamma}}$ and to apply $\bm{\mathsf{\Gamma}}$ to a vector $\bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptm}$, cf.
-Camano2025_OSI; martinsson2020; woodruff2015.
-While the embedding dimensions in Table 1 guarantee the oblivious subspace embedding (OSE) property (21), it has been observed that smaller values of $d$ are often sufficient for many linear algebra tasks in practice when utilizing any of the randomized DRMs of Table 1 (with the exception of CountSketch), e.g., $d=k+10$, cf.
+Table 1: Summary of coordinate-mixing randomized DRMs commonly used for (row-)sketching rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$.
+Theoretical lower bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy the OSE property are provided, as well as the cost to store $\bm{\mathsf{\Gamma}}$ and to apply $\bm{\mathsf{\Gamma}}$ to a vector.
+While the embedding dimensions in Table 1 guarantee the OSE property (21), it has been observed that smaller values of $d$ are often sufficient for many linear algebra tasks in practice when utilizing any of the randomized DRMs of Table 1 (with the exception of CountSketch), e.g., $d=k+10$, cf.
 Sect.
 4.1.
-Recently, this observation was thoroughly investigated from the perspective of oblivious subspace injections, versus embeddings, which relaxes the OSE property to the following weaker one.
-A random matrix $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an oblivious subspace injection (OSI) of $\bm{\mathsf{A}}$ with injectivity $\alpha$ if it satisfies the injectivity property
+Recently, this observation was thoroughly investigated from the perspective of oblivious subspace injections, versus embeddings, which relaxes the OSE property as follows.
+Random matrix $\bm{\mathsf{\Gamma}}:\mathbb{F}^{\hskip 0.56905ptm}\rightarrow\mathbb{F}^{\hskip 0.56905ptd}$ is an oblivious subspace injection (OSI) of $\bm{\mathsf{A}}$ with injectivity $\alpha$ if it satisfies the injectivity property
 $\alpha\|\bm{\mathsf{A}}\bm{\mathsf{x}}\|\leq\|\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}\bm{\mathsf{x}}\|\ \ \forall\ \bm{\mathsf{x}}\in\mathbb{F}^{\hskip 0.56905ptn},$ (24)
@@
 With the OSI property guiding our choice of random sketching matrix, we enjoy more favorable sketching dimensions for the random matrices considered above.
-In the OSI framework, Gaussian random matrices are still the gold standard for approximation accuracy, but suffer the same performance drawbacks in terms of storage and matrix-vector product, or mat-vec, costs.
-Structured random matrices offer significantly better performance, and specifically, sparse random matrices exhibit the best empirical performance, as demonstrated by many practical investigations Chen2025_sparseGPU; ChenakkodDereZinskiDongRudelson2024_OptimalEmbeddingSparse; EpperlyMeierNakatsukasa2024_FastRandomizedLeastSquares; halko2011; KaneNelson2014; martinsson2020; nelson2013; Tropp2015_MatrixConcentration; woodruff2015, and now also supported by the theoretical guarantees of Camano2025_OSI; Tropp2025_OSI summarized in Table 2.
+In the OSI framework, Gaussian random matrices are still the gold standard for accuracy, but suffer the same performance drawbacks in terms of storage and matrix-vector product, or mat-vec, costs.
+Structured random matrices offer significantly better performance, and specifically, sparse random matrices exhibit the best empirical performance, as demonstrated by many practical investigations Chen2025_sparseGPU; ChenakkodDereZinskiDongRudelson2024_OptimalEmbeddingSparse; EpperlyMeierNakatsukasa2024_FastRandomizedLeastSquares; halko2011; KaneNelson2014; martinsson2020; nelson2013; Tropp2015_MatrixConcentration; woodruff2015, and now also supported by the theoretical guarantees of Camano2025_OSI; Tropp2025_OSI in Table 2.
 In general, sparse random matrices are recommended for sketching a dense unstructured input matrix, due to the reduced storage costs and significant acceleration afforded by high-performance sparse arithmetic libraries.
-To that end, sketching with SparseRTT matrices, or sparsified SRTTs, is more computationally efficient than with standard SRTTs or Gaussians.
 Compared to the SRTT, the SparseRTT matrix can reduce data movement and better facilitate parallelism at the hardware level.
 However, the SparseStack matrix is typically faster and thus recommended over SparseRTTs whenever sparse linear algebra packages are available.
-As an additional boon, the OSI values of $d$ that are reported in Table 2 are even still too pessimistic for practical implementation purposes.
-In particular, the results of Camano2025_OSI provide compelling evidence that SparseStack matrices are the superior choice of DRM in practice; it is conjectured that SparseStack matrices can achieve the $\frac{1}{2}$-OSI property with $d=O(k)$ and constant sparsity $\zeta=O(1)$.
+As an additional boon, the OSI values of $d$ that are reported in Table 2 are even still too pessimistic in practice.
+The results of Camano2025_OSI provide compelling evidence that SparseStack matrices are the superior choice of DRM; it is conjectured that SparseStack matrices can achieve the $\frac{1}{2}$-OSI property with $d=O(k)$ and constant sparsity $\zeta=O(1)$.
 This conjecture is supported numerically, with $d=2k$ and $\zeta=4$ used in the experiments of Camano2025_OSI to demonstrate that the low-rank approximations obtained from the SparseStack sketches are of comparable accuracy to those obtained from Gaussian sketches.
@@
 Sect.
-5.3), we can construct Khatri-Rao sketching matrices that exploit underlying Kronecker product structure Camano2025_OSI.
-As we show in Sect.
-6.1, these sketching matrices are formed as Kronecker products of isotropic random vectors.
-Often, these random vectors are drawn from (real) Gaussian and Rademacher distributions, but it is advised in Camano2025_OSI that the complex spherical distribution be used to avoid near-zero injectivity $\alpha$, which can occur in worst-case scenarios.
-For our sketching matrix constructions in Sect.
-6, we do not assume any specific isotropic distribution unless otherwise specified.
+5.3), we can construct Khatri-Rao sketching matrices that exploit underlying structure Camano2025_OSI, formed as Kronecker products of isotropic random vectors.
+Often, these random vectors are drawn from (real) Gaussian or Rademacher distributions, but it is advised in Camano2025_OSI that the complex spherical distribution be used to avoid near-zero injectivity in the worst-case.
+We do not assume any specific isotropic distribution unless otherwise specified.
 DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$ Dimension $d$ for $\varepsilon$-OSE (21) Dimension $d$ for $1/2$-OSI (24)
@@
 SparseRTT $d=O(k),\zeta=O(\log^{3}k)^{\dagger}$ $d=O(k),\zeta=O(\log k)$
-Table 2: Summary of coordinate-mixing randomized DRMs commonly used for row-sketching rank-$k$ $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, which are also oblivious subspace injections (OSI).
+Table 2: Summary of randomized DRMs for row-sketching rank-$k$ $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$.
 Theoretical bounds on the dimension $d$ for $\bm{\mathsf{\Gamma}}$ to satisfy the OSI property (24) vs.
@@
 3.2 Randomized Matrix Sampling
-In the previous section, we observed that a random sketch $\bm{\mathsf{\Gamma}}\bm{\mathsf{A}}$ (or $\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$) is a new matrix whose rows (or columns) are random linear combinations of the rows (or columns) of $\bm{\mathsf{A}}$.
+In the previous section, we observed that a random sketch has rows (or columns) that are random linear combinations of the rows (or columns) of $\bm{\mathsf{A}}$.
 However, we can also construct randomized DRMs that sample coordinates of $\bm{\mathsf{A}}$ according to a probability distribution.
-We use the separate term randomized sampling to refer to these types of maps.
+We use the separate term randomized sampling for these maps.
 To illustrate, suppose that we want to draw $d$ random i.i.d.
@@
 Common sampling methods include:
-  • Uniform ((martinsson2020, Section 9.6.2), (tropp2011b, Lemma 3.4), Kannan2017):
-    $p_{i}=1/m.$
-  • Squared (row) norms (Frieze2004):
-    $p_{i}=\|\bm{\mathsf{a}}_{i}\|_{2}^{2}/\|\bm{\mathsf{A}}\|^{2}_{\text{F}},$
-    where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$,
+  • Uniform ((martinsson2020, Section 9.6.2), (tropp2011b, Lemma 3.4), Kannan2017): $p_{i}=1/m.$
+  • Squared (row) norms (Frieze2004): $p_{i}=\|\bm{\mathsf{a}}_{i}\|_{2}^{2}/\|\bm{\mathsf{A}}\|^{2}_{\text{F}},$ where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$.
   • Row leverage scores (DrineasKannanMahoney2006; DrineasMahoneyMuthukrishnan2006, (martinsson2020, Section 9.6.3)):
@@
     where $\bm{\mathsf{a}}_{i}^{*}$ is the $i$th row of $\bm{\mathsf{A}}$ and the norm is defined as $\|\bm{\mathsf{v}}\|_{\bm{\mathsf{B}}}=\sqrt{\bm{\mathsf{v}}^{*}\bm{\mathsf{B}}\bm{\mathsf{v}}}$.
-  • Determinantal point processes (DPP) ((DerezinskiMahoney2021, Section 5.2), kulesza2012determinantal): If $S\subseteq[m]$, $|S|=d$,
+  • Determinantal point processes (DPP) ((DerezinskiMahoney2021, Section 5.2), kulesza2012determinantal): If $S\subseteq[m]$,
     $p_{S}=\frac{\det\left((\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}\right)}{\det\left(\bm{\mathsf{I}}+\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\right)},$ (26)
-    where $(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}$ is the sub-matrix of $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\in\mathbb{F}^{m\times m}$ indexed by $S$ and all indices $S$ are selected at once.
-However, in practice, vectors (rows) corresponding to indices $S$ are selected iteratively, as in the algorithm presented in Hough_DPP_2005.
+    where $|S|=d$, $(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})_{S,S}$ is the sub-matrix of $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}\in\mathbb{F}^{m\times m}$ indexed by $S$, and $S$ is selected all at once.
+However, in practice, rows corresponding to $S$ are selected iteratively, as in Hough_DPP_2005.
 Let $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}=\sum_{i}\lambda_{i}\bm{\mathsf{u}}_{i}\bm{\mathsf{u}}_{i}^{*}$ denote the eigendecomposition with $\lambda_{i}\in[0,1]$ for all $i$.
@@
       Update the probabilities and repeat d1d-1 times.
-  • Nuclear norm maximization (Altschuler2016GreedyCSS; Fornace2024; Ordozgoiti2018IterativeCSS; Ordozgoiti2019RegularizedCSS): While neither a traditional sketching or sampling method, we include nuclear norm maximization in this list because of its comparable performance to DPP sampling.
-Let $\bm{\mathsf{K}}=\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$, $S\subseteq[m]$, $|S|=d$, and define
-    $\mathcal{L}_{\bm{\mathsf{K}}}(S)=\textup{trace}\left(\bm{\mathsf{K}}_{:,S}(\bm{\mathsf{K}}_{S,S})^{-1}\bm{\mathsf{K}}_{S,:}\right)=\textup{trace}\left((\bm{\mathsf{K}}^{2})_{S,S}(\bm{\mathsf{K}}_{S,S})^{-1}\right).$
-    Maximizing $\mathcal{L}_{\bm{\mathsf{K}}}(S)$ over all possible subsets of $[m]$ of size $d$ is infeasible as in the DPP setting, so we compute approximate nuclear scores using random sketches of $\bm{\mathsf{K}}$ and $\bm{\mathsf{A}}$, and the identities
-    $\textup{diag}(\bm{\mathsf{K}}^{2})=\mathbb{E}_{\bm{\mathsf{x}}\sim\mathcal{N}(\bm{\mathsf{0}},\bm{\mathsf{I}})}[(\bm{\mathsf{Kx}})^{\odot 2}],\hskip 5.69054pt\textup{diag}(\bm{\mathsf{K}})=\mathbb{E}_{\bm{\mathsf{x}}\sim\mathcal{N}(\bm{\mathsf{0}},\bm{\mathsf{I}})}[(\bm{\mathsf{\bm{\mathsf{A}}\bm{\mathsf{x}}}})^{\odot 2}],$
-    where $\odot 2$ denotes element-wise squaring. Namely, we compute the averages
-    $\textup{diag}(\bm{\mathsf{K}}^{2})/\textup{diag}(\bm{\mathsf{K}})\approx\frac{(\bm{\mathsf{K}}\bm{\mathsf{\Omega}})^{\odot 2}\bm{\mathsf{1}}}{(\bm{\mathsf{A}}\bm{\mathsf{\Phi}})^{\odot 2}\bm{\mathsf{1}}}\Rightarrow\bm{\mathsf{K}}^{2}_{\ell,\ell}/\bm{\mathsf{K}}_{\ell,\ell}\approx\frac{\|[\bm{\mathsf{K}}\bm{\mathsf{\Omega}}]_{\ell,:}\|^{2}_{2}}{\|[\bm{\mathsf{A}}\bm{\mathsf{\Phi}}]_{\ell,:}\|^{2}_{2}},$
-    where $\bm{\mathsf{\Omega}},\bm{\mathsf{\Phi}}$ are Gaussian matrices with $O(\log m)$ columns.
-We then update $\bm{\mathsf{K}}$ and $\bm{\mathsf{A}}$ by Schur complements with respect to the selected index, and repeat d1d-1 times.
 We can express the action of randomized sampling as a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{\hskip 0.56905ptd\times m}$.
-The lower bounds on the minimal dimension required for a random sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) are summarized in Table 3, along with the computational complexity of computing the associated probabilities.
-Unlike the bounds in Table 1, the values of $d$ required for a randomized sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) depend not only on the rank $k$ of $\bm{\mathsf{A}}$ and the distortion parameter $\varepsilon$, but also, in some cases, on quantities associated with the matrix $\bm{\mathsf{A}}$ that are usually very expensive to compute.
+The lower bounds on the minimal dimension required for a random sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) are summarized in Table 3, along with the cost of computing the associated probabilities.
+The values of $d$ required for a randomized sampling matrix $\bm{\mathsf{\Gamma}}$ to satisfy (21) depend not only on the rank $k$ of $\bm{\mathsf{A}}$ and distortion $\varepsilon$, but also, in some cases, on quantities associated with the matrix $\bm{\mathsf{A}}$ that are usually very expensive to compute.
 For uniform sampling, the requisite embedding dimension $d$ depends on the coherence $\mu$ of $\bm{\mathsf{A}}$, defined as $\mu(\bm{\mathsf{A}})=\max_{1\leq i\leq m}\frac{n}{m}l_{i}$, or the maximum leverage score $l_{i}$ in (25).
-The coherence is a measure of how evenly information in the matrix is distributed; high coherence indicates that its key information is very localized and that uniform sampling may perform poorly as a result; see, e.g., Ipsen14 or (martinsson2020, Section 9.6.2).
-In spite of this drawback, uniform sampling has become a popular choice for large-scale linear algebra computations as a cost-effective oblivious subspace embedding (OSE) method.
+The coherence is a measure of how evenly information in the matrix is distributed; high coherence indicates that its key information is very localized and that uniform sampling may perform poorly as a result; see, e.g., Ipsen14.
+In spite of this drawback, uniform sampling is a popular choice for large-scale linear algebra computations as a cost-effective OSE.
 However, it is unclear how to best choose the embedding dimension $d$ in practice since the coherence is typically unavailable, cf.
@@
 For squared norm sampling, the value of $d$ depends on the condition number $\kappa$ of $\bm{\mathsf{A}}^{*}\bm{\mathsf{A}}$, defined as the ratio of its largest to smallest eigenvalue, which is not readily available in practice.
-The embedding dimension for leverage score sampling only depends on $k$ and $\epsilon$, but this independence comes at great computational expense to compute the sampling probabilities.
+The embedding dimension for leverage score sampling only depends on $k$ and $\epsilon$, but the sampling probabilities are very expensive.
 In particular, we note that leverage scores are as expensive to directly compute as $\texttt{svd}(\bm{\mathsf{A}})$, rendering their direct computations infeasible for most practical purposes.
@@
 AlaouiMahoney15; Chen21; DrineasMagdonMahoneyWoodruff12; MuscoMusco17.
-DPP sampling is known to have nearly optimal theoretical guarantees, cf.
-derezinski2020improved.
-However, naive implementations incur a cost of $O(m^{3})$, which makes DPP sampling practically infeasible.
-In derezinski2019exact, the DPP-VFX algorithm is presented, which, given access to $\bm{\mathsf{A}}\bm{\mathsf{A}}^{\top}$, samples exactly from a DPP distribution with a pre-processing cost of $O(mk^{10}+k^{15})$.
-In particular, the DPP-VFX algorithm was the first to achieve a cost of $O(m\cdot\textup{poly}(k))$ for the first sample and $O(\textup{poly}(k))$ for subsequent samples; however, there still exists a runtime bottleneck of $\Omega(m)$ due to the computation of $m$ marginals for the associated sampling distribution.
-The same authors improved on DPP-VFX with the $\alpha$-DPP algorithm of calandriello2020sampling, which avoids the $\Omega(m)$ bottleneck by incorporating uniform sub-sampling before computing marginals.
-While nuclear norm maximization is a greedy method of choosing columns based on posterior error minimization, the authors of Fornace2024 present a “matrix-free” algorithm for fast implementation with randomization.
-Its performance is comparable to that of DPP sampling, and nuclear norms could be adapted into a sampling distribution, so we include it in this section for side-by-side comparisons in Table 3.
-Namely, in Fornace2024, randomized sketching is used to compute the approximate scores that dictate column selection.
-We note that the requisite embedding dimension $d$ for nuclear norm maximization depends on the quantity $\nu=\frac{\textup{trace}(\bm{\mathsf{K}})-\textup{trace}^{(k)}(\bm{\mathsf{K}})}{\textup{trace}^{(k)}(\bm{\mathsf{K}})}$, where $\bm{\mathsf{K}}=\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$ and $\textup{trace}^{(k)}(\bm{\mathsf{K}})$ is the sum of its $k$ largest eigenvalues.
-For the complexity, we also include the dependence on the quantity $\phi$, which is the maximum cost of a mat-vec with $\bm{\mathsf{K}}$ or $\bm{\mathsf{A}}$, cf.
-(Fornace2024, Corollary 4.1).
+DPP sampling has nearly optimal theoretical guarantees, but naive implementations incur a cost of $O(m^{3})$ (after forming $\bm{\mathsf{A}}^{*}\bm{\mathsf{A}}$).
+In derezinski2019exact, the DPP-VFX algorithm is presented, which, given access to $\bm{\mathsf{A}}\bm{\mathsf{A}}^{*}$, samples exactly from a DPP distribution with a pre-processing cost of $O(mk^{10}+k^{15})$.
+In particular, the DPP-VFX algorithm was the first to achieve $O(m\cdot\textup{poly}(k))$ cost for the first sample and $O(\textup{poly}(k))$ for subsequent samples; however, there is still a runtime bottleneck of $\Omega(m)$ due to the $m$ marginals needed for the sampling distribution.
+The $\alpha$-DPP algorithm of calandriello2020sampling improves on DPP-VFX by incorporating uniform sub-sampling before computing marginals to avoid the $\Omega(m)$ bottleneck.
 Sampling distribution Dimension $d$ Complexity of distribution
@@
 […]
-Nuclear norm max $d=\Omega(k/\varepsilon+k-1)(\log(\nu^{-1})+\log(1/\varepsilon)))$ $O(md^{2}\log m+d\log m\psi)$
 Table 3: Summary of randomized sampling distributions commonly used to sample rows of rank-$k$ matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$.
-Theoretical lower bounds on the embedding dimension $d$ are given for $\bm{\mathsf{\Gamma}}$ to satisfy (21), as well as the asymptotic complexity of computing the associated probability distribution, cf.
+Theoretical lower bounds on the embedding dimension $d$ are given for (21), as well as the asymptotic complexity of computing the associated probability distribution, cf.
 DerezinskiMahoney2021; martinsson2020.
@@
 A natural idea is to apply a randomized DRM to $\bm{\mathsf{A}}$ and $\bm{\mathsf{b}}$ in (27), a strategy known as “sketch-and-solve” sarlos2006improved.
-Namely, we draw a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ and solve a smaller sketched problem
+Namely, we draw a randomized DRM $\bm{\mathsf{\Gamma}}\in\mathbb{F}^{d\times m}$ and solve
 $\widehat{\bm{\mathsf{x}}}=\arg\min_{\bm{\mathsf{x}}\in\mathbb{R}^{n}}\|\bm{\mathsf{\Gamma}}(\bm{\mathsf{A}}\bm{\mathsf{x}}-\bm{\mathsf{b}})\|_{2}.$ (28)
-If $\bm{\mathsf{\Gamma}}$ is a randomized subspace embedding for the $k$-dimensional column space of $\begin{bmatrix}\bm{\mathsf{A}}&\bm{\mathsf{b}}\end{bmatrix}$ with distortion $\varepsilon\in(0,1)$, then
+If $\bm{\mathsf{\Gamma}}$ embeds the $k$-dimensional column space of $\begin{bmatrix}\bm{\mathsf{A}}&\bm{\mathsf{b}}\end{bmatrix}$ with distortion $\varepsilon\in(0,1)$, then
 $(1-\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}_{*}-\bm{\mathsf{b}}\|_{2}\leq\|\bm{\mathsf{A}}\widehat{\bm{\mathsf{x}}}-\bm{\mathsf{b}}\|_{2}\leq(1+\varepsilon)\|\bm{\mathsf{A}}\bm{\mathsf{x}}_{*}-\bm{\mathsf{b}}\|_{2}$
@@
 Sketched GMRES (NakatsukasaTropp2024, Algorithm 1.1) includes a $k$-truncated Arnoldi process to quickly compute an approximate basis of the Krylov subspace.
-Sketch-and-precondition methods are also commonly used, but we note that these methods were recently shown to be numerically unstable when initialized with the zero vector, cf.
-MeierNakatsukasaTownsendWebb2024.
-With iterative sketching methods, sketched least squares problems are solved until converged, and these methods were recently proven to be forward stable in Epperly2024.
-We discuss alternative randomized DRMs for overdetermined least squares problems for computing tensor decompositions in Sect.
-6, which are based on the randomized algorithms for low-rank matrix decompositions that we discuss next.
+Sketch-and-precondition methods are also popular Avron2010; MeierNakatsukasaTownsendWebb2024; MengSaundersMahoney2014.
+We discuss alternative randomized DRMs for overdetermined least squares problems in computing tensor decompositions in Sect.
+6.
 4 Randomized Algorithms for Low-Rank Matrix Decompositions
 In this section, we illustrate several ways that randomized DRMs are often used to compute low-rank matrix decompositions.
-We first introduce the rangefinder problem and show how it can be solved with randomized DRMs in Sect.
+We introduce the rangefinder problem and show how it can be solved with randomized DRMs in Sect.
 4.1.
-We then explain how randomized DRMS can be used to estimate the error in the corresponding low-rank approximation, which can be utilized for adaptive rangefinding.
+We then explain how randomized DRMS can be used to estimate the error in the corresponding low-rank approximation for adaptive rangefinding.
 These ideas are foundational for randomized algorithms that have become ubiquitous for fast and reliable low-rank matrix approximation: the randomized SVD (Sect.
@@
 4.1 The Randomized Rangefinder
-At the heart of many linear algebraic tasks lies the rangefinder problem, the computation of a lower-dimensional space that sufficiently captures the action of a given matrix.
-More specifically, to solve the rangefinder problem for an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and target rank $k\leq\min(m,n)$, we compute an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns approximately span the column space of $\bm{\mathsf{A}}$.
-We then an obtain a rank-$k$ approximation given by $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$, with approximation error
+At the heart of many linear algebraic tasks lies the rangefinder problem, finding a lower-dimensional space that captures the action of a given matrix.
+To do this, for an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and target rank $k\leq\min(m,n)$, we compute an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns approximately span the column space of $\bm{\mathsf{A}}$.
+We then obtain a rank-$k$ approximation given by $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$, with approximation error
 $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|=\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|,$ (29)
-measured in either the Frobenius or spectral norm in our work.
-The Frobenius norm is often preferable for ease of analysis, but error bounds in the Frobenius norm can be less informative than bounds in the spectral norm, cf.
-(martinsson2020, Remark 2.1).
-The rangefinder problem can be handled effectively with randomized sketching as shown in Algorithm 1.
-This algorithm produces an orthonormal matrix $\bm{\mathsf{Q}}$ via QR factorization of an $m\times\ell$ random sketch of $\bm{\mathsf{A}}$ for desired embedding dimension $\ell$.
+measured in either the Frobenius or spectral norm in our work. The Frobenius norm is often preferable for ease of analysis but can be less informative the spectral norm.
+The rangefinder problem can be handled effectively with randomized DRMs, as in Algorithm 1, which produces an orthonormal matrix $\bm{\mathsf{Q}}$ via QR factorization of an $m\times\ell$ random sketch of $\bm{\mathsf{A}}$ for desired embedding dimension $\ell$.
 The total cost of Algorithm 1 includes the cost of simulating the $n\times\ell$ random test matrix, $O(mn\ell)$ operations to form the sketch, and $O(mk^{2})$ operations for the QR factorization.
@@
 5:Compute $\bm{\mathsf{Q}}=\texttt{col}(\bm{\mathsf{Y}},k)$.
-Implementation details
-There are many considerations to be made when implementing the randomized rangefinder algorithm in practice.
-We summarize the important details here and refer to halko2011; martinsson2020 for more thorough treatments.
-First, sketching, as opposed to sampling, should be used whenever possible.
-Gaussian matrices are popular for easy implementation and strong theoretical results, but structured random matrices (e.g., SparseStack) perform almost as well in practice.
-Second, the numerical accuracy of the rangefinder algorithm can be improved with power iteration Hestenes1952methods; IpsenWills2006; Saad1986GMRES or Krylov methods IpsenMeyer1998; Saad2022; SimonciniSzyld2003, in addition to numerically stable orthogonalization techniques like Householder reflections or rank-revealing QR Chan1987; Ilse1994.
-One last important detail concerns the subspace dimension.
-If the target rank $k$ of the low-rank approximation to $\bm{\mathsf{A}}$ is known a priori, then a randomized DRM of size $\ell=k+p$ is usually sufficient with a small amount of oversampling $p$, such as $p=5$.
-(We deal with the situation where $k$ is unknown in Sect.
-4.1.) For Gaussian test matrices with slight oversampling, (halko2011, Theorem 10.1) guarantees that the expected error of the randomized rangefinder is comparable to the best rank-$k$ approximation error from the truncated SVD of $\bm{\mathsf{A}}$, so long as its trailing singular values are small.
-This is generally achievable, for instance, by powering $\bm{\mathsf{A}}$ in the sketch $\bm{\mathsf{Y}}=(\bm{\mathsf{A}}\bm{\mathsf{A}}^{*})^{q}\bm{\mathsf{\Omega}}$ some small number of iterations $q$ RokhlinSzlamTygert2009.
-The theoretical results for Gaussian random matrices often furnish the performance heuristics for structured random sketching matrices, which behave similarly to Gaussians in practice despite the more pessimistic bounds that appear in their analysis (halko2011, Theorem 11.2).
-This claim is strengthened by the simple addition of a posteriori “certificates of accuracy” upon termination of Algorithm 1, as explained in the next section.
+We mention a few implementation details here and refer to halko2011; martinsson2020 for thorough treatments.
+For Gaussian sketching matrices with dimension $k+p$, (halko2011, Theorem 10.1) guarantees the expected error of the randomized rangefinder is comparable to the best rank-$k$ approximation error from the truncated SVD, for very small $p$, e.g., $p=5$, so long as the trailing singular values of $\bm{\mathsf{A}}$ are small.
+The theory for Gaussians often furnishes performance heuristics for other randomized DRMs, which frequently behave like Gaussians despite the more pessimistic theoretical bounds, cf.
+Table 1.
+The simple addition of a posteriori “certificates of accuracy” provides additional assurance.
 Error estimation and certificates of accuracy
-We now demonstrate how the approximation error (30) can be estimated reliably and efficiently with randomized sketching.
-This method of error estimation also serves as a “certificate of accuracy” for rangefinding, an assurance that the orthornomal basis given by $\bm{\mathsf{Q}}$ produces a satisfactory low-rank approximation of $\bm{\mathsf{A}}$.
-Suppose that we have an orthonormal matrix $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ whose columns potentially form a basis for the range of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, e.g., the output of Algorithm 1 for a given $k$.
-We estimate the approximation error $\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|$ using the following method.
+The approximation error (29) can be estimated reliably and efficiently with randomized sketching, which serves as a “certificate of accuracy” that the basis given by $\bm{\mathsf{Q}}$ approximately spans the range of $\bm{\mathsf{A}}$.
+Let $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ be an orthonormal matrix whose columns potentially form a basis for the range of $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$.
 Let $\bm{\mathsf{\Phi}}\in\mathbb{F}^{\hskip 0.56905ptn\times s}$ be a Gaussian matrix that is independent from any random matrix used to compute $\bm{\mathsf{Q}}$.
 In practice, $s$ is a small fixed value, such as $s=10$.
-We form a new random sketch, known as the auxiliary sample
+We form a sketch, often called the auxiliary sample,
 $\bm{\mathsf{Z}}=\bm{\mathsf{A}}\bm{\mathsf{\Phi}},$ (30)
-incurring only $s$ additional matrix-vector products with $\bm{\mathsf{A}}$. Using $\bm{\mathsf{Z}}$, we can obtain an inexpensive Gaussian sketch of the approximation error (30):
+From $\bm{\mathsf{Z}}$, we obtain an inexpensive Gaussian sketch of the approximation error (29):
 $(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{Z}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\bm{\mathsf{\Phi}}.$ (31)
-We then take advantage of randomized methods to estimate the norm of (32), such as randomized trace estimation for the Frobenius norm or randomized Krylov methods for the Schatten $2p$-norm (a proxy for the the spectral norm with sufficiently large $p$).
-Recall that the Schatten $p$-norm of a matrix is the $\ell_{p}$-norm of its singular values, so that the Schatten $\infty$-norm coincides with the spectral norm and the Schatten 2-norm with the Frobenius norm.
-To estimate the error in the Frobenius norm, we use the approximation
+We can then estimate the Frobenius norm of (31) as
 $\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|_{\text{F}}^{2}\approx\frac{1}{s}\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{Z}}\|_{\text{F}}^{2},$ (32)
-relying on the theory of randomized trace estimation (recall that $\textup{trace}(\bm{\mathsf{X}}^{*}\bm{\mathsf{X}})=\|\bm{\mathsf{X}}\|_{\text{F}}^{2}$).
-Namely, we exploit the fact that for any matrix $\bm{\mathsf{B}}$ and isotropic random vector $\omega$, $\text{trace}(\bm{\mathsf{B}})=\bm{\mathsf{\omega}}^{*}\bm{\mathsf{B}}\bm{\mathsf{\omega}}$ in expectation.
-To reduce the variance, we take the average of $s$ random variables in (33), using the auxiliary sample $\bm{\mathsf{Z}}$.
-Estimating the error in the spectral norm is not quite as simple.
-It is possible to construct unbiased estimators for the Schatten 2pp-norm for any positive integer $p$, as for the Frobenius norm above, but these estimators are expensive with large variance KongValiant2017; LiWoodruff2014.
-Randomized Krylov methods are often used instead to estimate the spectral norm LeykWoz1998; MartinssonRokhlinTygert2011; MuscoMusco2015; Tropp2022; WangZhangZhang2015.
-For example, we can estimate the spectral norm error with a randomized block Krylov method, letting $\bm{\mathsf{B}}=\begin{bmatrix}(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\end{bmatrix}\begin{bmatrix}(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\end{bmatrix}^{*}$,
-$\|(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\|^{2}_{2}\approx\max_{\bm{\mathsf{v}}\in\mathcal{K}(\bm{\mathsf{B}};\bm{\mathsf{\Phi}})}\frac{\bm{\mathsf{v}}^{*}\bm{\mathsf{B}}\bm{\mathsf{v}}}{\bm{\mathsf{v}}^{*}\bm{\mathsf{v}}},$ (34)
-where $\mathcal{K}(\bm{\mathsf{B}};\bm{\mathsf{\Phi}})=\textup{span}\begin{bmatrix}\bm{\mathsf{\Phi}}\ &\bm{\mathsf{B}}\bm{\mathsf{\Phi}}\ &\bm{\mathsf{B}}^{2}\bm{\mathsf{\Phi}}\ &\cdots&\bm{\mathsf{B}}^{q}\bm{\mathsf{\Phi}}\end{bmatrix}$ is a block Krylov subspace of depth $q$.
-In general, $q\approx\log m$ is necessary to achieve the fast convergence rate of $1/q^{2}$ for Krylov methods, but smaller values of $q$ can be used in practice for matrices with fast spectral decay.
-Although Krylov methods enjoy quick convergence, numerical instability is a known shortcoming, so careful implementation is needed, cf.
-MuscoMusco2015; Tropp2018; WangZhangZhang2015.
-Regardless of the choice of norm, an auxiliary sketch of the error with a Gaussian random test matrix $\bm{\mathsf{\Phi}}$ can provide us with an a posteriori “certificate of accuracy” for rangefinding (martinsson2020, Section 12.2).
-For instance, if a structured random matrix (e.g., SRTT, sparse sign) is used in Algorithm 1 instead of a Gaussian matrix, the computations are generally faster but lack formal guarantees on the accuracy of the computed solution.
-More generally, if we are unsure of the accuracy of a given solution to the rangefinder problem, we can take a small (independent) Gaussian sketch of the error as in (31).
-This auxiliary sketch is inexpensive and avails us of the powerful theory of Gaussian matrices, regardless of how the column space approximation was obtained.
-In the next section, we demonstrate how this method of error estimation can be utilized for adaptive rangefinding when the target rank is unknown a priori.
-Adaptive rangefinding
-Up to this point, we have assumed that the target approximation rank could be passed as an argument to the randomized rangefinder algorithm.
-Unfortunately, a satisfactory target rank for the approximation is frequently unknown a priori.
-In that case, we can implement the randomized rangefinder algorithm adaptively, using the error estimation techniques of the previous section to approximate the range of $\bm{\mathsf{A}}$ to a desired precision $\varepsilon$, forming $\bm{\mathsf{Q}}$ incrementally until that precision is achieved.
-For ease of notation, let norm_est($\bm{\mathsf{Z}}$) denote the rangefinder error estimate obtained by one of the techniques in Sect.
-4.1, e.g., (33).
-An adaptive version of the randomized rangefinder is given in Algorithm 2, which we quickly summarize.
+noting that $\text{trace}(\bm{\mathsf{B}})=\bm{\mathsf{\omega}}^{*}\bm{\mathsf{B}}\bm{\mathsf{\omega}}$ in expectation for any matrix $\bm{\mathsf{B}}$ and isotropic random vector $\omega$).
+Then to reduce the variance, we take the average of $s$ random variables in (32), using the auxiliary sample $\bm{\mathsf{Z}}$.
+The error sketch avails us of the powerful theory of Gaussian matrices, regardless of how the column space approximation was obtained.
+For spectral norm estimation, see (martinsson2020, Sects.
+5-6).
+Rank-adaptive rangefinding
+We have so far assumed the target approximation rank can be passed as an argument to the randomized rangefinder.
+Unfortunately, the target rank for an approximation is often unknown a priori.
+The method of error estimation above also enables a rank-adaptive randomized rangefinder algorithm (Algorithm 2).
+For ease of notation, let norm_est($\bm{\mathsf{Z}}$) denote the rangefinder error estimated as in (32).
 Algorithm 2 Adaptive Randomized Rangefinder
@@
 […]
-7:while $\texttt{norm\_est}(\bm{\mathsf{Y}})>\varepsilon$ do $\triangleright$ Error estimation of Sect. 4.1
+7:while $\texttt{norm\_est}(\bm{\mathsf{Y}})>\varepsilon$ do $\triangleright$ Error estimation as in (32)
 8:  $i=i+1$
@@
 […]
-Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and let $b$ be the number of columns to be processed at a time.
-The choice of block size $b$ affects the computational efficiency of the algorithm.
-If $b$ is too small, we are not fully reaping the benefits of blocked matrix operations that help reduce memory traffic and avoid communication bottlenecks Heavner2022; Orti1998.
-With a larger block size $b$, more floating-point operations can be cast as matrix-matrix multiplication (BLAS-3) vs.
-matrix-vector multiplication (BLAS-2) Blackford2002; Dongarra1990; Gallivan1987; Higham1990.
-However, a larger block size runs the risk of overshooting the target rank, meaning we have done more work than necessary for rangefinding.
-In practice, a block size $10\leq b\leq 100$ is usually appropriate.
-In each iteration of the while-loop in Algorithm 2, the matrix $\bm{\mathsf{Y}}$ holds a sample of the residual error after the execution of line 9:
+Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ and let $b$ be the number of columns processed at a time.
+In practice, $10\leq b\leq 100$ is usually appropriate.
+In each iteration of Algorithm 2, after line 9,
 $\bm{\mathsf{Y}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}\bm{\mathsf{\Omega}},$ (33)
-where $\bm{\mathsf{Q}}$ is the basis that has been built up to the current iteration.
-Critically, the random test matrix $\bm{\mathsf{\Omega}}$ used in (35) is independent from the random matrices used to compute the current basis $\bm{\mathsf{Q}}$.
-We note that to compute the very first residual error norm_est($\bm{\mathsf{Y}}$) in line 5, we can draw a small auxiliary sample that then can be recycled in the second batch of samples in line 8.
-Variations of Algorithm 2 have been designed to optimize performance for specific types of input matrices.
-If the input matrix $\bm{\mathsf{A}}$ is small enough to store in RAM, it can be explicitly updated in each iteration via $\bm{\mathsf{A}}=(\bm{\mathsf{I}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*})\bm{\mathsf{A}}$ as in (martinsson2020, Algorithm 14).
-Alternatively, the version of YuGuLiLiuLi2017 effectively handles input matrices stored out-of-core, as well as sparse input matrices.
-If the spectrum of the input matrix decays slowly, power iteration can also be incorporated into Algorithm 2 to improve performance MartinssonVoronin2016.
+where $\bm{\mathsf{Q}}$ is the currently constructed basis. Critically, the random test matrix $\bm{\mathsf{\Omega}}$ in (33) is independent of the random matrices used to compute $\bm{\mathsf{Q}}$.
 4.2 Randomized SVD
-Now equipped with the randomized rangefinder, we present a foundational algorithm for low-rank matrix approximation known as the randomized SVD halko2011.
-Pseudocode is provided in Algorithm 3.
-We quickly go through its steps to show that the error of the resulting approximate SVD is precisely the randomized rangefinder error.
+The randomized SVD of halko2011 (Algorithm 3) is an important application of rangefinding, with approximation error equal to the randomized rangefinder error.
 Algorithm 3 Randomized SVD
@@
 6:Form $\bm{\mathsf{U}}=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$
-Given an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, our goal is to compute an approximate rank-$k$ truncated singular value decomposition
-$\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ (36)
-where $\bm{\mathsf{U}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ and $\bm{\mathsf{V}}\in\mathbb{F}^{\hskip 0.56905ptn\times k}$ are orthonormal matrices, and $\bm{\mathsf{\Sigma}}\in\mathbb{F}^{\hskip 0.56905ptk\times k}$ is a diagonal matrix of entries approximating the largest singular values of $\bm{\mathsf{A}}$.
-Suppose that $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ is a solution to the rangefinder problem (not necessarily from Algorithm 1), which corresponds to a rank-$k$ approximation of $\bm{\mathsf{A}}$ given by $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$.
-Let $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptk\times n}$ and compute singular value decomposition
-$\bm{\mathsf{B}}=\underset{k\times k}{\widehat{\bm{\mathsf{U}}}}\ \underset{k\times k}{\bm{\mathsf{\Sigma}}}\ \underset{k\times n}{\bm{\mathsf{V}}^{*}}.$ (37)
-We observe that
-$\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ $=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{B}}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|,$ (38)
-for $\bm{\mathsf{U}}:=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$; i.e.
-the approximation error of $\bm{\mathsf{A}}\approx\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$ is exactly the approximation error of $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$.
-Thus, if $\bm{\mathsf{Q}}=\texttt{RandomizedRangefinder}(\bm{\mathsf{A}},k)$ is the rangefinder solution computed by Algorithm 1, the accuracy of the randomized SVD is equal to that of the randomized rangefinder.
-In particular, this means that the error estimation techniques of Sect.
-4.1 can also be used to estimate the error of the randomized SVD.
-Then if the target rank $k$ is unknown a priori, the adaptive randomized rangefinder (Algorithm 2) can be carried out first, which determines $k$ in the randomized SVD.
-The total cost of Algorithm 3 includes the cost of the randomized rangefinder, plus $O(mnk)$ floating-point operations for $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$ in line 2.
-We note that the cost of the SVD in line 3 is just $O(k^{2}n)$, so the cost of Algorithm 3 after rangefinding is usually dominated by the matrix-matrix multiplication.
-Executing Algorithm 3 requires storage for $O((m+n)k)$ numbers.
-Recent work has also been done on cost-effective a posteriori error estimation methods that use only the information at hand for computing the low-rank decomposition.
-In EpperlyTropp2024, an algorithm for a posteriori error estimation is presented for the randomized SVD in a matrix-free setting, where the input matrix $\bm{\mathsf{A}}$ is only accessible via mat-vecs.
-Given the rank-$k$ randomized SVD $\bm{\mathsf{X}}=\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\in\mathbb{F}^{m\times n}$ corresponding to Gaussian random matrix $\bm{\mathsf{\Omega}}=\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\bm{\mathsf{\omega}}_{2}&\ldots&\bm{\mathsf{\omega}}_{k}\end{bmatrix}$, we want to estimate the mean-square error
+Given an input matrix $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$, our goal is to compute an approximate rank-$k$ truncated singular value decomposition $\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}.$ Suppose that $\bm{\mathsf{Q}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ is the output of Algorithm 1, corresponding to rank-$k$ approximation $\bm{\mathsf{A}}\approx\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$.
+Let $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptk\times n}$ and compute its SVD $\bm{\mathsf{B}}=\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$.
+We observe that for $\bm{\mathsf{U}}:=\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}$,
+$\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ $=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{B}}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\widehat{\bm{\mathsf{U}}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|=\|\bm{\mathsf{A}}-\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\|.$ (34)
+Thus, the approximation error of $\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}$ is exactly the approximation error of $\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$.
+If the target rank $k$ is unknown a priori, Algorithm 2 can be used instead.
+The total cost of Algorithm 3 includes the cost of rangefinding, plus $O(mnk)$ floating-point operations for $\bm{\mathsf{B}}=\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}$ in line 2.
+We note the cost of the SVD in line 3 is only $O(k^{2}n)$.
+The storage cost is $O((m+n)k)$.
+The recent work of EpperlyTropp2024 introduces a cost-effective a posteriori error estimation method for the randomized SVD.
+Given rank-$k$ randomized SVD $\bm{\mathsf{X}}=\bm{\mathsf{U}}\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*}\in\mathbb{F}^{m\times n}$ corresponding to Gaussian DRM $\bm{\mathsf{\Omega}}=\begin{bmatrix}\bm{\mathsf{\omega}}_{1}&\ldots&\bm{\mathsf{\omega}}_{k}\end{bmatrix}$, the mean-square error is defined
 $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})=\mathbb{E}\|\bm{\mathsf{A}}-\bm{\mathsf{X}}_{k-1}\|^{2}_{\text{F}},$
-where $\bm{\mathsf{X}}_{k-1}\in\mathbb{F}^{m\times(k-1)}$ is a rank-(k1)(k-1) randomized SVD.
-By (EpperlyTropp2024, Theorem 2.1), the following is an unbiased estimator for $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})$, which we use as a proxy for the mean-square error of the rank-$k$ approximation $\bm{\mathsf{X}}$, called the leave-one-out error estimator:
+where $\bm{\mathsf{X}}_{k-1}$ is a rank-(k1)(k-1) randomized SVD.
+By (EpperlyTropp2024, Theorem 2.1), the following is an unbiased estimator for $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})$, called the leave-one-out error estimator:
 $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})=\mathbb{E}\left[\frac{1}{k}\sum_{j=1}^{k}\|(\bm{\mathsf{A}}-\bm{\mathsf{X}}^{(j)})\bm{\mathsf{\omega}}_{j}\|^{2}\right],$ (35)
-where $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{Q}}^{(j)}(\bm{\mathsf{Q}}^{(j)})^{*}\bm{\mathsf{A}}$ is called a replicate, formed from the matrix $\bm{\mathsf{Q}}^{(j)}$ outputted by the randomized rangefinder using $\bm{\mathsf{\Omega}}$ without its $j$-th column $\bm{\mathsf{\omega}}_{j}$.
-Implemented naively, the right-hand side of (39) is too expensive for practical use, as it would require $k$ additional randomized rangefinder applications.
+where $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{Q}}^{(j)}(\bm{\mathsf{Q}}^{(j)})^{*}\bm{\mathsf{A}}$ is called a replicate, formed from the matrix $\bm{\mathsf{Q}}^{(j)}$ outputted by Algorithm 1 using $\bm{\mathsf{\Omega}}$ without its $j$-th column $\bm{\mathsf{\omega}}_{j}$.
+Implemented as written, the right-hand side of (35) is too expensive for practical use.
 However, by (EpperlyTropp2024, Equation 4.2),
@@
 where $\bm{\mathsf{t}}_{j}$, $j=1,\ldots,k$, are the normalized columns of $(\bm{\mathsf{R}}^{*})^{-1}$ in the QR factorization defined by $\bm{\mathsf{Q}}$ in step 1 of the randomized SVD (now also storing the $\bm{\mathsf{R}}$ matrix).
-Then we can obtain each replicate $\bm{\mathsf{X}}^{(j)}$ inexpensively as
-$\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{U}}(\bm{\mathsf{I}}-\widehat{\bm{\mathsf{U}}}^{*}\bm{\mathsf{t}}_{j}\bm{\mathsf{t}}_{j}^{*}\widehat{\bm{\mathsf{U}}})\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ (41)
-which allows us to compute (39) without any additional mat-vecs or applications of Algorithm 3, via
+Then we can obtain each replicate from (36) inexpensively as $\bm{\mathsf{X}}^{(j)}=\bm{\mathsf{U}}(\bm{\mathsf{I}}-\widehat{\bm{\mathsf{U}}}^{*}\bm{\mathsf{t}}_{j}\bm{\mathsf{t}}_{j}^{*}\widehat{\bm{\mathsf{U}}})\bm{\mathsf{\Sigma}}\bm{\mathsf{V}}^{*},$ which allows us to compute (35) without any additional rangefinding via
 $\textup{MSE}(\bm{\mathsf{X}}_{k-1},\bm{\mathsf{A}})\approx\frac{1}{k}\sum_{j=1}^{k}\left\|[(\bm{\mathsf{R}}^{*})^{-1}](:,j)\right\|^{-2}.$
@@
 In both the randomized rangefinder and the randomized SVD, orthonormal matrices are computed whose columns form an approximate basis for the row or column space of an input matrix $\bm{\mathsf{A}}$.
-The entries of these orthonormal matrices are linear combinations of the entries of $\bm{\mathsf{A}}$.
-However, for some problems (e.g., if $\bm{\mathsf{A}}$ is sparse or non-negative), it may be preferable to have a “natural basis,” consisting of subsets of rows or columns of $\bm{\mathsf{A}}$, as in the ID/CUR factorizations of Sect.
-2.4.
-In this section, we provide an overview of randomized methods to compute approximate ID/CUR factorizations based on the randomized rangefinder, including recent work to improve skeleton selection and performance DongRBRP2025; dong2022; Pearce2025; Pritchard2025.
-With the randomized rangefinder
-Recall from Sect.
-2.4 that pivoting algorithms, such as column-pivoted QR (CPQR) or LU with partial pivoting, are frequently used in the deterministic setting to select skeletons for an ID/CUR factorization.
-CPQR produces skeletons that are close to optimal in practice Cheng2005; martinsson2006interpolation.
-In particular, this is true when CPQR is applied to a random sketch of $\bm{\mathsf{A}}$ with appropriate embedding dimension.
-It is also conjectured in (trefethen1990, Remark 4) and exhibited empirically in dong2022; Pearce2025 that randomized sketching equips LUPP with near-optimal skeleton selection capability as well, even in adversarial cases.
-We illustrate how randomized sketching with CPQR or LUPP can be used to compute a row ID, then discuss alternate skeleton selection strategies that are rank-adaptive or based on sampling instead of sketching.
-More specifically, let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be an input matrix of exact rank $k$ for notational convenience.
-Let $\bm{\mathsf{Y}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ have columns that span the range of $\bm{\mathsf{A}}$, e.g., the output of the rangefinder algorithm.
-By definition,
-$\bm{\mathsf{A}}=\underset{m\times k}{\bm{\mathsf{Y}}}\ \underset{k\times n}{\bm{\mathsf{F}}}$
-for some matrix $\bm{\mathsf{F}}$.
-Compute a row ID of $\bm{\mathsf{Y}}$, either using CPQR via $\sim,\sim,I]=\texttt{qr}(\bm{\mathsf{Y}}^{*},k)$ as in Sect.
-2.4 or LUPP via $\sim,\sim,I]=\texttt{lu}(\bm{\mathsf{Y}},k)$ as in Sect.
-2.4, to obtain
-$\bm{\mathsf{Y}}=\underset{m\times k}{\bm{\mathsf{X}}}\underset{k\times k}{\bm{\mathsf{Y}}(I,:)}.$ (42)
+However, for some problems (e.g., if $\bm{\mathsf{A}}$ is sparse or non-negative), it may be preferable to have a “natural basis,” consisting of rows or columns of $\bm{\mathsf{A}}$, as in the ID/CUR of Sect.
+2.5.
+In this section, we provide an overview of randomized methods to compute approximate ID/CUR factorizations, including recent work to improve skeleton selection and performance DongRBRP2025; dong2022; Pearce2025; Pritchard2025.
+We first illustrate how a random sketch can be used as a proxy for the input matrix in ID/CUR approximations.
+Let $\bm{\mathsf{A}}\in\mathbb{F}^{\hskip 0.56905ptm\times n}$ be of exact rank $k$ for notational convenience.
+Let $\bm{\mathsf{Y}}\in\mathbb{F}^{\hskip 0.56905ptm\times k}$ have columns that span the range of $\bm{\mathsf{A}}$, not necessarily orthogonal.
+By definition, $\bm{\mathsf{A}}={\bm{\mathsf{Y}}}{\bm{\mathsf{F}}}$ for some matrix $\bm{\mathsf{F}}\in\mathbb{F}^{k\times n}$.
+Compute a row ID of $\bm{\mathsf{Y}}$, e.g., with CPQR via $\sim,\sim,I]=\texttt{qr}(\bm{\mathsf{Y}}^{*},k)$ or LUPP via $\sim,\sim,I]=\texttt{lu}(\bm{\mathsf{Y}},k)$, so that $\bm{\mathsf{Y}}=\bm{\mathsf{X}}\bm{\mathsf{Y}}(I,:)$.
 We now observe that index set $I$ and row interpolation matrix $\bm{\mathsf{X}}$ form a row ID of $\bm{\mathsf{A}}$:
 $\bm{\mathsf{A}}=\bm{\mathsf{Y}}\bm{\mathsf{F}}=\left[\bm{\mathsf{X}}\bm{\mathsf{Y}}(I,:)\right]\bm{\mathsf{F}}=\bm{\mathsf{X}}\left[\bm{\mathsf{Y}}(I,:)\bm{\mathsf{F}}\right]=\bm{\mathsf{X}}\bm{\mathsf{A}}(I,:).$
-Thus, to compute a row ID of $\bm{\mathsf{A}}$, it suffices to compute a row ID of a matrix $\bm{\mathsf{Y}}$ whose columns span the range of $\bm{\mathsf{A}}$.
-We can construct such a matrix $\bm{\mathsf{Y}}\in\mathbb{F}^{m\times\ell}$ for a rank-$k$ approximation of $\bm{\mathsf{A}}$ using the randomized rangefinder with high probability with, e.g., $\ell=k+10.$ Additionally, since we do not need $\bm{\mathsf{Y}}$ to be an orthonormal matrix here, we can save on computational cost by observing that the range of $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is itself an accurate approximation to the range of $\bm{\mathsf{A}}$ with high probability (halko2011, Theorems 10.5 and 10.6).
-The randomized rangefinder error $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ using $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is equal to $\|\bm{\mathsf{A}}-\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}\bm{\mathsf{A}}\|$, where $\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}$ is the orthogonal projector onto the column space of $\bm{\mathsf{Y}}$ or, equivalently, the column space of $\bm{\mathsf{Q}}$.
-Pseudocode for this randomized row ID algorithm (with sketched CPQR) is provided in Algorithm 4.
+Thus, to compute a row ID of $\bm{\mathsf{A}}$, it suffices to compute a row ID of a matrix $\bm{\mathsf{Y}}$ whose columns span the range of $\bm{\mathsf{A}}$, and since we do not need $\bm{\mathsf{Y}}$ to be orthogonal, we can save on computational cost by using the fact that the range of $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is itself an accurate approximation to the range of $\bm{\mathsf{A}}$ with high probability (halko2011, Theorems 10.5 and 10.6).
+We observe that the randomized rangefinder error $\|\bm{\mathsf{A}}-\bm{\mathsf{Q}}\bm{\mathsf{Q}}^{*}\bm{\mathsf{A}}\|$ using $\bm{\mathsf{Y}}=\bm{\mathsf{A}}\bm{\mathsf{\Omega}}$ is equal to $\|\bm{\mathsf{A}}-\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}\bm{\mathsf{A}}\|$, where $\bm{\mathsf{Y}}\bm{\mathsf{Y}}^{\dagger}$ is the orthogonal projector onto the column space of $\bm{\mathsf{Y}}$ (or, equivalently, the column space of $\bm{\mathsf{Q}}$).
+Pseudocode for this randomized row ID procedure (with sketched CPQR) is provided in Algorithm 4.
 Algorithm 4 Randomized Row ID with Sketched CPQR
@@
 7:Set $\bm{\mathsf{X}}(P,:)=\begin{bmatrix}\bm{\mathsf{I}}\ &\ \ \bm{\mathsf{R}}_{1:k,1:k}^{-1}\bm{\mathsf{R}}_{1:k,k+1:m}\end{bmatrix}^{*}$
-Adaptive algorithms for randomized ID/CUR
+With rank-adaptive randomized sketching and greedy pivoting
 Frequently in applications, the target rank for a low-rank approximation is unknown a priori, in which case we turn to rank-adaptive methods.
 If using CPQR for skeleton selection, the adaptive version of the randomized rangefinder may be used to build an orthonormal basis incrementally from independent random sketches or samples.
-However, if we are using LUPP as the skeleton selection method for improved efficiency, we want to estimate the approximation error without computing an orthonormal basis.
+However, if we are using LUPP as the skeleton selection method for improved efficiency and better parallelizability, we want to estimate the approximation error without computing an orthonormal basis.
 We briefly illustrate how a row ID can be constructed with the rank-adaptive, LUPP-based method of Pearce2025.
@@
 $\bm{\mathsf{C}}^{\dagger}\bm{\mathsf{A}}\bm{\mathsf{R}}^{\dagger}$ as in Sect.
-2.4.) The column residual is initialized as a row sketch, $\bm{\mathsf{S}}^{(C)}_{0}=\bm{\mathsf{G}}\bm{\mathsf{A}}\in\mathbb{F}^{1.1b\times n}$, where typically $5\leq b\leq 250$, and the row and column skeletons are empty sets.
+2.5.) The column residual is initialized as a row sketch, $\bm{\mathsf{S}}^{(C)}_{0}=\bm{\mathsf{G}}\bm{\mathsf{A}}\in\mathbb{F}^{1.1b\times n}$, where typically $5\leq b\leq 250$, and the row and column skeletons are empty sets.
 While the error criterion is unmet, column skeletons $J$ are first computed from $\bm{\mathsf{S}}^{(C)}_{0}$, and the row residual is formed as $\bm{\mathsf{S}}^{(R)}_{1}=\bm{\mathsf{A}}(:,J)-\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}\bm{\mathsf{R}}_{1}(:,J)=\bm{\mathsf{A}}(:,J)$.
@@
 This procedure is repeated, with new skeletons appended to previously selected ones, until the desired approximation accuracy is achieved.
-Algorithms based on coordinate sampling
+With rank-adaptive random pivoting
 There have been several recent advances in coordinate sampling-based methods for ID/CUR decompositions.
 In ChenEpperlyTroppWebber2025, the randomly pivoted (RP) Cholesky algorithm is introduced for positive semi-definite matrices $\bm{\mathsf{A}}\in\mathbb{F}^{m\times m}$.
-Given a target rank $k$, the algorithm outputs a set of pivot indices $S=\{s_{1},\ldots,s_{k}\}$ akin to skeletons, and a matrix $\bm{\mathsf{F}}\in\mathbb{F}^{m\times k}$ that defines the Nyström approximation $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{F}}\bm{\mathsf{F}}^{*}$.
+Given a target rank $k$, the algorithm outputs a set of pivot indices $S=\{s_{1},\ldots,s_{k}\}$ akin to skeletons, and a matrix $\bm{\mathsf{F}}\in\mathbb{F}^{m\times k}$ for approximate Cholesky factorization $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{F}}\bm{\mathsf{F}}^{*}$, which is the factorized form of the column Nystr”om approximation $\widehat{\bm{\mathsf{A}}}=\bm{\mathsf{A}}(:,S)\bm{\mathsf{A}}(S,S)^{{\dagger}}\bm{\mathsf{A}}(S,:)$.
+By selecting pivots, or skeletons, with probability proportional to the magnitudes of the diagonal entries of the residual matrix in each iteration, the RP Cholesky algorithm returns the Cholesky-factorized form of the Nyström approximation in $O(km)$ matrix entry evaluations and $O(k^{2}m)$ additional arithmetic operations, with $O(km)$ storage requirements.
 We begin by defining $\bm{\mathsf{d}}=\textup{diag}(\bm{\mathsf{A}})$ and initializing the matrix $\bm{\mathsf{F}}$ as $\bm{\mathsf{0}}$.
@@
 5 Tensor Preliminaries
-This section marks the beginning of our exploration of tensors.
+This section marks the beginning of our exploration of tensors, where we present tensors as natural higher-order generalizations of matrices, from 2-way to $d$-way arrays.
 We begin with formal notation and definitions in Sect.
-5.1, presenting tensors as natural higher-order generalizations of matrices, from 2-way to $d$-way arrays.
-As a result, many of the randomized methods in the previous sections developed for matrices can be extended to tensors, though there are subtleties with higher-order inputs, and many concepts can be obscured by the unwieldy notation that frequently accompanies them.
-We try to mitigate this by providing visualizations, including a table of notation (Table 4) and illustrations of the matrix and tensor products in Sect.
+5.1 and discuss important matrix and tensor operations in Sect.
 5.2.
-In Sect.
-5.3, we point out some peculiarities specific to higher-order tensors, such as tensor rank and uniqueness of decompositions.
 While there are many different ways to represent tensors, we focus on two fundamental tensor decomposition formats: the canonical polyadic (CP) and Tucker.
 In Sect.
-5.3, we lay out key concepts and practicalities for each of them, including algorithms commonly used to compute them.
+5.3, we lay out key concepts and practicalities for each of them, including algorithms to compute them.
 We then discuss particular low-rank tensor decompositions that can be represented in each format, such as higher-order extensions of the matrix SVD and ID/CUR.
-Throughout this section, we try to emphasize that the classical matrix methods covered in previous sections underlie those for tensors.
-Symbol Definition
-$\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$ Tensor with $d$ modes
-$\|\bm{\mathcal{X}}\|$ Frobenius norm $\left(\sum_{i_{1}=1}^{N_{1}}\ldots\sum_{i_{d}=1}^{N_{d}}\bm{\mathcal{X}}(i_{1},\ldots,i_{d})^{2}\right)^{1/2}$
-$N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$ Number of entries of $\bm{\mathcal{X}}$
-$N^{(-j)}=\prod_{k\neq j}N_{j}$ Number of entries of $\bm{\mathcal{X}}$ excluding mode $j$
-$\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$ Mode-$j$ unfolding of $\bm{\mathcal{X}}$
-$\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}$ Kronecker product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J},\bm{\mathsf{B}}\in\mathbb{F}^{K\times L}$
-$\bm{\mathsf{A}}\odot\bm{\mathsf{B}}$ Khatri-Rao product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times K},\bm{\mathsf{B}}\in\mathbb{F}^{J\times K}$
-$\bm{\mathsf{A}}*\bm{\mathsf{B}}$ Hadamard product of $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J},\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$
-Table 4: Summary of tensor notation.
+Throughout this section, we show how the classical matrix methods of Sect.
+2 form the foundation of those for tensors.
 5.1 Notation and Definitions
@@
 The mode-$j$ slices of a tensor are the sub-tensors $\bm{\mathcal{X}}(:,\ldots,:,i_{j},:,\ldots,:)$ obtained by fixing a single index $i_{j}\in[N_{j}]$, e.g., slices of 3-mode tensors are matrices.
-We define $N^{(-j)}=\prod_{k=1,k\neq j}^{d}N_{k}$ as the product of all dimensions except mode $j$, and $N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$.
+We define $N=\prod_{j=1}^{d}N_{j}=\textup{size}(\bm{\mathcal{X}})$, and $N^{(-j)}=\prod_{k=1,k\neq j}^{d}N_{k}=N/N_{j}$.
 Through a rearrangement of entries known as matricization, a $d$-mode tensor can be “unfolded” into a matrix.
 The mode-$j$ unfolding of a $d$-mode tensor $\bm{\mathcal{X}}$, denoted $\bm{\mathsf{X}}_{(j)}\in\mathbb{F}^{N_{j}\times N^{(-j)}}$, $j=1,\ldots,d$, is a matrix whose columns are the mode-$j$ fibers of $\bm{\mathcal{X}}$.
-Throughout this section, we assume that the $d$ modes are ordered $1,\ldots,d$ for notational simplicity, but for computational reasons, it may be beneficial to re-order the modes for processing.
+We assume the $d$ modes are ordered $1,\ldots,d$ for simplicity, but for computational reasons, it may be beneficial to re-order the modes for processing.
 In our work, the (Frobenius) norm of a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times\cdots\times N_{d}}$, denoted by $\|\bm{\mathcal{X}}\|$, is the square root of the sum of squares of all elements, analogous to the Frobenius norm for matrices.
@@
 $(\bm{\mathsf{A}}\otimes\bm{\mathsf{B}})(K(i-1)+k,L(j-1)+\ell)=\bm{\mathsf{A}}(i,j)\bm{\mathsf{B}}(k,\ell).$
-The Kronecker product then takes the form
-$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,J}\end{bmatrix}}\otimes\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,L}\\ \vdots&\ddots&\vdots\\ b_{K,1}&\cdots&b_{K,L}\end{bmatrix}}$ $=\underset{\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&\cdots&a_{1,1}b_{1,L}&\cdots\cdots&a_{1,J}b_{1,1}&\cdots&a_{1,J}b_{1,L}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ a_{1,1}b_{K,1}&\cdots&a_{1,1}b_{K,L}&\cdots\cdots&a_{1,J}b_{K,1}&\cdots&a_{1,J}b_{K,L}\\ \vdots&&&&&&\vdots\\ \vdots&&&\ddots&&&\vdots\\ \vdots&&&&&&\vdots\\ a_{I,1}b_{1,1}&\cdots&a_{I,1}b_{1,L}&\cdots\cdots&a_{I,J}b_{1,1}&\cdots&a_{I,J}b_{1,L}\\ \vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ a_{I,1}b_{K,1}&\cdots&a_{I,1}b_{K,L}&\cdots\cdots&a_{I,J}b_{K,1}&\cdots&a_{I,J}b_{K,L}\end{bmatrix}}$
-$=\begin{bmatrix}a_{1,1}\bm{\mathsf{B}}&\cdots&a_{1,J}\bm{\mathsf{B}}\\ \vdots&\ddots&\vdots\\ a_{I,1}\bm{\mathsf{B}}&\cdots&a_{I,J}\bm{\mathsf{B}}\end{bmatrix}$.
 The Khatri-Rao product of matrices $\bm{\mathsf{A}}\in\mathbb{F}^{I\times K}$ and $\bm{\mathsf{B}}\in\mathbb{F}^{J\times K}$, denoted $\bm{\mathsf{A}}\odot\bm{\mathsf{B}}$, is the $(IJ)\times K$ matrix with entries
 $(\bm{\mathsf{A}}\odot\bm{\mathsf{B}})(:,k)=\bm{\mathsf{A}}(:,k)\otimes\bm{\mathsf{B}}(:,k)$
-for $k=1,\ldots,K.$ The Khatri-Rao product takes the form
-$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,K}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,K}\end{bmatrix}}\odot\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,K}\\ \vdots&\ddots&\vdots\\ b_{J,1}&\cdots&b_{J,K}\end{bmatrix}}$ $=\underset{\bm{\mathsf{A}}\odot\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&a_{1,2}b_{1,2}&\cdots&a_{1,K}b_{1,K}\\ \vdots&\vdots&\ddots&\vdots\\ a_{1,1}b_{J,1}&a_{1,2}b_{J,2}&\cdots&a_{1,K}b_{J,K}\\ \vdots&\vdots&&\vdots\\ \vdots&\vdots&\ddots&\vdots\\ \vdots&\vdots&&\vdots\\ a_{I,1}b_{1,1}&a_{I,2}b_{1,2}&\cdots&a_{I,K}b_{1,K}\\ \vdots&\vdots&\ddots&\vdots\\ a_{I,1}b_{J,1}&a_{I,2}b_{J,2}&\cdots&a_{I,K}b_{J,K}\\ \end{bmatrix}}$
-$=\begin{bmatrix}\bm{\mathsf{a}}_{1}\otimes\bm{\mathsf{b}}_{1}&\bm{\mathsf{a}}_{2}\otimes\bm{\mathsf{b}}_{2}&\cdots&\bm{\mathsf{a}}_{K}\otimes\bm{\mathsf{b}}_{K}\end{bmatrix}$,
-where $\bm{\mathsf{a}}_{i}$ and $\bm{\mathsf{b}}_{i}$ are the columns of $\bm{\mathsf{A}}$ and $\bm{\mathsf{B}}$.
-There is a useful property relating the Kronecker and Khatri-Rao products: if $\bm{\mathsf{A}}\in\mathbb{F}^{K\times J}$, $\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$, $\bm{\mathsf{C}}\in\mathbb{F}^{I\times K}$, and $\bm{\mathsf{D}}\in\mathbb{F}^{J\times I}$,
-$(\bm{\mathsf{C}}\otimes\bm{\mathsf{D}})(\bm{\mathsf{A}}\odot\bm{\mathsf{B}})=(\bm{\mathsf{C}}\bm{\mathsf{A}})\odot(\bm{\mathsf{D}}\bm{\mathsf{B}}).$
-Additionally we have the following inequality of matrix norms:
-$\|\bm{\mathsf{A}}\odot\bm{\mathsf{B}}\|^{2}_{\text{F}}\leq\|\bm{\mathsf{A}}\otimes\bm{\mathsf{B}}\|^{2}=\|\bm{\mathsf{A}}\|^{2}\|\bm{\mathsf{B}}\|^{2}.$
+for $k=1,\ldots,K.$
 The Hadamard product of matrices $\bm{\mathsf{A}}\in\mathbb{F}^{I\times J}$ and $\bm{\mathsf{B}}\in\mathbb{F}^{I\times J}$, denoted $\bm{\mathsf{A}}*\bm{\mathsf{B}}$, is the $I\times J$ matrix of entries
 $(\bm{\mathsf{A}}*\bm{\mathsf{B}})(i,j)=\bm{\mathsf{A}}(i,j)\bm{\mathsf{B}}(i,j).$
-The Hadamard product can be visualized as element-wise multiplication,
-$\underset{\bm{\mathsf{A}}}{\begin{bmatrix}a_{1,1}&\cdots&a_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}&\cdots&a_{I,J}\end{bmatrix}}*\underset{\bm{\mathsf{B}}}{\begin{bmatrix}b_{1,1}&\cdots&b_{1,J}\\ \vdots&\ddots&\vdots\\ b_{I,1}&\cdots&b_{I,J}\end{bmatrix}}=\underset{\bm{\mathsf{A}}*\bm{\mathsf{B}}}{\begin{bmatrix}a_{1,1}b_{1,1}&\cdots&a_{1,J}b_{1,J}\\ \vdots&\ddots&\vdots\\ a_{I,1}b_{I,1}&\cdots&a_{I,J}b_{I,J}\end{bmatrix}}.$
 The tensor-times-matrix (TTM) product is the tensor resulting from the mode-$j$ product of a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{j}\times\cdots\times N_{d}}$ with a matrix $\bm{\mathsf{U}}\in\mathbb{F}^{K\times N_{j}}$, denoted by $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{j}\bm{\mathsf{U}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{j-1}\times K\times N_{j+1}\times\cdots\times N_{d}}$, with entries
@@
 The order of multiplication does not matter, cf. (Ballard_Kolda_2025, Proposition 3.19). Equivalently, after mode-$j$ matricization,
-$\bm{\mathsf{Y}}_{(j)}=\bm{\mathsf{U}}_{j}\bm{\mathsf{X}}_{(j)}\left(\bm{\mathsf{U}}_{d}\otimes\cdots\otimes\bm{\mathsf{U}}_{j+1}\otimes\bm{\mathsf{U}}_{j-1}\otimes\cdots\otimes\bm{\mathsf{U}}_{1}\right)^{\top}.$
+$\bm{\mathsf{Y}}_{(j)}=\bm{\mathsf{U}}_{j}\bm{\mathsf{X}}_{(j)}\left(\bm{\mathsf{U}}_{d}\otimes\cdots\otimes\bm{\mathsf{U}}_{j+1}\otimes\bm{\mathsf{U}}_{j-1}\otimes\cdots\otimes\bm{\mathsf{U}}_{1}\right)^{*}.$
 5.3 Tensor Decompositions
 This work considers two fundamental formats of tensor decompositions: the canonical polyadic (CP) decomposition and the Tucker decomposition.
-We introduce them in Sect.
-5.3 and 5.3, along with prototypical algorithms to compute them.
-The CP decomposition has been known by many different names since its introduction in 1927: the polyadic form Hitchcock1927TheEO, PARAFAC (parallel factors) Harshman1970, CANDECOMP (canonical decomposition) carroll1970analysis; see also (KoldaBader2009, Table 3.1).
-It is often viewed as a higher-order extension of the matrix SVD KoldaBader2009.
+The CP decomposition has been known by many different names since its introduction in 1927 in Hitchcock1927TheEO; Hitchcock1927Multiple, cf.
+(KoldaBader2009, Table 3.1).
+This decomposition is often viewed as a higher-order extension of the matrix SVD, as it is formed from rank-one components.
 The Tucker decomposition Tuck1963a; tucker1964; Tucker1966 is a representation of a tensor as a Multi-TTM product of a core tensor with a factor matrix for each mode.
-Like the CP decomposition, the Tucker decomposition has been known by several different names in the literature (KoldaBader2009, Table 4.1), and the Tucker decomposition can also be viewed as an extension of the matrix SVD, in the form of higher-order principal component analysis (PCA) kapteyn1986approach; kroonenberg1980principal; LeiboviciSabatier1998; Vasilescu2002.
-We briefly summarize a few key properties of these decompositions in Sect. 5.3 before discussing low-rank tensor approximations that can be expressed in each format in Sect. 5.3.
+Like the CP decomposition, the Tucker decomposition has been known by several different names in the literature, cf.
+(KoldaBader2009, Table 4.1), and the Tucker decomposition can also be viewed as an extension of the matrix SVD, in the form of higher-order principal component analysis (PCA).
 The Canonical Polyadic (CP) Decomposition
@@
 We can express (37) in terms of mode-$j$ unfoldings using the Khatri-Rao product:
-$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}.$
+$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}.$
 The notation of Kolda2006 is frequently used as shorthand for the CP decomposition:
@@
 (39)
-In practice, the CP decomposition is often computed by the alternating least squares (ALS) method carroll1970analysis; Harshman1970.
-Pseudocode for the general ALS procedure is given in Algorithm 5, which we quickly summarize.
-Given a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the ALS procedure outputs a CP decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$ with $R$ rank-one components, once some convergence criteria is met.
-This is accomplished through a sequence of linear least squares problems, where all but one factor matrix is fixed:
-$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{X}}_{(j)}-\widehat{\bm{\mathsf{A}}}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right\|^{2}_{\text{F}}.$ (46)
+In practice, the CP decomposition is often computed by the alternating least squares (ALS) method carroll1970analysis; Harshman1970 in Algorithm 5.
+Given a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, CP-ALS outputs $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$ with $R$ rank-one components through a sequence of linear least squares problems, where all but one factor matrix is fixed:
+$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{X}}_{(j)}-\widehat{\bm{\mathsf{A}}}\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}\right\|^{2}_{\text{F}}.$ (40)
 The solution of (40) is given by
-$\bm{\mathsf{A}}^{(j)}:=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right]^{\dagger},$ (47)
+$\bm{\mathsf{A}}^{(j)}:=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}\right]^{\dagger},$ (41)
 where the columns of $\bm{\mathsf{A}}^{(j)}$ are normalized to determine $\lambda_{r}$ for $r=1,\ldots,R$.
@@
 If Algorithm 5 is naively implemented, the per-iteration cost to compute a CP decomposition is dominated by $N=\textup{size}(\bm{\mathcal{X}})$, which is computationally undesirable.
-Fortunately, since its introduction in 1970, CP-ALS has benefited from many modifications for improved efficiency, cf.
-Sect.
-6.1 for recent algorithmic developments.
-One strategy that has garnered attention for efficient CP-ALS incorporates a preprocessing step to compress the input tensor before executing Algorithm 5; this method is known as CANDELINC BroAndersson1998PartII; Carroll1980.
-The compression is often achieved through a Tucker decomposition, which we discuss next.
+Fortunately, since its introduction in 1970, CP-ALS has benefited from many modifications for improved efficiency.
+One strategy that has garnered attention incorporates a preprocessing step to compress the input tensor before executing Algorithm 5, known as CANDELINC BroAndersson1998PartII; Carroll1980, which may be achieved through a Tucker decomposition.
 Algorithm 5 Alternating Least Squares for CP Decomposition (CP-ALS)
@@
 5:  for $j=1,\ldots,d$ do
-6:   $\bm{\mathsf{A}}^{(j)}=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{\top}\right]^{\dagger}$
+6:   $\bm{\mathsf{A}}^{(j)}=\bm{\mathsf{X}}_{(j)}\left[\left(\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\right)^{*}\right]^{\dagger}$
 7:   $\lambda_{r}=\left\|\bm{\mathsf{A}}^{(j)}(:,r)\right\|_{2}$, $r=1,\ldots,R$
@@
 Tucker Decomposition
-The Tucker decomposition expresses a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$ as a $d$-way Multi-TTM product:
+The Tucker decomposition expresses a tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$ as a $d$-way Multi-TTM
 $\bm{\mathcal{X}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\times_{2}\bm{\mathsf{A}}^{(2)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}=:\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\bm{\mathsf{A}}^{(2)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket,$ (42)
@@
 We can express (42) in terms of mode-$j$ unfoldings by
-$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\bm{\mathsf{G}}_{(j)}\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)^{\top}.$ (49)
+$\bm{\mathsf{X}}_{(j)}=\bm{\mathsf{A}}^{(j)}\bm{\mathsf{G}}_{(j)}\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)^{*}.$ (43)
 A Tucker decomposition of a 3-mode tensor is visualized in Fig. 2.
@@
 For an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots N_{d}}$, each factor matrix $\bm{\mathsf{A}}^{(j)}$ is computed as the leading-$R_{j}$ left singular vectors of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$.
-If $R_{j}<\textup{rank}(\bm{\mathsf{X}}_{(j)})$, the decomposition is said to be “truncated.” When $R_{j}\ll N_{j}$, the Tucker decomposition is said to be a compressed representation of the input tensor.
-Once the factor matrices are computed, the core tensor is obtained as a Multi-TTM product of the original tensor with the factor matrix transposes.
-The computational cost of the HOSVD is $O(dn^{d+1}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity.
-For improved efficiency, the sequentially truncated HOSVD (ST-HOSVD) is often used Hackbusch2019; Vannieuwenhoven2012.
-In the ST-HOSVD, once a factor matrix is computed for mode $j$, the input tensor is compressed via mode-$j$ TTM before the next factor matrix is computed; i.e., for input tensor $\bm{\mathcal{X}}$, we first compute $\bm{\mathsf{A}}^{(1)}$ and form $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{A}}^{(1)}$, then compute $\bm{\mathsf{A}}^{(2)}$ from $\bm{\mathcal{Y}}$, then form $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{2}\bm{\mathsf{A}}^{(2)}$, and so on.
+If $R_{j}<\textup{rank}(\bm{\mathsf{X}}_{(j)})$, the decomposition is said to be “truncated.” When $R_{j}\ll N_{j}$, we say the tensor is compressed.
+The core tensor is a Multi-TTM of the original tensor with the factor matrix transposes.
+The total cost is $O(dn^{d+1}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity.
+For improved efficiency, the sequentially truncated HOSVD (ST-HOSVD) is often used Vannieuwenhoven2012.
+In the ST-HOSVD, once the mode-$j$ factor matrix is computed, the tensor is compressed via mode-$j$ TTM before the next factor matrix is computed; i.e., we first compute $\bm{\mathsf{A}}^{(1)}$ and form $\bm{\mathcal{Y}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{A}}^{(1)}$, then compute $\bm{\mathsf{A}}^{(2)}$ from $\bm{\mathcal{Y}}$, then form $\bm{\mathcal{Z}}=\bm{\mathcal{Y}}\times_{2}\bm{\mathsf{A}}^{(2)}$, and so on.
 ST-HOSVD successively reduces the size of the tensor used in subsequent computations, and the core tensor is available immediately after computing the factor matrix for the last mode.
-The computational cost of ST-HOSVD is $O(\sum_{j=1}^{d}k^{j-1}n^{d-j+2}+k^{j}n^{d-j})$.
+The total cost is $O(\sum_{j=1}^{d}k^{j-1}n^{d-j+2}+k^{j}n^{d-j})$.
 Algorithm 6 Higher-order SVD for Tucker Decomposition (HOSVD)
@@
 (Ballard_Kolda_2025, Chapter 7).
-However, we can use the HOSVD output as a starting point for an ALS-style algorithm to compute Tucker decompositions, known as higher-order orthogonal iteration (HOOI) kroonenberg1980principal; DeLathauwer2000a.
+However, we can use the HOSVD output as a starting point for an ALS-style algorithm to compute Tucker decompositions, known as higher-order orthogonal iteration (HOOI) or Tucker-ALS kroonenberg1980principal; DeLathauwer2000a.
 Pseudocode for HOOI is given in Algorithm 7.
-As in CP-ALS, in HOOI, we seek to minimize the approximation error $\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|$, now over $\bm{\mathcal{G}}$ and $\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}$, subject to the constraints that $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1},\ldots,R_{d}}$ and that $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$ is orthogonal for $j=1,\ldots,d$.
-This tensorial minimization problem can be transformed into matrix least squares problems in each mode $j$, where we solve for each factor matrix via
-$\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\widehat{\bm{\mathsf{A}}}^{\top}-\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}.$
-Note the optimal solution to this least squares problem is given by the leading left singular vectors of each unfolding $\bm{\mathsf{X}}_{(j)}$ Andersson1998; DeLathauwer2000; Kolda2006.
+We seek to minimize the approximation error $\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|$, now over $\bm{\mathcal{G}}$ and $\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}$, subject to the constraints that $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1},\ldots,R_{d}}$ and that $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$ is orthogonal for $j=1,\ldots,d$.
+This minimization problem can be transformed into a matrix least squares problem in each mode $j$, so that
+$\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)}\right)\bm{\mathsf{G}}^{*}_{(j)}\widehat{\bm{\mathsf{A}}}^{*}-\bm{\mathsf{X}}^{*}_{(j)}\right\|^{2}_{F}.$
+The solution to this least squares problem is given by the leading left singular vectors of each unfolding $\bm{\mathsf{X}}_{(j)}$ DeLathauwer2000; Kolda2006.
 The core tensor can be computed as a column vector $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{x}}\right\|^{2}_{2}$, where $\bm{\mathsf{x}}$ is the vectorization or unfolding of tensor $\bm{\mathcal{X}}$ into a large column vector of size $N\times 1$.
 To get $\bm{\mathcal{G}}$, we “re-fold” $\bm{\mathsf{g}}\in\mathbb{F}^{\prod_{j}R_{j}}$ back into a tensor of dimension $\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$.
-If the factor matrices are computed as the leading $R_{j}$ left singular vectors of unfoldings, then the optimal core tensor is given by $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\left(\bm{\mathsf{A}}^{(1)}\right)^{\top}\times\cdots\times_{d}\left(\bm{\mathsf{A}}^{(d)}\right)^{\top}$.
+If the factor matrices are computed as the leading $R_{j}$ left singular vectors of unfoldings, then the optimal core tensor is given by $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\left(\bm{\mathsf{A}}^{(1)}\right)^{*}\times\cdots\times_{d}\left(\bm{\mathsf{A}}^{(d)}\right)^{*}$.
 The per-iteration cost of Algorithm 7 is $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$, letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$.
@@
 5:  for $j=1,\ldots,d$ do
-6:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}$
+6:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{*}_{(j)}\bm{\mathsf{A}}^{*}-\bm{\mathsf{X}}^{*}_{(j)}\right\|^{2}_{F}$
 7:  end for
@@
 $\textup{rank}(\bm{\mathcal{X}})=\min\{R\in\mathbb{N}\ |\ \bm{\mathcal{X}}=\llbracket\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket,\ \bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R}\ \forall j\in[d]\}.$ (44)
-Mode ordering and TTM products with nonsingular matrices do not affect tensor rank. However, it is NP-hard to determine Hrastad1990.
+Mode ordering and TTM products with nonsingular matrices do not affect tensor rank.
+However, it is NP-hard to determine Hrastad1990.
 In practice, the target ranks for low-rank tensor decompositions are chosen heuristically.
@@
 (Ballard_Kolda_2025, Section 9.4.1).
-Generally, the optimization procedure in CP-ALS is repeated until the error no longer decreases or a maximum number of iterations is reached.
+Generally, the optimization procedure in CP-ALS is repeated until the error no longer decreases significantly.
 For Tucker decompositions, it is useful to consider the $j$-rank of a tensor, denoted by $\textup{rank}_{j}(\bm{\mathcal{X}})$ and defined as the column rank of its mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$ kruskal1989rank; DeLathauwer2000a.
@@
 In practice, we often prescribe a core $j$-rank for $j=1,\ldots,d$ that meets some error tolerance or compression ratio (Ballard_Kolda_2025, Section 4.2).
-The (ST-)HOSVD algorithm has the nice property that it computes a rank-$(R_{1},\ldots,R_{d})$ decomposition with an approximation error that is within $\sqrt{d}$ of the optimal rank-$(R_{1},\ldots,R_{d})$ approximation error.
 Whether it is better to represent an input tensor in the CP or Tucker format is generally problem-dependent.
 Under mild assumptions, the CP decomposition of a tensor is unique, up to scaling and permutation of its rank-one components tenberge2002uniqueness; harshman1984data; Harshman1970; kruskal1977three; kruskal1989rank; sidiropoulos2000uniqueness.
-The same cannot be said for Tucker decompositions; the mode-$j$ factor matrix can be multiplied by a nonsingular matrix and the core tensor updated via mode-$j$ product of the original core tensor with the inverse of the nonsingular matrix.
-As a result, the factors of a Tucker decomposition are less interpretable in the context of the original data than those of a CP decomposition, cf.
-(Ballard_Kolda_2025, Section 9.6–9.8).
-However, as previously mentioned, the Tucker decomposition is highly advantageous for data compression, even as a preprocessing step for a CP decomposition, e.g., CANDELINC BroAndersson1998PartII; Carroll1980.
+The same cannot be said for Tucker decompositions; the mode-$j$ factor matrix can be multiplied by a nonsingular matrix and the core tensor updated via mode-$j$ TTM with its inverse.
+As a result, the factors of a Tucker decomposition are less interpretable in the context of the original data than those of a CP decomposition.
 Despite its utility in data compression, the Tucker decomposition still suffers from the curse of dimensionality: letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the cost of storing a Tucker decomposition is $O(k^{d}+dkn)$.
-By contrast, the storage cost of a CP decomposition is linear in its rank, but its rank may be very large to achieve the desired approximation accuracy.
+By contrast, the storage cost of a CP decomposition is linear in its rank, but its rank may be very large to achieve the desired accuracy.
 In short, the notion of rank for tensor decompositions is more nuanced than for matrices, and the choice of representation and target rank (or $j$-rank) should be informed by the desired trade-off between approximation accuracy and computational performance.
@@
 In the previous sections, we introduced two fundamental tensor representations, the CP and the Tucker decomposition, and we summarized their key properties.
-In this section, we survey low-rank tensor factorizations that can be expressed in each format, which we characterize primarily by whether or not the factor matrices are orthogonal.
-Namely, we assume that the standard Tucker decomposition has orthonormal factor matrices, whereas CP decompositions generally do not.
-Higher-order versions of ID/CUR may be expressed using orthogonal factor matrices or factor matrices that contain entries of the input tensor, so we further characterize them by whether the core subtensor or the factor matrices contain entries of the original tensor.
+In this section, we survey low-rank tensor factorizations that can be expressed in each format.
 Orthogonal factor matrices
@@
 As in the matrix setting (Sect.
-2.4), a tensor ID or CUR factorization comprises actual entries of the input tensor.
+2.5), a tensor ID or CUR factorization comprises actual entries of the input tensor.
 The advantages of ID/CUR factorizations for tensors are similar to those for matrices.
@@
 The “structure-preserving ST-HOSVD” of MinsterSaibabaKilmer2020, initially developed for sparse rank-$(R_{1},\ldots,R_{d})$ tensors, may also be interpreted as a tensor CUR factorization of the first category.
-The “Chidori” and “Fiber” tensor CUR factorizations with rank-$(R_{1},\ldots,R_{d})$ core tensors were characterized in Cai2021, inspired by those in Caiafa2010; these factorizations were utilized for robust tensorial principal component analysis in Cai2024.
+The “Chidori” and “Fiber” tensor CUR factorizations with rank-$(R_{1},\ldots,R_{d})$ core tensors were characterized in Cai2021, inspired by those in Caiafa2010.
 The recently proposed “CoreID” in Zhang2025 is similar to the CUR factorizations of Cai2021.
@@
 The early algorithm of Drineas2007tensorSVD to compute a $\textup{rank-}(R_{1},\ldots,R_{d})$ “approximate tensor SVD” forms the factor matrices in a Tucker decomposition as randomly sampled columns of unfoldings; this factorization would later be termed a “higher-order ID” in Saibaba2016, where the mode-$j$ factor matrix in a Tucker decomposition is computed as column skeletons of the mode-$j$ unfolding selected by greedy pivoting.
-In FriedlandMehrmannMiedlarNkengla2011, an HOOI procedure is developed for 3-mode Tucker tensors that yields factor matrices of original tensor entries, and the “CURT” algorithm of ZhaoSong2018RETL; Song2019 computes tensor CUR decompositions in the Tucker format with strong accuracy guarantees.
-The CURT algorithm was recently generalized in Zhang2025 to compute rank-$(R_{1},\ldots,R_{d})$ Tucker decompositions where the core tensor is unconstrained and the factor matrices are column skeletons of mode unfoldings; this algorithm is also suitable for input tensors expressed in the CP format.
-The tensor ID algorithm of Biagioni2015 computes a CP decomposition of a $d$-mode tensor, in which factor matrices are comprised of entries of the input tensor, as does the more recent work of MalikBecker2020.
-We detail several algorithms that leverage randomization to compute tensor ID/CUR factorizations efficiently in Sect.
+We detail several recent algorithms that leverage randomization to compute tensor ID/CUR approximations efficiently in Sect.
 6.
 6 Randomized Low-Rank Tensor Decompositions
-As in the matrix setting, we are interested in algorithms that compute low-rank tensor decompositions accurately and efficiently, with provable performance guarantees that are substantiated by empirical evidence.
+As in the matrix setting, we are interested in algorithms that compute low-rank tensor decompositions accurately and efficiently, with provable performance guarantees, and with demonstrated high practical speed.
 Because many tensor algorithms fundamentally rely on classical matrix algorithms, we can frequently fall back on the ideas introduced in Sect.
@@
 However, given the rapidly increasing number of applications that involve large-scale tensorial data, it is often impractical or infeasible to rely on basic implementations of the randomized rangefinder, e.g., with Gaussian sketching.
-The curse of dimensionality for $d$-way arrays is even more pernicious than for matrices, so great care must be taken to develop randomized algorithms that are not only accurate, but also scalable.
 In this section, we survey new developments in fast randomized algorithms for low-rank tensor decompositions in the CP and Tucker formats.
-We draw connections and refer back to our presented matrix algorithms when relevant, augmenting them with emerging ideas from the recent proliferation of work on randomized algorithms for tensors.
 Sect.
-6.1 explores how CP-ALS can be accelerated with randomization, specifically through fast sampling methods that exploit the structure of the overdetermined least squares problems solved in each iteration.
+6.1 explores how CP-ALS can be accelerated with randomization, specifically through fast sampling methods that exploit the structure of the overdetermined least squares problems.
 Sect.
-6.2 summarizes randomized methods to compute Tucker decompositions having orthogonal factor matrices, which naturally extend the randomized rangefinder to the tensor setting via the standard randomized (ST-)HOSVD and randomized HOOI, as well as a newly proposed algorithm for more efficient sketching.
-These randomized algorithms are contrasted with those for tensor ID/CUR decompositions in Sect.
+6.2 summarizes randomized methods to compute Tucker decompositions with orthogonal factors, which naturally extend the randomized rangefinder to the tensor setting.
+These algorithms are contrasted with those for tensor ID/CUR decompositions in Sect.
 6.3, which includes both the CP and Tucker formats.
-Our intent throughout this section is not to provide an exhaustive summary of all recent developments, but rather to broadly survey them to direct the interested reader to the relevant literature, and then hone in on key ideas that may stimulate future work in these emergent areas.
+Our goal is to broadly survey relevant literature and detail recent developments that may stimulate future work.
 6.1 Randomized Algorithms for CP Decomposition
@@
 (41), for fast sketching and sampling.
-In Wang2015, the TensorSketch algorithm of Pham2013 is used to sketch the input tensor without explicitly forming it, then the KRP is efficiently computed.
+In Wang2015, the input tensor is sketched without being explicitly formed, then the KRP is computed.
 The “randomized ALS” algorithm of Reynolds2016 incorporates projections onto random tensors to improve the conditioning of the least squares problems in CP-ALS.
-A procedure to sample rows of the KRP near-optimally via fast leverage score approximations is presented in Cheng2016; however, the least squares problem is not sketched in this work (only the KRP).
-In BattaglinoBallardKolda2018, a Kronecker fast Johnson-Lindenstrauss transform (KFJLT) is applied to the KRP to reduce the coherence for better sampling performance, cf.
-Jin2021; MALIK2020120, which is then utilized in a sketched ALS problem without explicitly forming the KRP.
+A procedure to sample rows of the KRP near-optimally via fast leverage score approximations is presented in Cheng2016.
+In BattaglinoBallardKolda2018, a Kronecker fast Johnson-Lindenstrauss transform (KFJLT, cf.
+Jin2021) is applied to the KRP to reduce the coherence for better sampling performance.
 The algorithm of Aggour2020 builds on BattaglinoBallardKolda2018 with options for adaptive sketch sizes and regularization terms.
-We now expand on the recent work of Larsen2022 on approximate leverage score sampling that exploits the structure of the KRP.
+Below, we expand on the recent work of Larsen2022 on improved leverage score sampling that exploits KRP structure.
 Randomized CP-ALS with Fast Leverage Score Sampling
@@
 More precisely, given input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, in each iteration of CP-ALS, we solve a least squares problem in each mode $j=1,\ldots,d$, in which every factor matrix except $\bm{\mathsf{A}}^{(j)}$ is fixed, and $\bm{\mathsf{A}}^{(j)}$ is computed as the solution of
-$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\widehat{\bm{\mathsf{A}}}\bm{\mathsf{Z}}^{\top}-\bm{\mathsf{X}}_{(j)}\right\|^{2}_{\text{F}}=\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{Z}}\widehat{\bm{\mathsf{A}}}^{\top}-\bm{\mathsf{X}}_{(j)}^{\top}\right\|^{2}_{\text{F}},$ (51)
+$\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\widehat{\bm{\mathsf{A}}}\bm{\mathsf{Z}}^{*}-\bm{\mathsf{X}}_{(j)}\right\|^{2}_{\text{F}}=\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\left\|\bm{\mathsf{Z}}\widehat{\bm{\mathsf{A}}}^{*}-\bm{\mathsf{X}}_{(j)}^{*}\right\|^{2}_{\text{F}},$ (45)
 where $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(d)}\odot\cdots\odot\bm{\mathsf{A}}^{(j+1)}\odot\bm{\mathsf{A}}^{(j-1)}\odot\cdots\odot\bm{\mathsf{A}}^{(1)}\in\mathbb{F}^{N^{(-j)}\times R}$.
 If we solve (45) without exploiting any inherent problem structure, the cost of computing the QR decomposition of $\bm{\mathsf{Z}}$ is $O(R^{2}N^{(-j)})$, and the cost of applying it $N_{j}$ times brings the total cost to $O(R^{2}N^{(-j)}+RN)$.
-In Larsen2022, a randomized sampling strategy based on fast leverage score approximation, CP-ARLS-LEV, is used to accelerate each least squares sub-problem (51).
+In Larsen2022, the CP-ARLS-LEV algorithm is developed to accelerate each least squares sub-problem (45).
 Namely, $S\ll N^{(-j)}$ rows of $\bm{\mathsf{Z}}$ are sampled according to an approximate leverage score distribution.
-The corresponding $S$ rows of $\bm{\mathsf{X}}_{(j)}^{\top}$ are then selected for a smaller “sketched” problem whose solution is a good approximation of the solution of (51) with high probability.
+The corresponding $S$ rows of $\bm{\mathsf{X}}_{(j)}^{*}$ are then selected for a smaller “sketched” problem, whose solution approximates the solution of (45).
 Sketching reduces the cost of solving the $j$th least squares problem to $O(R^{2}S+RSN_{j})$.
 However, forming $\bm{\mathsf{Z}}$ or $\bm{\mathsf{X}}$ and computing leverage scores would incur a cost of $O(R^{2}N^{(-j)})$.
-To avoid this, in CP-ARLS-LEV, the leverage scores are estimated as the product of leverage scores of the respective factor matrices in the KRP, and the KRP structure is exploited so that only the $S$ selected rows of $\bm{\mathsf{Z}}$ are formed explicitly.
+To avoid this, in CP-ARLS-LEV, the leverage scores are estimated as the product of leverage scores of the respective factor matrices in the KRP, and only the $S$ sampled rows of $\bm{\mathsf{Z}}$ are formed explicitly.
 To this end, let $j\in[d]$ be fixed.
-We first observe that there is a bijection mapping each row $i\in[N]$ of $\bm{\mathsf{Z}}$ to a (d1)(d-1)-tuple $(i_{d},\ldots,i_{j+1},i_{j-1},\ldots,i_{1})$, with $i_{k}\in[N_{k}]$, $k\neq j,$ which indexes rows of factor matrices in the KRP:
-$\bm{\mathsf{Z}}(i,:)=\bm{\mathsf{A}}^{(d)}(i_{d},:)*\cdots*\bm{\mathsf{A}}^{(j+1)}(i_{j+1},:)*\bm{\mathsf{A}}^{(j-1)}(i_{j-1},:)*\cdots*\bm{\mathsf{A}}^{(1)}(i_{1},:),$ (52)
-where * again denotes the Hadamard product.
+We first observe that there is a bijection mapping each row $i\in[N^{(-j)}]$ of $\bm{\mathsf{Z}}$ to a (d1)(d-1)-tuple $(i_{d},\ldots,i_{j+1},i_{j-1},\ldots,i_{1})$, with $i_{k}\in[N_{k}]$, $k\neq j,$ which indexes rows of the factor matrices in the KRP:
+$\bm{\mathsf{Z}}(i,:)=\bm{\mathsf{A}}^{(d)}(i_{d},:)*\cdots*\bm{\mathsf{A}}^{(j+1)}(i_{j+1},:)*\bm{\mathsf{A}}^{(j-1)}(i_{j-1},:)*\cdots*\bm{\mathsf{A}}^{(1)}(i_{1},:).$ (46)
 Using (Cheng2016, Theorem 3.3), the leverage scores of the rows of $\bm{\mathsf{Z}}$ are bounded above by the product of the leverage scores of the rows of its constituent factor matrices:
 $\ell_{i}(\bm{\mathsf{Z}})\leq\hat{\ell_{i}}(\bm{\mathsf{Z}}):=\prod_{k\neq j}\ell_{i_{k}}\left(\bm{\mathsf{A}}^{(k)}\right).$ (47)
-From (53), we define a probability distribution on the rows of $\bm{\mathsf{Z}}$; namely, row $i$ is selected with probability
+From (47), we define a probability distribution on the rows of $\bm{\mathsf{Z}}$:
 $p_{i}=\hat{\ell_{i}}(\bm{\mathsf{Z}})/R^{d-1}.$ (48)
-Computing the probabilities in (54) explicitly would require the Kronecker product of leverage scores, costing $O(N^{(-j)})$ operations.
-However, (Larsen2022, Lemma 9) establishes that each $i_{k}$ in the (d1)(d-1)-tuple corresponding to row $i$ of $\bm{\mathsf{Z}}$ can be sampled independently, using the leverage scores of the rows of $\bm{\mathsf{A}}^{(k)}$, $k\neq j$, which cost $O(N_{k}R^{2})$ to compute.
+Computing the probabilities in (48) explicitly would cost $O(N^{(-j)})$ operations.
+However, (Larsen2022, Lemma 9) establishes that each $i_{k}$ in the (d1)(d-1)-tuple can be sampled independently, using the leverage scores of the rows of $\bm{\mathsf{A}}^{(k)}$, $k\neq j$, which cost $O(N_{k}R^{2})$ to compute.
 This gives a multi-index corresponding to the $i$-th row of $\bm{\mathsf{Z}}$ without computing the Kronecker product of leverage scores, and $\bm{\mathsf{Z}}(i,:)$ can be formed in O(R(d1))O(R(d-1)) operations using (46).
-There are additional modifications that can be made to the CP-ARLS-LEV algorithm to further improve performance; we refer the reader to Larsen2022 for details.
-The basic pseudocode for this procedure is given in Algorithm 8.
-Algorithm 8 Alternating Randomized Least Squares for CP with Leverage Scores (CP-ARLS-LEV)
-1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, number of rank-one components $R$, number of samples $S$
-2:CP decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\lambda};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
-3:Initialize $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R}$, $j=1,\ldots,d$ $\triangleright$ e.g., $\bm{\mathsf{A}}^{(j)}=\texttt{randn}(N_{j},R)$
-4:for $j=1,\ldots,d$ do
-5:  Initialize $\bm{\mathsf{p}}_{j}=\bm{\mathsf{\ell}}(\bm{\mathsf{A}}^{(j)})/R$, the vector of row leverage scores for $\bm{\mathsf{A}}^{(j)}$
-6:end for
-7:while not converged do
-8:  for $j=1,\ldots,d$ do
-9:   $I=\emptyset$
-10:   while $|I|<S$ do
-11:     for $k=1,\ldots,j-1,j+1,\ldots,d$ do
-12:      Sample $i_{k}\sim\bm{\mathsf{p}}_{k}$ from $N_{k}]$
-13:     end for
-14:     Add $i=(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$ to II
-15:   end while
-16:   Form submatrix $\widehat{\bm{\mathsf{Z}}}=\bm{\mathsf{Z}}(I,:)$
-17:   Form submatrix $\widehat{\bm{\mathsf{X}}}_{(j)}^{*}=\bm{\mathsf{X}}_{(j)}^{*}(I,:)$
-18:   Compute $\bm{\mathsf{A}}^{(j)}=\arg\min_{\widehat{\bm{\mathsf{A}}}\in\mathbb{F}^{N_{j}\times R}}\|\widehat{\bm{\mathsf{Z}}}\widehat{\bm{\mathsf{A}}}^{*}-\widehat{\bm{\mathsf{X}}}_{(j)}^{*}\|^{2}_{\text{F}}$
-19:   Update $\bm{\mathsf{p}}_{j}=\bm{\mathsf{\ell}}(\bm{\mathsf{A}}^{(j)})/R$
-20:  end for
-21:end while
-The CP-ARLS-LEV algorithm enjoys strong theoretical guarantees.
-Namely, the resulting rank-$R$ CP decomposition is an $\varepsilon$-accurate approximation to the input tensor with high probability when the number of samples $S=O\left(R^{d-1}/\varepsilon\right)$.
+There are additional modifications that can be made to CP-ARLS-LEV to further improve performance; see Larsen2022 for details.
+The CP-ARLS-LEV algorithm enjoys nice theoretical guarantees.
+Namely, the resulting rank-$R$ CP decomposition is an $\varepsilon$-accurate approximation to the input tensor when the number of samples $S=O\left(R^{d-1}/\varepsilon\right)$, under certain assumptions on the structure of the randomized DRM (pertaining to its injectivity parameter and the sketched residual norm); cf.
+(Larsen2022, Appendix A.1).
 In particular, CP-ARLS-LEV improves on the sampling complexity required by the KFJLT-based method of Jin2021.
 A distributed-memory implementation of CP-ARLS-LEV is also developed in Bharadwaj2024, along with a distributed-memory version of an algorithm that performs random walks on a binary tree for fast leverage score estimation instead.
-The recent work of Malik2022 improves on the sampling complexity of CP-ARLS-LEV by avoiding exponential dependence on $d$, using the fast leverage score estimation technique of DrineasMagdonMahoneyWoodruff12, but replacing the FJLT with recursive sketching Ahle2020.
-This strategy yields a sampling distribution closer to the exact leverage score distribution, but performs similarly to CP-ARLS-LEV in practice, and its implementation details are beyond the scope of the current manuscript; we refer the reader to Ahle2020; Malik2022.
+The recent work of Malik2022 improves on the sampling complexity of CP-ARLS-LEV by avoiding exponential dependence on $d$, using the fast leverage score estimation technique of DrineasMagdonMahoneyWoodruff12, but using recursive sketching Ahle2020.
+This strategy yields a sampling distribution closer to the exact leverage score distribution, but performs similarly to CP-ARLS-LEV in practice, and its implementation details are beyond the scope of the current manuscript.
 There has also been success in using stochastic gradient descent on the least squares objective function in (45), cf.
 KoldaHong2020; Wang2023; Yu2024.
-Randomized algorithms to compute low-rank Tucker decompositions as a pre-processing step for CP-ALS have also been investigated Erichson_2020; Yang2018; Zhou2014.
-We discuss several of these algorithms in the next section.
-6.2 Randomized Algorithms for Tucker Decompositions with Orthogonal Factors
-Because the Tucker decomposition provides a natural framework for data compression, methods based on the HOSVD (Algorithm 6) and HOOI (Algorithm 7) have been utilized in many data science and scientific computing applications.
-However, the Tucker decomposition suffers from the curse of dimensionality, thus classical algorithms can be prohibitively expensive for large problems.
-As randomized algorithms, particularly the randomized SVD (Algorithm 3), have become mainstays for fast and reliable matrix compression, it is natural to consider how randomization can be leveraged for fast and reliable tensor compression in the Tucker format.
-In this section, we survey recent work on randomized algorithms for Tucker decompositions with orthogonal factor matrices (as opposed to interpolating, cf.
-Sect.
-5.3).
-We begin with an overview of major developments in randomized Tucker decompositions, before detailing two algorithms based on the randomized SVD: the randomized HOSVD and the randomized HOOI.
+Randomized algorithms to compute low-rank Tucker decompositions as a pre-processing step for CP-ALS have also been investigated; we discuss several of these algorithms next.
+6.2 Randomized Algorithms for Tucker Decompositions
+The Tucker decomposition provides a natural framework for data compression.
+As randomized algorithms, particularly the randomized SVD (Algorithm 3), have become mainstays for fast and reliable matrix compression, it is natural to consider how randomization can be leveraged for fast and reliable tensor compression.
 Randomized algorithms for Tucker decompositions have been investigated in many problem settings, generally involving either randomized sampling or sketching.
-One of the earliest works on randomized HOSVD and HOOI algorithms is Tsourakakis2010, called the MACH-HOSVD and MACH-HOOI, so named for the work of Achlioptas-McSherry Achlioptas2001 that inspired these randomized sampling algorithms.
+One of the earliest works on randomized HOSVD and HOOI algorithms is Tsourakakis2010, called the MACH-HOSVD and MACH-HOOI.
 In MACH-HOSVD and MACH-HOOI, the input tensor is sparsified according to a coin flip for each nonzero entry; with probability $p$, a nonzero entry is kept and reweighted by $1/p$, and with probability 1p1-p, the entry is set to 0.
-The HOSVD and HOOI algorithms are then performed on the sparsified tensor, outputting a Tucker decomposition that approximates the original tensor well in practice, even for small values of $p$, e.g., $p=0.1$.
 Leverage score sampling of matrix unfoldings to sparsify the tensor is explored in Perros2015, and Tucker decompositions of sub-sampled tensors are also investigated in Oseledets2008; Vervliet2014BreakingTC.
@@
 Formal analysis of RandTucker is performed in Erichson_2020, which incorporates power iteration into RandTucker for faster spectral decay; it is noted, though, that uniform sampling instead of Gaussian sketching performs similarly and may be preferable for large problems.
-Power iteration and randomized sketching are utilized in the related work of CheWeiYan2020; CheWeiYan2021 for randomized HOSVD and HOOI, which build on the rank-adaptive algorithm of CheWei2019 that iteratively computes an orthonormal basis for the range of each mode unfolding with $\varepsilon$-accuracy.
-Rank-adaptive versions of the randomized (ST-)HOSVD are also presented in MinsterSaibabaKilmer2020, which apply the adaptive randomized rangefinder to each mode unfolding.
+Power iteration and randomized sketching are utilized in the related work of CheWeiYan2020; CheWeiYan2021 for randomized HOSVD and HOOI, which build on the rank-adaptive algorithm of CheWei2019 that computes a basis for the range of each mode unfolding with $\varepsilon$-accuracy.
 In sun2020low, a single-pass algorithm based on the HOSVD is developed that significantly improves on the cost of storing the random sketching matrices, via Khatri-Rao products of random dimension reduction maps (called tensor random projections); details can also be found in TroppStreaming2019.
@@
 Randomized (ST-)HOSVD
-In the HOSVD (Algorithm 6), we compute factor matrices for a Tucker decomposition as the left singular vectors of each mode unfolding.
-The randomized HOSVD is a natural extension, where we substitute the SVD of each unfolding with the randomized SVD (Algorithm 3).
-Pseudocode is given in Algorithm 9, this time for the ST-HOSVD to highlight the slight difference in its computation vs.
-the HOSVD; the R-HOSVD algorithm of MinsterSaibabaKilmer2020 is precisely the HOSVD algorithm, but using the randomized SVD.
-Note that this algorithm can be made adaptive by using the adaptive randomized rangefinder with a prescribed error tolerance for the low-rank decomposition of each unfolding.
-Algorithm 9 Randomized Sequentially Truncated Higher-Order SVD (R-STHOSVD)
-1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, target ranks $R_{1},\ldots,R_{d}$, oversampling parameter $p$
-2:Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
-3:Initialize $\bm{\mathcal{G}}=\bm{\mathcal{X}}$
-4:for $j=1,\ldots,d$ do
-5:  Draw Gaussian matrix $\bm{\mathsf{\Omega}}^{(j)}\in\mathbb{R}^{N^{(-j)}\times(R_{j}+p)}$
-6:  Form sketch $\bm{\mathsf{Y}}^{(j)}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}^{(j)}$
-7:  $\widehat{\bm{\mathsf{U}}},\widehat{\bm{\mathsf{\Sigma}}},\widehat{\bm{\mathsf{V}}}]=\texttt{svd}(\bm{\mathsf{Y}}^{(j)},R_{j})$
-8:  Set $\bm{\mathsf{A}}^{(j)}=\widehat{\bm{\mathsf{U}}}$
-9:  Update $\bm{\mathsf{G}}_{(j)}=\widehat{\bm{\mathsf{\Sigma}}}\widehat{\bm{\mathsf{V}}}^{\top}$
-10:end for
-If $\widehat{\bm{\mathcal{X}}}$ is the output of Algorithm 9 (or the randomized HOSVD, with Gaussian sketching), it is proven in (MinsterSaibabaKilmer2020, Theorem 3.1, 3.2) that the expected error is
-$\mathbb{E}_{\{\bm{\mathsf{\Omega}}^{(j)}\}_{j=1}^{d}}\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}\|\leq\left(d+\frac{\sum_{j=1}^{d}R_{j}}{p-1}\right)^{1/2}\|\bm{\mathcal{X}}-\widehat{\bm{\mathcal{X}}}_{opt}\|.$ (55)
-Letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots=R_{d}$ for notational simplicity, the computational cost of R-STHOSVD as in Algorithm 9 is $O(\sum_{j=1}^{d}k^{j}n^{d-j+1}+k^{j}n^{d-j})$.
-By contrast, R-HOSVD (substituting the SVD with the randomized SVD in Algorithm 6) has computational cost $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$.
-In other words, the cost is reduced roughly by a factor of $n$ when using randomized (ST-)HOSVD vs.
-deterministic.
-Algorithmic performance can also be improved in the R-STHOSVD by processing the largest modes first, and while the error of a given output does depend on processing order, the worst-case expected error does not, cf.
+The randomized HOSVD is a natural extension of the HOSVD, where we substitute the SVD of each unfolding with the randomized SVD (Algorithm 3).
+The R-HOSVD algorithm of MinsterSaibabaKilmer2020 is precisely Algorithm 6, but with a sketch instead of the unfolding.
+If $\widehat{\bm{\mathcal{X}}}$ is the output of the randomized HOSVD, with Gaussian sketching of each mode-$j$ unfolding, the expected error is within $\approx\sqrt{d(k+1)}$ of the optimal rank-$(k,\ldots,k)$ Tucker approximation error, cf.
+(MinsterSaibabaKilmer2020, Theorem 3.1).
+Letting $n$ be the uniform mode size again, the computational cost of R-STHOSVD is $O(\sum_{j=1}^{d}k^{j}n^{d-j+1}+k^{j}n^{d-j})$, where the first term is the cost of computing SVDs of mode unfoldings and the second term is the cost of forming the core tensor.
+By contrast, the R-HOSVD algorithm costs $O(dkn^{d}+\sum_{j=1}^{d}k^{j}n^{d-j+1})$ operations.
+In other words, the cost is reduced roughly by a factor of $n$ with Gaussian sketching.
+Performance can be improved by processing the largest modes first; while the resulting error does depend on processing order, the worst-case expected error does not, cf.
 (MinsterSaibabaKilmer2020, Sect.
@@
 Randomized HOOI
-Recall from Section 5.3 that higher-order orthogonal iteration (HOOI) is a method of computing Tucker decompositions by iteratively solving least-squares problems for each mode’s factor matrix, which is typically more accurate than the HOSVD but more computationally expensive.
-Similar to our goal in CP-ALS, for a given input $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, we are seeking a rank-$(R_{1},\ldots,R_{d})$ Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}$ that minimizes
+Recall from Section 5.3 that HOOI computes Tucker decompositions iteratively by solving least-squares problems in each mode for the factor matrix, generally more accurate than the HOSVD but more expensive.
+Similar to our goal in CP-ALS, for a given input $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, we are seeking $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}$ that minimizes
 $\min_{\bm{\mathcal{G}},\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}}\|\bm{\mathcal{X}}-\bm{\mathcal{G}}\times_{1}\bm{\mathsf{A}}^{(1)}\cdots\times_{d}\bm{\mathsf{A}}^{(d)}\|,$ (49)
 where $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$ and $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$.
-Through an appropriate transformation of the objective function, we can re-write the HOOI procedure in Algorithm 7 as two main steps that are repeated until convergence; cf.
+We can re-write the HOOI procedure in Algorithm 7 as two main steps that are repeated until convergence; cf.
 Malik2018:
   1. 1.
-    For $j=1,\ldots,d$, solve $\bm{\mathsf{A}}^{(j)}=\underset{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}{\arg\min}\|\bm{\mathsf{Z}}^{(j)}\bm{\mathsf{G}}_{(j)}^{\top}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{X}}_{(j)}^{\top}\|^{2}$, where
+    For $j=1,\ldots,d$, solve $\bm{\mathsf{A}}^{(j)}=\underset{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}{\arg\min}\|\bm{\mathsf{Z}}^{(j)}\bm{\mathsf{G}}_{(j)}^{*}\bm{\mathsf{A}}^{*}-\bm{\mathsf{X}}_{(j)}^{*}\|^{2}$, where
     $\bm{\mathsf{Z}}^{(j)}=\bm{\mathsf{A}}^{(d)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(j+1)}\otimes\bm{\mathsf{A}}^{(j-1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(1)},$
@@
     Update $\bm{\mathcal{G}}=\underset{\bm{\mathcal{Y}}\in\mathbb{F}^{R_{1}\times R_{d}}}{\arg\min}\|\bm{\mathsf{Z}}\bm{\mathsf{y}}-\bm{\mathsf{x}}\|^{2}$, where $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(d)}\otimes\cdots\bm{\mathsf{A}}^{(1)}$, and $\bm{\mathsf{y}},\bm{\mathsf{x}}$ are (column) vectorizations of $\bm{\mathcal{Y}}$, $\bm{\mathcal{X}}$, respectively.
-We observe that each of these two steps involve large overdetermined least squares problems, and similar to the strategy in CP-ARLS-LEV, we will exploit the matrix product structures to solve smaller sketched least squares problems efficiently, now with TensorSketch, cf.
-(Malik2018, Algorithm 2 (TUCKER-TS)).
-TensorSketch   We briefly summarize TensorSketch pagh2013compressed; Pham2013; AvronNguyenWoodruff2014; DiaoSongSunWoodruff2018, a specialized version of the CountSketch operator of Charikar2004.
-A CountSketch operator can be defined as a linear map $\bm{\mathsf{S}}:\mathbb{R}^{I}\rightarrow\mathbb{R}^{J}$ given by $\bm{\mathsf{S}}=\bm{\mathsf{P}}\bm{\mathsf{D}}$, where $\bm{\mathsf{P}}\in\mathbb{R}^{J\times I}$ is a matrix with $\bm{\mathsf{P}}(h(i),i)=1$ and all other entries equal to 0, for a random map $h:[I]\rightarrow[J]$ such that $\mathbb{P}[h(i)=j]=1/J$ for all $i\in[I]$, $j\in[J]$.
-The matrix $\bm{\mathsf{D}}\in\mathbb{R}^{I\times I}$ is a diagonal matrix with each entry equal to $\pm 1$ with equal probability.
-In particular, a CountSketch operator $\bm{\mathsf{S}}$ can be rapidly applied to a matrix $\bm{\mathsf{A}}$ in $O(\textup{nnz}(\bm{\mathsf{A}}))$ operations, without explicitly forming $\bm{\mathsf{S}}$ or $\bm{\mathsf{A}}$.
-More specifically, if $\bm{\mathsf{S}}\in\mathbb{R}^{J\times I}$ and $\bm{\mathsf{A}}\in\mathbb{F}^{I\times L}$, typically with $J\ll I$, then we can form $\bm{\mathsf{Y}}=\bm{\mathsf{S}}\bm{\mathsf{A}}$ by hashing each row $i$ of $\bm{\mathsf{A}}$ with an integer $h$ sampled uniformly from $[J]$, assigning it a value $s$ of $\pm 1$ with equal probability, then summing the rows with the same hash value (i.e.
+We observe that each of these two steps involve large overdetermined least squares problems, and similar to the strategy in CP-ARLS-LEV, we will exploit the matrix product structures to solve smaller sketched least squares problems efficiently, now with TensorSketch.
+TensorSketch   TensorSketch pagh2013compressed; Pham2013; AvronNguyenWoodruff2014; DiaoSongSunWoodruff2018 is a specialized version of CountSketch.
+A CountSketch operator can be defined as a linear map $\bm{\mathsf{S}}:\mathbb{R}^{I}\rightarrow\mathbb{R}^{J}$ given by $\bm{\mathsf{S}}=\bm{\mathsf{P}}\bm{\mathsf{D}}$, where $\bm{\mathsf{P}}\in\mathbb{R}^{J\times I}$ is a matrix with $\bm{\mathsf{P}}(h(i),i)=1$ and all other entries zero, for a random map $h:[I]\rightarrow[J]$ such that $\mathbb{P}[h(i)=j]=1/J$ for all $i\in[I]$, $j\in[J]$.
+The matrix $\bm{\mathsf{D}}\in\mathbb{R}^{I\times I}$ is a diagonal matrix of $\pm 1$ with equal probability.
+The CountSketch operator $\bm{\mathsf{S}}$ can be rapidly applied to a matrix $\bm{\mathsf{A}}$ in $O(\textup{nnz}(\bm{\mathsf{A}}))$ operations, without explicitly forming $\bm{\mathsf{S}}$ or $\bm{\mathsf{A}}$.
+If $\bm{\mathsf{S}}\in\mathbb{R}^{J\times I}$ and $\bm{\mathsf{A}}\in\mathbb{F}^{I\times L}$, typically $J\ll I$, we form $\bm{\mathsf{Y}}=\bm{\mathsf{S}}\bm{\mathsf{A}}$ by hashing each row $i$ of $\bm{\mathsf{A}}$ with an integer $h$ sampled uniformly from $[J]$, assigning it a value $s$ of $\pm 1$ with equal probability, then summing the rows with the same hash value (i.e.
 adding $s\bm{\mathsf{A}}_{i,:}$ to the $h$th row of $\bm{\mathsf{Y}}$ if $\bm{\mathsf{A}}_{i,:}$ has hash value $h$).
 A TensorSketch operator is a specific type of CountSketch operator, which can be rapidly applied to Kronecker products, e.g., $\bm{\mathsf{Z}}=\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)}\in\mathbb{F}^{n^{d}\times k^{d}}$.
-(We again assume here that $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$.) If $\bm{\mathsf{T}}:\mathbb{R}^{n^{d}}\rightarrow\mathbb{R}^{\ell}$ is a TensorSketch operator with embedding dimension $\ell\ll n^{d}$, then the cost of forming $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is only $O(\ell dnk+\ell k^{d})$, suppressing log factors, instead of $O(\ell n^{d}k^{d})$ for naive multiplication.
+If $\bm{\mathsf{T}}:\mathbb{R}^{n^{d}}\rightarrow\mathbb{R}^{\ell}$ is a TensorSketch operator, then the cost of sketching $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is $O(\ell dnk+\ell k^{d})$, suppressing log factors, instead of $O(\ell n^{d}k^{d})$ for naive multiplication.
 This is accomplished by sketching each factor matrix with independent CountSketch operators, then convolving with Fast Fourier Transforms (FFTs).
-To that end, let $h_{j}:[n]\rightarrow[\ell]$ for $j=1,\ldots,d$ be 3-wise independent hash functions, i.e.
-for any distinct $i_{1},i_{2},i_{3}\in[n]$, the hash codes $h_{j}(i_{1})$, $h_{j}(i_{2})$, $h_{j}(i_{3})$ are independent random variables, and $h_{j}(i)$ is uniformly distributed in $\ell$] for any fixed $i\in[n]$.
-Let $s_{j}:[n]\rightarrow\{+1,-1\}$ be 4-wise independent sign functions (so that analogously, $s_{j}(i_{1})$, $s_{j}(i_{2})$, $s_{j}(i_{3})$, $s_{j}(i_{4})$ are independent random variables with $s_{j}(i)\sim\textup{Uniform}\{\pm 1\}$.
+Let $h_{j}:[n]\rightarrow[\ell]$ for $j=1,\ldots,d$ be 3-wise independent hash functions, i.e.
+for any distinct $i_{1},i_{2},i_{3}\in[n]$, the hash codes $h_{j}(i_{1})$, $h_{j}(i_{2})$, $h_{j}(i_{3})$ are $\textup{Uniform}\{[\ell]\}$.
+Let $s_{j}:[n]\rightarrow\{+1,-1\}$ be 4-wise independent sign functions (i.e.
+$\textup{Uniform}\{\pm 1\}$).
 We form independent CountSketch matrices $\bm{\mathsf{S}}^{(j)}=\bm{\mathsf{P}}^{(j)}\bm{\mathsf{D}}^{(j)}$ for $j=1,\ldots,d$, using the hash functions $h_{j}$ to form $\bm{\mathsf{P}}^{(j)}$ and $s_{j}(i)=\bm{\mathsf{D}}^{(j)}_{i,i}$ for $i\in[n]$.
-The application of $\bm{\mathsf{S}}^{(j)}$ to each column $\bm{\mathsf{A}}^{(j)}_{:,r_{j}}$, for $r_{j}=1,\ldots,k$, can be represented by an $(\ell-1)$-degree polynomial
+Applying $\bm{\mathsf{S}}^{(j)}$ to each column $\bm{\mathsf{A}}^{(j)}_{:,r_{j}}$, for $r_{j}=1,\ldots,k$, can be represented by an $(\ell-1)$-degree polynomial
 $\mathcal{P}^{(j)}_{r_{j}}(\omega)=\sum_{i=1}^{n}s_{j}(i)\bm{\mathsf{A}}^{(j)}_{i,r_{j}}\omega^{h_{j}(i)-1}=:\sum_{l=1}^{\ell}c^{(j)}_{l,r_{j}}\omega^{j-1},$ (50)
-and we define $\bm{\mathsf{c}}_{r_{j}}^{(j)}$ to be the vector of coefficients $(c^{(j)}_{1,r_{j}},\ldots,c^{(j)}_{\ell,r_{j}})$, which are grouped in (57) by hash value.
-Let $\bm{\mathsf{T}}$ be the TensorSketch operator, which is defined as the CountSketch operator formed from hash function $H$ and sign function $S$ given by
+where $\bm{\mathsf{c}}_{r_{j}}^{(j)}$ are coefficients $(c^{(j)}_{1,r_{j}},\ldots,c^{(j)}_{\ell,r_{j}})$, grouped in (50) by hash value.
+The TensorSketch operator $\bm{\mathsf{T}}$ is defined as the CountSketch operator comprised of hash function $H$ and sign function $S$ given by
 $H:[n]^{d}\rightarrow\ell,\ \ (i_{1},\ldots,i_{d})\mapsto\left(\sum_{j=1}^{d}(h_{j}(i_{j})-1)\mod\ell\right)+1,$ (51)
 $S:[n]^{d}\rightarrow\{+1,-1\},\ \ (i_{1},\ldots,i_{d})\mapsto\prod_{j=1}^{d}s_{j}(i_{j}).$ (52)
-Then the $r$th column of $\bm{\mathsf{T}}(\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)})=\bm{\mathsf{T}}\bm{\mathsf{Z}}$ for $r=1,\ldots,k^{d}$ can also be represented as a polynomial
+The $r$-th column of $\bm{\mathsf{T}}(\bm{\mathsf{A}}^{(1)}\otimes\cdots\otimes\bm{\mathsf{A}}^{(d)})=\bm{\mathsf{T}}\bm{\mathsf{Z}}$, $r=1,\ldots,k^{d}$, is the polynomial
 $\mathcal{P}_{r}(\omega)$ $=\sum_{i=1}^{n^{d}}S(i_{1},\ldots,i_{d})\bm{\mathsf{Z}}_{i,r}\omega^{H(i_{1},\ldots,i_{d})},$ (53)
@@
 $=\textup{FFT}^{-1}\left(\textup{FFT}(\bm{\mathsf{c}}^{(1)}_{r_{1}})*\cdots*\textup{FFT}(\bm{\mathsf{c}}_{r_{d}}^{(d)}\right),$ (55)
-where * denotes the Hadamard product, and $i$ in (60) corresponds to the tuple $(i_{1},\ldots,i_{d})\in[n]^{d}$, and similarly $r$ corresponds to the tuple $(r_{1},\ldots,r_{d})\in[k]^{d}$.
-In other words,
-$\bm{\mathsf{T}}\bm{\mathsf{Z}}=\textup{FFT}^{-1}\left(\left({\bigodot}_{j=1}^{d}\left(\textup{FFT}\left(\bm{\mathsf{S}}^{(j)}\bm{\mathsf{A}}^{(j)}\right)\right)^{\top}\right)^{\top}\right),$ (63)
+where * denotes the Hadamard product, $i$ in (53) corresponds to the tuple $(i_{1},\ldots,i_{d})\in[n]^{d}$, and $r$ corresponds to the tuple $(r_{1},\ldots,r_{d})\in[k]^{d}$. In other words,
+$\bm{\mathsf{T}}\bm{\mathsf{Z}}=\textup{FFT}^{-1}\left(\left({\bigodot}_{j=1}^{d}\left(\textup{FFT}\left(\bm{\mathsf{S}}^{(j)}\bm{\mathsf{A}}^{(j)}\right)\right)^{*}\right)^{*}\right),$ (56)
 where $\odot$ is the KRP, and the FFT is applied column-wise.
 Instead of the usual $O(\ell n^{d}k^{d})$ cost of matrix multiplication, the cost of $\bm{\mathsf{T}}\bm{\mathsf{Z}}$ is $O(\ell dnk+\ell k^{d}\log(\ell))$ using (56).
-Using TensorSketch in HOOI We now turn our attention back to the randomized HOOI algorithm of Malik2018, which uses TensorSketch to reduce the size of each least squares subproblem in the two primary steps of HOOI listed above.
-The least squares problem to find the factor matrix for each mode $j=1,\ldots,d$ is highly overdetermined, involving a Kronecker product $\bm{\mathsf{Z}}^{(j)}$ of size $N^{(-j)}\times(\prod_{k\neq j}R_{k})$.
-The least squares problem to find the core tensor is also highly overdetermined, involving a Kronecker product of size $\prod_{j=1}^{d}N_{j}\times\prod_{j=1}^{d}R_{j}$, or $n^{d}\times k^{d}$ for convenience.
-To compute the factor matrices and core tensor, we form independent TensorSketches to solve smaller sketched least squares problems corresponding to both of these steps.
-Pseudocode for R-HOOI, equivalently (Malik2018, TUCKER-TS), is given in Algorithm 10.
-This algorithm can be implemented as a single-pass algorithm through the input tensor, cf.
+Using TensorSketch in HOOI We now turn our attention back to the randomized HOOI algorithm of Malik2018, called TUCKER-TS, which uses TensorSketch in each least squares subproblem in the two main steps above.
+In Step 1, the factor matrix sub-problems for modes $j=1,\ldots,d$ are highly overdetermined, involving an $n^{d-1}\times k^{d-1}$ Kronecker product $\bm{\mathsf{Z}}^{(j)}$.
+The core tensor sub-problem in Step 2 is also highly overdetermined, involving a $n^{d}\times k^{d}$ Kronecker product.
+For both steps, we form independent TensorSketches to solve smaller sketched least squares problems, as in Sect.
+3.3.
+TUCKER-TS can also be made single-pass; cf.
 (Malik2018, Algorithm S3).
-The dominant cost of Algorithm 10 (single-pass) is $O(dnk^{2}+\ell dk^{d}+\ell_{c}k^{d}+k^{2d})$, or roughly $k^{O(d)}$, versus the dominant cost of $O(n^{d})$ for deterministic HOOI (Algorithm 7), if we suppress $\log$ factors and assume sketching dimension $\ell=\ell_{1}=\cdots=\ell_{d}$.
-We refer the reader to (Malik2018, S3.2.6) for details on the complexity analysis.
-However, we observe that the single-pass R-HOOI algorithm as presented in Malik2018 uses the same TensorSketches for sketching the updated Kronecker products until some convergence criteria is met.
-In practice, the performance of R-HOOI as written in Algorithm 10 is curiously better in terms of error reduction than if new independent TensorSketch operators were drawn in every iteration, though it is unclear why this is the case.
-However, for the theoretical guarantees of DiaoSongSunWoodruff2018 to hold for TensorSketched least squares, it is technically necessary to form new independent TensorSketch operators in each iteration of R-HOOI.
-In Malik2018, the numerical results suggest that sketching dimensions $\ell=Ck^{d-1}$ and $\ell_{c}=Cr^{d}$ for constant $C>4$ (e.g., $C=10$) work well in practice.
-Algorithm 10 Randomized Higher-order Orthogonal Iteration with TensorSketch (R-HOOI)
-1:Input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, factor matrix target ranks $R_{1},\ldots,R_{d}$, sketching dimensions $\ell_{1},\ldots,\ell_{d}$, $\ell_{c}$
-2:Tucker decomposition $\widehat{\bm{\mathcal{X}}}=\llbracket\bm{\mathcal{G}};\bm{\mathsf{A}}^{(1)},\ldots,\bm{\mathsf{A}}^{(d)}\rrbracket$
-3:Initialize $\bm{\mathsf{A}}^{(j)}\in\mathbb{F}^{N_{j}\times R_{j}}$, $j=2,\ldots,d$
-4:Draw TensorSketch matrices $\bm{\mathsf{T}}^{(j)}\in\mathbb{R}^{\ell_{j}\times(\prod_{i\neq j}N_{i})}$ for $i,j\in[d]$ and $\bm{\mathsf{T}}^{(c)}\in\mathbb{R}^{\ell_{c}\times\prod_{j}N_{j}}$
-5:while not converged do
-6:  for $j=1,\ldots,d$ do
-7:   $\bm{\mathsf{A}}^{(j)}=\arg\min_{\bm{\mathsf{A}}\in\mathbb{F}^{N_{j}\times R_{j}}}\left\|\left(\bm{\mathsf{T}}^{(j)}\bigotimes_{i=d,i\neq j}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{G}}^{\top}_{(j)}\bm{\mathsf{A}}^{\top}-\bm{\mathsf{T}}^{(j)}\bm{\mathsf{X}}^{\top}_{(j)}\right\|^{2}_{F}$
-8:  end for
-9:  $\bm{\mathsf{g}}=\arg\min_{\bm{\mathsf{z}}\in\mathbb{F}^{\prod_{j}R_{j}\times 1}}\left\|\left(\bm{\mathsf{T}}^{(c)}\bigotimes_{j=d}^{1}\bm{\mathsf{A}}^{(j)}\right)\bm{\mathsf{z}}-\bm{\mathsf{T}}^{(c)}\bm{\mathsf{x}}\right\|^{2}_{2}$ $\triangleright$ Column vectorization $\bm{\mathsf{x}}$ of $\bm{\mathcal{X}}$
-10:  Re-fold vectorization $\bm{\mathsf{g}}$ into $\bm{\mathcal{G}}\in\mathbb{F}^{R_{1}\times\cdots\times R_{d}}$
-11:end while
+The dominant cost of randomized HOOI with TensorSketch (TUCKER-TS) is $O(dnk^{2}+\ell dk^{d}+\ell_{c}k^{d}+k^{2d})$, or roughly $k^{O(d)}$, versus the dominant cost of $O(n^{d})$ for classical HOOI (Algorithm 7), suppressing $\log$ factors and letting $\ell=\ell_{1}=\cdots=\ell_{d}$.
+See (Malik2018, S3.2.6) for details on the complexity analysis.
+We note that the single-pass algorithm presented in Malik2018 uses the same TensorSketch operators on the updated Kronecker products until some convergence criteria is met.
+For the theoretical guarantees of DiaoSongSunWoodruff2018 to hold, it is technically necessary to form new independent TensorSketch operators in each iteration; however, in practice, it is noted in Malik2018 that re-using TensorSketch operators is more efficient and works just as well.
+Their numerical results suggest $\ell=Ck^{d-1}$ and $\ell_{c}=Cr^{d}$ for constant $C>4$ (e.g., $C=10$) are usually appropriate.
 Randomized TSMS
-We end this section by briefly highlighting the recent work of Hashemi2025 on Randomized Tucker decompositions with Single-Mode Sketching (RTSMS).
+We end this section by highlighting the recent work of Hashemi2025 on Randomized Tucker decompositions with Single-Mode Sketching (RTSMS).
 The RTSMS algorithm sketches the input tensor one mode at a time, i.e.
 sketching on the left of the mode-$j$ unfolding vs.
-sketching on the right, to avoid the computational bottleneck of multiplying along the larger dimension $N^{(-j)}=\prod_{i\neq j}N_{i}$.
-The algorithm is also single-pass, so that the input tensor is not revisited after the first sketch, and rank-adaptive if the target ranks for the factor matrices are unknown a priori.
-To illustrate, suppose we want to compute a rank-$(R_{1},R_{2},R_{3})$ HOSVD that approximates an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$.
-(We omit rank-adaptivity and iterative refinement; the reader is referred to Hashemi2025 for practical implementation details.) We initialize tensors
-$\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{X}}\ \textup{and}\ \bm{\mathcal{G}}^{\textup{new}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}^{(1)}\iff\bm{\mathsf{G}}^{\textup{new}}_{(1)}=\left(\bm{\mathsf{\Omega}}^{(1)}\right)^{\top}\bm{\mathsf{X}}_{(1)}\in\mathbb{F}^{\tilde{R_{1}}\times(N_{2}N_{3})}$
+sketching on the right, to avoid the computational bottleneck of multiplying along the larger dimension $N^{(-j)}\equiv n^{d-1}$.
+The algorithm is also single-pass and rank-adaptive.
+To illustrate, suppose we want to compute a rank-$(R_{1},R_{2},R_{3})$ HOSVD approximating $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times N_{2}\times N_{3}}$.
+(We omit rank-adaptivity and iterative refinement; see Hashemi2025 for practical implementation details.) We initialize tensors
+$\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{X}}\ \textup{and}\ \bm{\mathcal{G}}^{\textup{new}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}^{(1)}\iff\bm{\mathsf{G}}^{\textup{new}}_{(1)}=\left(\bm{\mathsf{\Omega}}^{(1)}\right)^{*}\bm{\mathsf{X}}_{(1)}\in\mathbb{F}^{\tilde{R_{1}}\times(N_{2}N_{3})}$
 for some sketching dimension $\tilde{R_{1}}>R_{1}$.
@@
 (Hashemi2025, Section 4.2).
-We then update $\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{G}}^{\textup{new}}$.
-For the next mode, we first update $\bm{\mathcal{G}}^{\textup{new}}$ via mode-2 sketching:
-$\bm{\mathsf{G}}^{\textup{new}}_{(2)}=\left(\bm{\mathsf{\Omega}}^{(2)}\right)^{\top}\bm{\mathsf{G}}_{(2)}^{\textup{old}}\in\mathbb{F}^{\tilde{R_{2}}\times(\tilde{R_{1}}N_{3})}$
+We update $\bm{\mathcal{G}}^{\textup{old}}=\bm{\mathcal{G}}^{\textup{new}}$, and for the next mode, update $\bm{\mathcal{G}}^{\textup{new}}$ via
+$\bm{\mathsf{G}}^{\textup{new}}_{(2)}=\left(\bm{\mathsf{\Omega}}^{(2)}\right)^{*}\bm{\mathsf{G}}_{(2)}^{\textup{old}}\in\mathbb{F}^{\tilde{R_{2}}\times(\tilde{R_{1}}N_{3})}$
 for some sketching dimension $\tilde{R_{2}}>R_{2}$.
@@
 The computational complexity of RTSMS is dominated by the first TTM $\bm{\mathcal{X}}\times_{1}\bm{\mathsf{\Omega}}_{1}$.
-Letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$, the dominant cost is $O(kn^{d})$ to compute the first TTM, but the sketch size is only $k\times n$ vs.
+The dominant cost is $O(kn^{d})$ to compute the first TTM, but the sketch size is only $k\times n$ vs.
 $k\times n^{d-1}$ for R-HOSVD.
 The least squares solves are also significantly reduced by sub-sampling and by computing a generalized Nyström decomposition vs.
-SVD for the factor matrices.
-The full tensor does not need to be sketched again in modes $j=2,\ldots,d$, and the core tensor is available immediately upon termination, as in the ST-HOSVD algorithm.
+SVD.
+The full tensor does not need to be sketched again in later modes, and the core tensor is available immediately upon termination.
 Additional details can be found in Hashemi2025.
@@
 However, there exists a natural categorization of tensor ID/CUR decompositions, based on whether the factor matrices or core tensors in the computed decompositions are comprised of entries of the input tensor.
-In this section, we explore how randomization can be leveraged in prototypical algorithms to compute a Tucker ID/CUR decomposition from each category.
-We also discuss recent work on randomized algorithms for tensor ID/CUR in the CP format, which can be viewed as a form of input tensor pruning.
+We discuss randomized algorithms to compute a Tucker ID/CUR decomposition from each category, as well as recent work on CP tensor ID/CUR.
 Randomized Algorithms for Tucker Core ID/CUR
-We first consider randomized algorithms for tensor ID/CUR factorizations in the Tucker format of the first category, where the core tensor in the compressed representation is a sub-tensor of the input tensor.
-In particular, we focus on the tensor ID/CUR decompositions that were recently characterized in Cai2024; Cai2021, termed the “Chidori” and “Fiber” tensor CUR factorizations.
-This work was inspired by numerous investigations of randomized sampling algorithms for tensor CUR decompositions, including CaiafaCichocki2009; Caiafa2010; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR.
-In Cai2021, given an input tensor $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the rank-$(R_{1},\ldots,R_{d})$ tensor CUR factorization is defined as
+We first consider randomized algorithms for Tucker tensor ID/CUR factorizations of the first type, where the core tensor is a sub-tensor of the input tensor.
+In particular, we focus on the tensor ID/CUR decompositions that were recently characterized in Cai2021, termed the “Chidori” and “Fiber” tensor CUR factorizations.
+This work was inspired by numerous investigations of tensor CUR decompositions, including Caiafa2010; Drineas2007tensorSVD; MahoneyMaggioniDrineas2008_tensorCUR.
+In Cai2021, given $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the rank-$(R_{1},\ldots,R_{d})$ tensor CUR factorization is
 $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{1}(\bm{\mathsf{C}}_{1}\bm{\mathsf{U}}_{1}^{\dagger})\times_{2}(\bm{\mathsf{C}}_{2}\bm{\mathsf{U}}_{2}^{\dagger})\times\cdots\times_{d}(\bm{\mathsf{C}}_{d}\bm{\mathsf{U}}_{d}^{\dagger}),$ (57)
@@
 $\bm{\mathcal{G}}=\bm{\mathcal{X}}(I_{1},\ldots,I_{d}),$ (58)
-for index sets $I_{j}\subset[N_{j}]$, $j=1,\ldots,d$. The mode-$j$ factor matrix $\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger}$ is defined in terms of the mode-$j$ unfolding:
+for index sets $I_{j}\subset[N_{j}]$, $j=1,\ldots,d$. The mode-$j$ factor matrix $\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger}$ is defined by
 $\begin{split}\bm{\mathsf{C}}_{j}&=\bm{\mathsf{X}}_{(j)}(:,J_{j}),\ \ \textup{and}\\ \bm{\mathsf{U}}_{j}&=\bm{\mathsf{C}}_{j}(I_{j},:),\end{split}$ (59)
@@
 If $J_{j}=\otimes_{k\neq j}I_{k}$, then (57) is called a “Chidori” CUR decomposition.
-By (Cai2021, Theorem 3.3), (HammHuang2020, Corollary 5.2), if the indices $I_{j}$ and $J_{j}$ are each sampled uniformly with $|I_{j}|=O(R_{j}\log N_{j})$ and $|J_{j}|=O(R_{j}\log N^{(-j)})$, then the computed Tucker decomposition is a rank-$(R_{1},\ldots,R_{d})$ approximation with high probability.
-We present pseudocode for either the Chidori or Fiber tensor CUR decomposition in Algorithm 11.
-With uniform sampling, the computational complexity of computing either tensor CUR decomposition is dominated by the cost of the pseudoinverse of $\bm{\mathsf{U}}_{j}$, which is $O(k^{d+1}\log^{d}(n))$ for Chidori and $O(k^{2}\log^{2}(n))$ for Fiber, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$ as before.
-Recall that the HOSVD has complexity $O(kn^{d})$ by comparison.
-Algorithm 11 Randomized Tucker ID/CUR with interpolating core
-[…]
-2:chidori or fiber decomposition $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{j=1}^{d}(\bm{\mathsf{C}}_{j}\bm{\mathsf{U}}_{j}^{\dagger})$
-3:for $j=1,\ldots,d$ do
-4:  if chidori then
-5:   Sample indices $I_{j}$ from $N_{j}]$, $|I_{j}|=S_{j}$ without replacement according to $\bm{\mathsf{p}}_{j}$
-6:   Set $J_{j}=\otimes_{i\neq j}I_{i}$
-7:  end if
-8:  if fiber then
-9:   Sample indices $I_{j}$ from $N_{j}]$, $|I_{j}|=S_{j}$ without replacement according to $\bm{\mathsf{p}}_{j}$
-10:   Sample indices $J_{j}$ from $N^{(-j)}]$, $|J_{j}|=T_{j}$ without replacement according to $\bm{\mathsf{q}}_{j}$
-11:  end if
-12:  Form $\bm{\mathsf{C}}_{j}=\bm{\mathsf{X}}_{(j)}(:,J_{j})$
-13:  Form $\bm{\mathsf{U}}_{j}=\bm{\mathsf{C}}_{j}(I_{j},:)$
-14:end for
-[…]
-The recent work of Zhang2025 proposes a similar factorization to (64), called the “CoreID.” The CoreID algorithm is reminiscent of the sequentially-truncated HOSVD, specifically the “structure-preserving” ST-HOSVD developed for sparse input tensors in MinsterSaibabaKilmer2020.
-Namely, in the CoreID, as each mode’s factor matrix is computed, the input tensor is compressed along that mode before proceeding to the next one, so that each subsequent unfolding is reduced in size.
+By (Cai2021, Theorem 3.3), (HammHuang2020, Corollary 5.2), if the indices $I_{j}$ and $J_{j}$ are each sampled uniformly with $|I_{j}|=O(R_{j}\log N_{j})$ and $|J_{j}|=O(R_{j}\log N^{(-j)})$, then the computed Tucker decomposition is a rank-$(R_{1},\ldots,R_{d})$ approximation with high probability, under certain incoherence assumptions on the input tensor; cf.
+Dong2021TensorCompletion.
+With uniform sampling, the computational complexity of computing either CUR decomposition is dominated by the cost of $\bm{\mathsf{U}}_{j}^{\dagger}$, which is $O(k^{d+1}\log^{d}(n))$ for Chidori and $O(k^{2}\log^{2}(n))$ for Fiber, letting $n=N_{1}=\cdots=N_{d}$ and $k=R_{1}=\cdots R_{d}$ as before.
+The recent work of Zhang2025 proposes a similar factorization to (57), called the “CoreID.” The CoreID algorithm is reminiscent of the “structure-preserving” ST-HOSVD developed for sparse input tensors in MinsterSaibabaKilmer2020.
+Namely, in the CoreID, as each factor matrix is computed, the input tensor is compressed along that mode before proceeding to the next one, so that each subsequent unfolding is reduced in size.
 Through matricization, computing the CoreID is reduced to computing a randomized matrix row ID in each mode, cf.
 Algorithm 4.
-For $j=1,\ldots,d$, the mode-$j$ unfolding of the tensor is sketched by a random matrix $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}_{j}.$ Then a row ID of $\bm{\mathsf{Y}}_{j}$ is computed by some skeleton selection method, e.g., greedy pivoting or squared-norm sampling.
+For $j=1,\ldots,d$, the mode-$j$ unfolding of the tensor is sketched by a random matrix $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{X}}_{(j)}\bm{\mathsf{\Omega}}_{j}.$ Then a row ID of $\bm{\mathsf{Y}}_{j}$ is computed by some skeleton selection method.
 The selected skeletons $I_{j}$ form the core sub-tensor, and the mode-$j$ factor matrix for the CoreID is defined as the associated interpolation matrix.
-The tensor is then compressed along mode $j$ as in ST-HOSVD before proceeding to the next mode.
-Sketching is the dominating cost in this algorithm, which is why many of the numerical investigations in Zhang2025 consider structured input tensors for fast sketching, e.g., a sparse input tensor or a CP-factorized input tensor.
+The tensor is then compressed along mode $j$ before proceeding to the next.
+Sketching is the dominant cost, so many of the numerical experiments in Zhang2025 consider structured input tensors for fast sketching, e.g., a sparse input tensor or a CP-factorized input tensor.
 Randomized Algorithms for Tucker Factor ID/CUR
 We now turn our attention to randomized algorithms for tensor ID/CUR factorizations in the Tucker format in which the factor matrices contain entries of the input tensor.
-We illustrate such factorizations with the “higher-order ID” of Saibaba2016, which we contrast with the “SatID” of Zhang2025.
+We detail the “higher-order ID” of Saibaba2016, contrasted with the “SatID” of Zhang2025.
 For a given $\bm{\mathcal{X}}\in\mathbb{F}^{N_{1}\times\cdots\times N_{d}}$, the higher-order ID (HOID) of Saibaba2016 is given by
@@
 $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{1}\bm{\mathsf{C}}_{1}^{\dagger}\times_{2}\bm{\mathsf{C}}_{2}^{\dagger}\times\cdots\times_{d}\bm{\mathsf{C}}_{d}^{\dagger},$ (61)
-and each factor matrix $\bm{\mathsf{C}}_{j}\in\mathbb{F}^{N_{j}\times R_{j}}$ consists of column skeletons of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$.
+and each factor $\bm{\mathsf{C}}_{j}\in\mathbb{F}^{N_{j}\times R_{j}}$ consists of column skeletons of the mode-$j$ unfolding $\bm{\mathsf{X}}_{(j)}$.
 The core tensor in (61) is optimal for the factorization in (60), cf.
 DeLathauwer2000.
-For the theoretical guarantees of Saibaba2016, the strong rank-revealing QR algorithm of gu1996 is used for skeleton selection on a random sketch $\bm{\mathsf{Y}}_{j}=\bm{\mathsf{\Omega}}_{j}\bm{\mathsf{X}}_{(j)}$, where $\bm{\mathsf{\Omega}}_{j}$ is an $(R_{j}+p)\times N_{j}$ random matrix for some small oversampling parameter $p$.
-However, in practice, CPQR or LUPP in the randomized matrix ID (e.g., Algorithm 4, with input $\bm{\mathsf{X}}_{(j)}^{*}$) performs similarly and is less computationally expensive.
-Pseudocode is provided in Algorithm 12.
-Algorithm 12 Randomized Tucker ID/CUR with interpolating factors
-[…]
-2:Tensor ID $\widehat{\bm{\mathcal{X}}}=\bm{\mathcal{G}}\times_{j=1}^{d}\bm{\mathsf{C}}_{j}$
-3:for $j=1,\ldots,d$ do
-4:  Compute rank-$R_{j}$ randomized column ID $\bm{\mathsf{X}}_{(j)}\approx\bm{\mathsf{C}}_{j}\bm{\mathsf{W}}_{j}$ $\triangleright$ e.g., Algorithm 4 on $\bm{\mathsf{X}}_{(j)}^{*}$
-5:end for
-6:Compute core $\bm{\mathcal{G}}=\bm{\mathcal{X}}\times_{j=1}^{d}\bm{\mathsf{C}}_{j}^{\dagger}$
-For each mode $j$, the cost of sketching is $O(R_{j}N_{j}N^{(-j)})=O(R_{j}N)$.
-Skeleton selection as in Saibaba2016 costs $O(R_{j}^{2}N^{(-j)})$.
-The cost of computing the core tensor is $O\left(\sum_{j=1}^{d}\left[R_{j}^{2}N_{j}+\prod_{i\leq j}R_{i}\cdot N^{(-j)}\right]\right)$.
-Letting $k=R_{1}=\cdots=R_{d}$ and $n=N_{1}=\cdots=N_{d}$ again, the total cost is roughly $O(dkn^{d}+dk^{d-1}n^{d-1})$.
+For the theoretical guarantees of Saibaba2016, the strong rank-revealing QR algorithm of gu1996 is used for skeleton selection on the random sketch of $\bm{\mathsf{X}}_{(j)}$.
+In practice, LUPP in the randomized matrix ID (i.e., Algorithm 4, with input $\bm{\mathsf{X}}_{(j)}^{*}$) performs similarly and is less expensive.
+For each mode $j$, the cost of sketching is $O(R_{j}N_{j}N^{(-j)})\equiv O(kn^{d})$.
+Skeleton selection as in Saibaba2016 costs $O(R_{j}^{2}N^{(-j)})\equiv O(k^{2}n^{d-1})$.
+The total cost is roughly $O(dkn^{d}+dk^{d-1}n^{d-1})$ after computing the core tensor.
 A sequentially-truncated version of the HOID (ST-HOID), analogous to the ST-HOSVD, is given in (Saibaba2016, Algorithm 4).
-A very similar decomposition is developed in Zhang2025, termed the “SatID,” which can be more efficiently computed.
-In particular, for an $n^{d}$-sized tensor, a QR-based algorithm to compute skeletons for the mode-$j$ factor matrix as in Saibaba2016 requires $n^{d-1}$ column norms, which is not optimal for structured tensors and can even have larger storage costs than the input tensor.
-By contrast, the SatID algorithm is based on a randomized matrix ID computed by marginalized norm sampling, which can be viewed as an extension of the original tensor CUR construction of Drineas2007tensorSVD, which uses norm sampling.
-We briefly summarize the marginalized norm sampling procedure of Zhang2025.
-As in Sect.
-6.1, for a fixed mode $j$, we can uniquely identify a column index $i\in[N^{(-j)}]$ of $\bm{\mathsf{X}}_{(j)}$ with a (d1d-1)-tuple $(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$.
+A very similar decomposition is developed in Zhang2025, termed the “SatID.” For an $n^{d}$-sized tensor, a QR-based algorithm to compute skeletons for the mode-$j$ factor matrix as in Saibaba2016 requires $n^{d-1}$ column norms, which is not optimal for structured tensors and can even have larger storage costs than the input.
+By contrast, the SatID algorithm is based on a randomized matrix ID with marginalized norm sampling, which can be viewed as an extension of the original tensor CUR of Drineas2007tensorSVD that uses norm sampling.
+We briefly summarize the marginalized norm sampling procedure.
+Recall that we can uniquely identify a column index $i\in[N^{(-j)}]$ of $\bm{\mathsf{X}}_{(j)}$ with a (d1d-1)-tuple $(i_{1},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$.
 Similar to CP-ARLS-LEV in Larsen2022, the idea is to accelerate sampling of $i\in[N^{(-j)}]$, now based on column norm sampling of $\bm{\mathsf{X}}_{(j)}$.
@@
 3.2) requires the computation of column scores
-$d_{i}^{(S)}=\|\bm{\mathsf{X}}_{:,i}-\bm{\mathsf{Q}}_{S}\bm{\mathsf{Q}}_{S}^{\top}\bm{\mathsf{X}}_{:,i}\|^{2}=\min_{\bm{\mathsf{v}}}\|\bm{\mathsf{X}}_{:,S}\bm{\mathsf{v}}-\bm{\mathsf{X}}_{:,i}\|^{2},\ \ i\in[N^{(-j)}],$ (69)
+$d_{i}^{(S)}=\|\bm{\mathsf{X}}_{:,i}-\bm{\mathsf{Q}}_{S}\bm{\mathsf{Q}}_{S}^{*}\bm{\mathsf{X}}_{:,i}\|^{2}=\min_{\bm{\mathsf{v}}}\|\bm{\mathsf{X}}_{:,S}\bm{\mathsf{v}}-\bm{\mathsf{X}}_{:,i}\|^{2},\ \ i\in[N^{(-j)}],$ (62)
 where $\bm{\mathsf{Q}}_{S}$ is an orthonormal basis for the range of $\bm{\mathsf{X}}_{:,S}$.
-If $X=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{d})$ is a random variable representing the next column index, then using (69),
-$\mathbb{P}(X=i)\propto d_{i}^{(S)}.$
-The key observation is that
+If $X=(X_{1},\ldots,X_{j-1},X_{j+1},\ldots,X_{d})$ is a random variable representing the next column index, from (62), $\mathbb{P}(X=i)\propto d_{i}^{(S)}.$ The key observation is that
 $\mathbb{P}(X_{1}=i_{1})\propto\sum_{i_{2},\ldots,i_{d}}d_{i_{1},i_{2},\ldots,i_{d}}^{(S)},$
-so that we can sample $i_{1}$ from its marginal distribution, then sample $i_{2}$ given $i_{1}$, and continue for all of $i=(i_{1},i_{2},\ldots,i_{j-1},i_{j+1},\ldots,i_{d})$.
-This sum can be computed efficiently via the equation derived in (Zhang2025, Eq.
-18):
-$\mathbb{P}(X_{1}=i_{1})\propto\left\|\widetilde{\bm{\mathsf{X}}}_{:,i_{1}}\right\|^{2},$ (70)
 […]
-For structured inputs, such as sparse tensors, this procedure can be modified for better performance.
-The total complexity (including randomized sketching) is $O(\textup{nnz}(\bm{\mathcal{X}})+\ell n^{2}k^{2})$ for a rank-$(k,\ldots,k)$ SatID of an $n^{d}$-sized tensor with sketch size $\ell$; see (Zhang2025, Algorithm 7) for details.
+For structured inputs, this procedure can be modified for better performance.
+The total complexity (including randomized sketching) is $O(\textup{nnz}(\bm{\mathcal{X}})+\ell n^{2}k^{2})$ for a rank-$(k,\ldots,k)$ SatID of an $n^{d}$-sized tensor with sketch size $\ell$; see (Zhang2025, Algorithm 7).
 Randomized Algorithms for CP Factor ID/CUR
@@
 We note that the matrix $\bm{\mathsf{Y}}$ can be computed efficiently without ever explicitly forming $\bm{\mathsf{R}}$ or $\bm{\mathsf{X}}$, since each $\bm{\mathsf{Y}}(l,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}^{(j)}_{l},\bm{\mathsf{a}}^{(j)}_{r}\rangle$.
-However, we note that the elements of $\bm{\mathsf{R}}$ are not independent, and so theoretical guarantees on the randomized matrix decomposition computed with $\bm{\mathsf{Y}}$ are not given, though it is demonstrated in Biagioni2015 that it performs well in practice.
-Pseudocode is provided in Algorithm 13.
-Algorithm 13 Randomized CP ID/CUR with interpolating factors
-1:CP tensor $\bm{\mathcal{X}}=\sum_{r=1}^{R}\bm{\mathsf{a}}^{(1)}_{r}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{r}$, target rank $k$, oversampling $p$
-2:Rank-$k$ CP tensor $\widehat{\bm{\mathcal{X}}}=\sum_{i=1}^{k}\bm{\mathsf{a}}^{(1)}_{s_{i}}\circ\cdots\circ\bm{\mathsf{a}}^{(d)}_{s_{i}}$
-3:Draw $\ell=k+p$ random rank-one tensors $\bm{\mathsf{r}}_{i}^{(1)}\circ\cdots\circ\bm{\mathsf{r}}_{i}^{(d)}$
-4:Form $\bm{\mathsf{Y}}\in\mathbb{F}^{\ell\times R}$ via $\bm{\mathsf{Y}}(\ell,r)=\prod_{j=1}^{d}\langle\bm{\mathsf{r}}_{\ell}^{(j)},\bm{\mathsf{a}}_{r}^{(j)}\rangle$
-5:Compute rank-$k$ column ID of $\bm{\mathsf{Y}}$ for skeletons $\{s_{1},\ldots,s_{k}\}$ $\triangleright$ e.g., Algorithm 4 applied to $\bm{\mathsf{Y}}^{*}$
+However, we observe that the elements of $\bm{\mathsf{R}}$ are not independent, and so theoretical guarantees on the randomized matrix decomposition computed with $\bm{\mathsf{Y}}$ are not given, though there is evidence that it performs well in practice Biagioni2015.
 To analyze the cost of this algorithm, let $n=N_{1}=\cdots=N_{d}$.
@@
 7 Concluding remarks
-We end our survey with an outlook on the future of RNLA, which will continue to be an indispensable tool for scientific computing and data science in the face of increasingly large problems.
-We have seen throughout this work that randomization is not only useful for dimension reduction, but also admits algorithmic implementations that can be optimized for specific hardware or problem constraints, such as parallelization for GPU computing or factorizing streaming data.
-Randomization is especially critical for the development of scalable tensor algorithms.
-Future work may touch on several different aspects of computing fast and accurate matrix and tensor decompositions with randomized algorithms.
-Probabilistic guarantees are still needed for the resulting error of many randomized algorithms that are used in practice, as well as tighter bounds that more closely align with observed behavior.
-For example, it is generally an open question as to what the weakest assumptions are for a particular type of randomized DRM to achieve a desired approximation accuracy.
-The question can be further refined by the specific problem setting, e.g., the weakest assumptions to obtain a certain accuracy of rank-structured matrix decomposition or sketched least squares solutions.
-As stronger theoretical foundations are developed for new and existing randomized algorithms, high-performance software is also needed to fully leverage available computing resources.
-For example, there remains a need for standardized RNLA kernels and interfaces that can offer both practical speed and front-end accessibility, e.g., in Python or Julia.
-RNLA software libraries like RandBLAS or RandLAPACK are continually being updated with new functionality, such as mixed-precision arithmetic or sparse linear algebra packages.
-Additionally, there is an ongoing effort to develop high-performance implementations of randomized tensor algorithms that take advantage of GPU acceleration.
-This includes the growing body of work on tensor decompositions other than those presented here, some of which can break the curse of dimensionality.
-In summary, RNLA has become an established area of applied mathematics that will prove increasingly important as larger problems are tackled and the size and complexity of scientific data grow.
-As such, there remain many interesting open questions to investigate, which build on the fundamental concepts surveyed in this work.
+The primary purpose of this survey was to review a set of recently developed randomized algorithms for processing large scale tensor data.
+These new methods were in many cases inspired by analogous techniques for processing matrices that exploit randomized dimension reducing maps to accelerate computations.
+While such methods for the matrix case are reaching a certain state of maturity, the tensor environment is still evolving rapidly.
+In our survey, our focus was on techniques that are relatively straight-forward generalizations of algorithms designed for matrices.
+We did not attempt to describe the plethora of algorithms that are designed for more tensor-specific representations, such as, for instance, tensor trains.
+Nor did we cover techniques that are designed for tensors that are so large they they cannot be formed or stored explicitly.
+We observe that much of the success of RNLA stems from the fact that randomized algorithms often involve less data movement, and therefore execute faster on modern hardware than many traditional algorithms that were designed to minimize floating point operations.
+Such considerations have proven to be even more essential in the tensor case, given the often enormous size of the data sets under consideration, and the stronger tension between data being located close by in the tensor on the one hand, and being close in physical memory on the other.
+Acknowledgments: The work reported was supported by the Office of Naval Research (N00014-18-1-2354), by the National Science Foundation (DMS-2313434, DMS-2401889), and by the Department of Energy ASCR (DE-SC0025312).
+PGM further expresses gratitude for an academic sabbatical in 2025/26 that was made possible through support provided by the Simons Foundation, and by the Swedish Royal Academy of Sciences through the Institute Mittag-Leffler.